### v1

# The Dynamic Net Architecture: Learning Robust and Holistic Visual Representations Through Self-Organizing Networks

Abstract
We present a novel intelligent-system architecture called “Dynamic Net Architecture” (DNA) that relies on recurrence-stabilized networks and discuss it in application to vision. Our architecture models a (cerebral cortical) area wherein elementary feature neurons encode details of visual structures, and coherent nets of such neurons model holistic object structures. By interpreting smaller or larger coherent pieces of an area network as complex features, our model encodes hierarchical feature representations essentially different than artificial neural networks (ANNs).
DNA models operate on a dynamic connectionism principle, wherein neural activations stemming from initial afferent signals undergo stabilization through a self-organizing mechanism facilitated by Hebbian plasticity alongside periodically tightening inhibition. In contrast to ANNs, which rely on feed-forward connections and backpropagation of error, we posit that this processing paradigm leads to highly robust representations, as by employing dynamic lateral connections, irrelevant details in neural activations are filtered out, freeing further processing steps from distracting noise and premature decisions.
We empirically demonstrate the viability of the DNA by composing line fragments into longer lines and show that the construction of nets representing lines remains robust even with the introduction of up to $59\%$ noise at each spatial location. Furthermore, we demonstrate the model’s capability to reconstruct anticipated features from partially obscured inputs and that it can generalize to patterns not observed during training. In this work, we limit the DNA to one cortical area and focus on its internals while providing insights into a standalone area’s strengths and shortcomings. Additionally, we provide an outlook on how future work can implement invariant object recognition by combining multiple areas.
Keywords net fragments  $\cdot$ neural code  $\cdot$ early commitment  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neuronal networks  $\cdot$ neuroscience  $\cdot$ machine learning
1 Introduction
We present the first implementation of a novel intelligent-system architecture called “Dynamic Net Architecture” (DNA). We motivate it here by contrasting it to the architectures of artificial neural networks (ANNs) that play an essential role in analyzing high-dimensional data (LeCun et al., 2015; Schmidhuber, 2015; Stadelmann et al., 2018). An ANN is a static, deeply nested function optimized with respect to a static cost function. An input to an ANN is mapped to an output by a succession of nested functions arranged in layers. Consequently, representations have the form of a hierarchy that integrates simpler low-level features (outputs of inner functions) to more complex high-level features (outputs of outer functions). The representations computed by the inner functions are independent of the outputs of outer functions. We posit that this is one of the leading causes of the lack of robustness witnessed in ANNs (Windrim et al., 2016; Li et al., 2019; Wang et al., 2019; Rusak et al., 2020; Drenkow et al., 2022) since the decisions of inner functions cannot be adjusted by plausibility checks based on outer functions. This assertion finds support in the formal explanation provided by Goodfellow et al. (2015), who argue that adversarial attacks rely on perturbations below the threshold of visibility when these are summed over a sufficiently high input dimensionality (cf. Ilyas et al. 2019). Following Marr (2010), we call this phenomenon of lacking robustness due to sequential information processing “the fallacy of early commitment,” i.e., the inability to factor the outputs of the outer functions (high-level features) into the computation of the inner functions (low-level features), so that low-level decisions are in danger of misleading high-level decisions.
With the proposed DNA, we overcome this limitation by dynamically integrating local and global features. The DNA is formulated in terms of cortical areas (two-dimensional sheets of neurons) and projections between them. Connectivity within areas is dominated by short-range lateral (“recurrent”) connections, different areas are connected bidirectionally. The neurons of a given area are elementary feature detectors. Afferent connections from a given input pattern activate a fraction of all neurons in the area. Cyclically rising inhibition then silences the majority of the initially activated neurons, leaving only those active that are supporting each other with sufficient strength through excitatory lateral connections. This mutual support can only be realized within specifically structured consistent networks, “nets” (von der Malsburg, 2018) and have to be generated by statistical learning. Nets can span the entire visual field and connect feature neurons that are spatially distant through chains of locally connected intermediary neurons. Nets are flexibly composed of “net fragments” that correspond to often-occurring feature constellations. Each position of an area contains many fragments (analogous to the code-book of image compression schemes, Taubman and Marcellin (2002)) from which the input can choose. Unlike the pieces of a jigsaw puzzle, each net fragment fits to a choice of possible neighboring fragments (and each neuron can be part of several net fragments), so that a virtually infinite number of nets can be formed from a finite set of neurons and net fragments.
According to this principle, a two-phase process admits first a transient initial volley of activity, while in a second phase, most of this activity is silenced so that only mutually supporting neurons stay active. What is responding to the input of a cortical area are thus not individual neurons but rather organized nets. From Gestalt psychology (Wertheimer et al., 1938; Köhler, 1992), we know the visual system can rapidly discern holistic structures, that is, recognize global patterns as arrays of local features that conform consistently to certain “local Gestalt rules” Köhler (1992): from all the local features present in the input, only those are selected by the visual system that fit consistently into a coherent global pattern (coherent in the sense of being composed throughout of overlapping fragments). Thus, local decisions are conditioned on global patterns, whereas global patterns need for their constitution local features. We offer our nets of lateral connections as the mechanism that mediates between local and global levels: local Gestalt rules are implemented by net fragments, and as long as an input pattern can be covered throughout by fragments, the response it elicits is a coherent net.
The two-phase process described is able to remove noise and filter out irrelevant details. It thus avoids the fallacy of early commitment by conditioning neuron activation on integrability into a globally consistent net. An additional effect is that neurons that aren’t receiving afferent input but have sufficient lateral support might get activated, thus complementing figures or textures and allowing the interpretation of occluded objects.
The connectivity necessary for stable second-phase activity is very specific and needs to be generated by learning. The goal is to find and stabilize those patterns of afferent activation that stand out statistically. This is possible with the help of Hebbian plasticity (Hebb, 1949) that strengthens short-range excitatory connections between neurons often co-activated by the input. In the initial stages of learning, before the establishment of supporting connections, learning has to rely on phase one activity, that is, on the actual statistics of input activation. In the course of learning, however, plasticity has to be biased more and more in favor of those neurons that stand out by lateral support, neurons whose activity survives into phase two.
How do our cortical areas compare to the layers of ANNs? Whereas the latter are mainly there to accommodate a hierarchy of feature detectors of higher and higher complexity, already a single cortical area provides for such hierarchy in the form of nested net fragments of growing size (von der Malsburg et al., 2022b), a hierarchy that extends up to input-activated nets representing entire objects. One role of higher cortical areas is to permit the construction of representations invariant to transformations affecting lower areas (translation, rotation, scaling, etc.), while the connection between variant and invariant representations is established by structured dynamic sets of connections (see below).
Areas are dominated by short-range lateral connections, which abound in the cortex Stettler et al. (2002); Suzuki et al. (2023) but are ignored in multi-layer perceptrons (MLPs) (Prince, 2023) and transformer architectures (Vaswani et al., 2017). Lateral, recurrent connections play important roles both for function and for learning. As for function, they are the basis for the establishment of nets as high-level (e.g., object-global) representations by a collective computational mechanism Forrest (1990), conditioning global and local decisions on each other, as noted above. Such mediation between levels is not possible in ANNs with their purely feed-forward flow of information (which is unavoidable due to their learning principle of backpropagation of error (Rosenblatt, 1962; Linnainmaa, 1970; Rumelhart and McClelland, 1987)). Regarding learning, the restriction to connectivity patterns that are self-consistent (connectivity structure supporting activity states that stabilize this same connectivity structure) reduces the search space of learning by many orders of magnitude and has the potential to produce connectivity patterns representing novel structures never seen as examples. Moreover, the stabilization of connection weights based on local rules (rather than backpropagation through many layers) decisively relaxes precision requirements put on them.
In this work, we present the first implementation of a cortical area structured by fragments and apply it to a simple vision task, showcasing its ability to be robust against noise and perform figure completion. We believe that this conceptual work lays the foundation for a novel learning system that can fulfill complex tasks such as invariant object recognition when scaled up to multiple areas in subsequent research (see Section 2.3). Specifically, we make the following contributions:
  • We demonstrate the viability of net fragments as neural code, highlighting their role in enhancing the robustness of computational frameworks.
  • We provide a summary of the relevant neuroscientific hypotheses of net fragments and dynamic linking.
  • We translate these hypotheses into a computational framework compatible with recent machine learning advances.
  • We discuss the introduced concepts and describe how extending it with multiple areas can foster invariant object detection.
2 Related Work
As outlined above, deep neural networks exhibit a deficiency in robustness. In the scientific literature, robustness is typically defined as immunity to noise (additive noise, or subtractive noise in the form of missing information; see Cao et al. (2021); Neururer et al. (2024)) and adversarial attacks (Drenkow et al., 2022), as well as the capability of dealing with ambiguous inputs (Simmler et al., 2021) and generalization within/across domains (e.g., over image translation, rotation, changing lighting, etc.).
This work focuses on analyzing net fragments that are formed in each DNA-based area concerning their robustness to small perturbations. Related work pertaining to this aspect is summarized in Section 2.1, followed by a review of representation learning. The subsequent Section 2.3 examines systems utilizing correspondence-based mapping, which can cope well with ambiguous inputs and exhibit good generalization capabilities. These systems closely align conceptually with the DNA proposed in this work, making them highly relevant for future research, especially for scaling to a multi-area DNA.
2.1 Robustness to Noise
Adversarial attacks exploit deep networks’ vulnerability to noise by calculating slight perturbations capable of altering the model’s predictions for specific samples (Szegedy et al., 2014; Goodfellow et al., 2015; Papernot et al., 2016a; Carlini and Wagner, ; Deriu et al., 2022) or perturbations applicable across the entire dataset (Moosavi-Dezfooli et al., 2017). Adversarial attacks can also be devised without direct access to the model and its parameters, for example, by analyzing a model’s decision boundary (Brendel et al., 2018; Engstrom et al., 2019) or training an auxiliary model (Chen et al., 2017; Cisse et al., 2017; Sarkar et al., 2017).
Current approaches aim to counter these attacks by modifying the network’s architecture to increase robustness (Papernot et al., 2016b; Gao et al., 2017; Ros and Doshi-Velez, 2018), including adversarial examples to the training data (Moosavi-Dezfooli et al., 2016; Goodfellow et al., 2015), transform the inputs before being processed (Dziugaite et al., 2016; Bhagoji et al., 2018), or adding external models (Meng and Chen, 2017; Xu et al., 2018) or attack detectors (Amirian et al., 2018).
The lack of robustness also manifests itself in poor generalization between domains, where a change in the statistical distribution of data prevents the applicability of a model trained on a source domain to a slightly different target domain (Tuggener et al., 2022; Yan et al., 2024). Addressing such domain shifts typically involves domain alignment (Erfani et al., 2016; Ghifary et al., 2017; Rahman et al., 2020; Sager et al., 2022), meta-learning (Dou et al., 2019; Zhou et al., 2020; Du et al., 2020), or data augmentation (Carlucci et al., 2019; Zhang et al., 2020; Tuggener et al., 2024).
Despite notable advances in improving robustness and preventing adversarial attacks, existing methods primarily address symptoms by reducing the number of parameters (Papernot et al., 2016b), removing specific features (Gao et al., 2017), regularizing gradients (Ros and Doshi-Velez, 2018), augment or extend the training data (Moosavi-Dezfooli et al., 2016; Dziugaite et al., 2016; Tuggener et al., 2024), align domains (Erfani et al., 2016), or meta-learn features (Dou et al., 2019). In contrast, we here attempt to fundamentally rethink the principles underlying learning, network architecture, and information flow, potentially leading to more robust representations. In our experiments (see Section 6), we substantiate the efficacy of the proposed net fragments in effectively mitigating random Gaussian noise and dealing with occluded objects (a type of subtractive noise).
2.2 Representation Learning
Other research relevant to this work employs other design principles to enhance robustness. A crucial principle involves sparse neural activation, which enhances robustness relative to dense representations, as shown for deep neural networks (Guo et al., 2018; Liao et al., 2022; Timpl et al., 2022; Prince, 2023). This is particularly true for binary sparse distributions (Ahmad and Scheinkman, 2019). Element-wise comparison of two vectors, each with only a few activated neurons, proves to be highly robust. Given sufficiently large vector dimensionality, accidental similarity across different binary sparse representations is highly improbable (Ahmad and Hawkins, 2015).
Also in this work, we exploit binary sparse representations to bolster model robustness. In our system, sparsity is not imposed by specific constraints but by design along established principles, including lateral connections (Kothari and Agyepong, 1996; Corchado et al., 2003) and self-organization through local Hebbian updates (Hebb, 1949), coupled with inhibitory mechanisms for activity regulation (Abbott and Nelson, 2000; Luz and Shamir, 2012).
2.3 Correspondence-Based Mapping
A highly robust vision framework for object detection is based on correspondence mapping between pixel patterns within an image and a corresponding prototype (Wolfrum et al., 2008). Here, “corresponding” refers to neurons relating to the same point on the object’s surface. A pivotal concept in this framework are maps implemented as sets of projection fibers (connections that dynamically turn on and off) known as shifter circuits (Anderson and van Essen, 1987; Olshausen et al., 1993) that are composed of maplets (Zhu and von der Malsburg, 2004). These maps are activated when point-to-point correlations are detected between coherent nets in two cortical areas. The composite nets in the two cortical areas explicitly represent spatial relations, and a map between them can only be formed if they agree on this spatial structure. Once a forward mapping to a coherent model of a visual pattern is established, backward mapping to the primary visual domain can stabilize the modeled structure there by complementing missing elements or by lacking to support visual noise elements.
So far, correspondence-based mapping has been applied only to human face detection and identification (Lades et al., 1993; Wolfrum et al., 2008; Fernandes and von der Malsburg, 2015), in which image and object model match with little deformation. For general classes of visual structure, mapping has to tolerate deformation and must be based on the matching of robust feature constellations (Biederman and Kalocsai, 1997). Net fragments can be seen as an attempt to model these (von der Malsburg, 2014, 2018; von der Malsburg et al., 2022a). So far, however, net fragments have only been proposed as a theoretical concept and have not been modeled in any concrete fashion.
In this work our focus is, however, not on correspondence-based mapping, but on the learning of fragment-based net formation, hoping to thereby contribute to the establishment of this neural code as a sound conceptual framework translating concepts of biological vision to applications in next-generation machine learning systems.
3 Net Fragments for Visual Perception
In the following, we restrict the discussion of DNA to the visual system. Light fields captured by the human eyes are translated to retinal images that appear as two-dimensional arrays of activated neurons in the primary visual cortex (Grill-Spector and Malach, 2004). There, neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004) and interact through short-range lateral connections (Gilbert et al., 1990). Visual patterns that appear with significant frequency within an image patch lead, through Hebbian plasticity, to the formation of net fragments, as visualized in Figure 1.
Figure 1: Net fragments in a nutshell: The eyes observe a visual scene and activate neurons in the primary visual cortex. Neurons that support each other inside net fragments remain active, while neurons initially activated by afferent excitation but lacking such support are silenced by inhibition.
Among afferent input-activated neurons, only a few are part of net fragments so as to support each other’s activity by exchanging spikes along excitatory lateral connections (Stettler et al., 2002). Within milliseconds, neurons that have sufficient lateral support are singled out by neural inhibition (Vogels et al., 2011) from all others that also receive afferent input but lack lateral support.
All statistically significant visual textures within image patches with the radius of the range of lateral connections are represented by local connectivity patterns (such as in image compression algorithms (Wallace, 1991) by codebook vectors). To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (“convolutional” structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift.
A given feature can be part of a large number of patterns. To avoid interference between these patterns, each feature (defined in the primary visual area by a particular receptive field shape in a particular position) is represented by a number of “alternative neurons,” which are free to carry different connections  [^1] Alternative neurons may be seen as analogous to transformer’s Vaswani et al. (2017) token embedding vector components, which support the different connections between tokens represented by the query-key mechanism.. In this way, features can vary their connections to other features by activating different alternative neurons: they are connected by “dynamic links.” For more details on alternative neurons, please refer to Appendix A.
To project net fragments and coherent nets out from the mass of all neurons that are individually activated by the sensory input needs dynamic processing: After neurons are activated by afferent input, they are subjected to progressively increasing inhibition, which eventually silences all those of them that don’t have sufficient lateral support. As many fragments can survive inhibition only if supported by overlapping other fragments, inhibition-resistant responses constitute nets that are coherently composed of overlapping fragments.
These fragment-composed nets differ fundamentally from associative memories as formalized, for instance, by Hopfield (1982); Cohen and Grossberg (1983). In these, entire neural patterns, “assemblies,” are stored (in a single step) and can be retrieved only as a rigid whole, while nets are flexibly composed of overlapping fragments (which have been learned in a statistical process) and are flexibly assembled to model the input. We discuss this distinction in more detail in Appendix A.
4 Computational Framework
Figure 2: The dynamics of a single area within the proposed DNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer. In the upper part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the lateral connections within $S2$. In the lower part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
We here restrict ourselves to modeling this functionality within a single cortical area, although the significance of coherent nets in primary sensory areas can be fully appreciated only in the context of different interacting areas.
A single area of a DNA model comprising net fragments is shown in Figure 2(A). Our system is based on an input layer representing observed images by simple feature detectors (which are perhaps to be localized within layer IV of the primary visual cortex, see, for instance, Usrey et al. (2003)) and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$) whose neurons are thought to be also localized in the same primary visual cortex area, presumably within layers II and III.
Feature Extraction.
The system’s input is a retinal 2D image represented by variables $\boldsymbol{x}_{k,i}$ (gray box in Figure 2), where $k\in(1,...,C)$ denotes the channel index of the image ($C=1$ for grayscale, $C=3$ for RGB), and the composite index $i=(h,w)\in(H\times W)$ stands for vertical and horizontal position of pixels within the input image.
Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their “membrane potential” $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2). The membrane potential $\boldsymbol{a}^{(S1)}$ is calculated by applying a convolutional operation (LeCun et al., 1989) (yellow box $S1$ in Figure 2):
$\boldsymbol{a}^{(S1)}_{c,j}=\sum_{k\in C}\sum_{i\in l_{j}}\boldsymbol{W}^{(S1)}_{c,k,j-i}\cdot\boldsymbol{x}_{k,i}$ (1)
Here, $c\in C^{(S1)}$ denotes the feature channel, the compound index $j$, expanded as $j=\{h,w\}$, designates again spatial position $h\in H$, $w\in W$, and the compound index $i=\{h,w\}$ is running through all $(h,w)$-pairs in the two-dimensional range $l_{j}=(h,....,h+h^{(W1)};w,....,w+w^{(W1)})$ where $h^{(W1)}$, $w^{(W1)}$ denote the size of the convolutional filter. The size of the convolutional kernel used for feature extraction is $\boldsymbol{W}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times C\times h^{(W1)}\times w^{(W1)}}$.
Feature Binarization.
The membrane potential $\boldsymbol{a}^{(S1)}$, together with a bias parameter $b^{(S1)}$, determines the binary output of the neurons in $S1$, denoted as variable $\boldsymbol{y}_{c,j}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$, where $1$ stands for firing, 00 for silence. The binary neuron model with the input-output function $\mathrm{B}:\boldsymbol{a},b\rightarrow\boldsymbol{y}$ is defined as
$\boldsymbol{y}_{c,j}^{(S1)}=\mathrm{B}(\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)})=\begin{cases}1,&\text{ if }\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)}>0\\ 0,&\text{otherwise}\end{cases}$ (2)
with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (cells fire if the membrane potential is above this threshold).
Dynamics.
Stage $S2$ builds net fragments by allowing active neurons to support laterally connected neurons over T time steps (depicted as recurrent connection in Figure 2). A sample image is introduced into the system at time step $t=0$ to activate neurons and remains unchanged until time step $t=T$. Throughout this period, neuronal activity undergoes sparsification due to growing inhibition. Through the sparsification process, the system forms an object representation by leaving active a set of neurons each of which is supported by active neighbors, that is, a net.
Short-Range Lateral Connections.
Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feed-forward connections between different layers, we employ them here to implement short-range lateral connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments. Given lateral connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$. In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state.
The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by lateral excitation from within $S2$. Before digitization, they are modified by saturation, normalization, and inhibition (yellow box $S2$ in Figure 2). Before those modifications, its raw form is computed as:
$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Lateral Connection}}.$ (3)
Similar to $S1$, $c\in C^{(S2)}$ denotes the feature channel of $S2$, the compound index $j$ expands to $j=\{w,h\}$, $w\in W$, $h\in H$, and the compound index $i$ is running in the range $l_{j}=(h,....,h+h^{(W2)};w,....,w+w^{(W2)})$, with $h^{(W2)}$ and $w^{(W2)}$ being the size of the convolution kernel.
The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the lateral connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e. $C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments, see Apendix A). In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the lateral connections determines which neurons can be connected within stage 2. The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix B). The weights are updated using Hebbian Learning (details in Appendix C).
For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide lateral support:
$\boldsymbol{y}_{i}^{(S2)}[t=0]=0$ (4)
Single Weight-Kernel Implementation.
This entire process is visually depicted in the upper part (A) of Figure 2. For simplicity in our algorithm, and since we use the same kernel size for the forward and lateral connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(S2)}_{c,k,j-i}\cdot\left(\boldsymbol{y}_{k,i}^{(S1)};\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)$ (5)
The net fragments are implicitly contained in this processing as the active neurons whose activity state is stored in $\boldsymbol{y}^{(S1)}$ and their connection to other neurons.
Inhibition.
The raw neural excitation $\boldsymbol{a}^{\prime}$ computed in eq. 3 (or eq. 5) is to be subjected to saturation, inhibition, and normalization so it can be binarized using a pre-defined threshold. These steps of processing the raw membrane potential $\boldsymbol{a}^{\prime}$ of neurons proved necessary in our simulation experiments to achieve the qualitative function we aimed for. The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix D. The differences between activation levels of neurons are accentuated by setting
$\boldsymbol{a}^{(S2)}_{c,j}[t]=\max\left(\left(\boldsymbol{a}^{\text{norm }(S2)}_{c,j}[t]\right)^{\gamma},0\right),$ (6)
The coefficient $\gamma$ regulates the inhibition strength and increases with the time-steps of the inner loop according to $\gamma=1.2+0.2t$, making it harder for neurons to fire (as $\boldsymbol{a}^{\text{norm }(S2)}\leq 1$). The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive lateral support, while invalid activations are less supported. While the inhibition coefficient increases over time, many neurons eventually get suppressed, leaving only those active that support each other and form a consistent net. Finally, the membrane potential is converted to a binary activation, neurons becoming active when their inhibited membrane potential is above the activation bias $b^{(S2)}$, which serves to control the firing rate of neurons:
$\boldsymbol{y}_{c,j}^{(S2)}[t]=\mathrm{B}\left(\boldsymbol{a}^{(S2)}_{c,j}[t]-b^{(S2)}\right)$ (7)
The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the lateral connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed. To allow features to be part of multiple net fragments and represent various patterns, they are represented by a set of alternative neurons (details in Appendix A). All neurons within such sets are free to learn different lateral connections to other neurons. They are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero). This selection process process is described in the Appendix E.
5 Experiments
We evaluate our framework by assessing the efficacy of net fragments to filter out random Gaussian noise, the capacity to reconstruct occluded objects, i.e., the efficacy of dealing with subtractive noise, and its generalization capacity by composing small net fragments into object structures not observed during training.
5.1 Dataset
The dataset comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
The images are generated when required and their set may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$. To that end, we probabilistically flip neurons within each channel with a probability of up to $20\%$. Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location. To evaluate subtractive noise, we create discontinuous lines by deleting a line segment in the middle.
To demonstrate the generalization capability of DNA areas, we generate validation samples that differ from the ones used during training. Specifically, we produce kinked lines by generating four random points within the image and draw straight lines between these points.
6 Results
In the following, we show that the proposed net fragments can suppress Gaussian noise in Section 6.1 and can reconstruct removed pixels, see Section 6.2. A description of the feature extraction mechanism to obtain $\boldsymbol{y}^{(S1)}$ is in Appendix F, and details about the used parameters are in Appendix G. Additionally, Appendix H contains visualizations of the average support strength received by cells, and Appendix I a quantitative evaluation of suppressed noise.
6.1 Filtering Gaussian Noise
Figure 3: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by lateral connections of net fragments. In each image, the background is shown in black, and the four different features in distinct colors. The plots are created using a bias $b^{(S2)}=0.5$ and an inhibition coefficient $\gamma=1.2+0.2t$.
The illustration in Figure 3 visually demonstrates the system’s effective capability of filtering out isolated noise or small clusters thereof. Noise of varying degrees (from $0\%$ to $59\%$ noise probability at each spatial location - see Appendix 5.1) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments. While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix I, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
Figure 4: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed. However, when inhibition increases over time, even more noise is filtered out in the course of iteration. The plot is created after adding $10\%$ noise to each feature channel (sing a bias $b^{(S2)}=0.7$ and an inhibition coefficient $\gamma=1.2+0.2t$).
Figure 3 visualizes that neural activity simulating visual clutter without coherent structure is removed after processing for $5$ time-steps. Figure 4 depicts this systematic filtering over time. The additionally introduced clutter doesn’t activate overlapping net fragments and cannot withstand increasing inhibition. The figure illustrates a substantial noise reduction already at time step $t=0$ due to the filtering by matrix $\boldsymbol{W}^{(F)}$ and initial inhibition. In each subsequent time step, additional clutter is removed. This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons. We thus observe a chain reaction of gradual suppression of incoherent structure as well as the preservation of net structure on the basis of internal support.
6.2 Reconstructing Subtractive Noise
Figure 5: Partial occlusion as subtractive noise: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
If a substantial number of neurons activate and support other neurons within a net fragment, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction. Figure 5 shows how inactive neurons activate and thus demonstrates that net fragments can deal with subtractive noise (occluded patterns). This reconstruction works only reliably for up to $3$ missing pixels (see Appendix I). In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
Membrane Potential Represents Expected Feature Activation.
Figure 6: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the “activation probabilities” (on the right) after the first time step. The activation probability is close to $1$ where a line has been observed and around $50\%$ in the middle where no line was observed but expected. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ plays a crucial role. The potential is higher at spatial locations where line features are observed and lower where they are not observed but expected. Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 6. Modeling this ambiguity enables having different internal interpretations, such as whether “it is an interrupted line that should be reconstructed” or “there are two co-linear line segments.” We find that the activation map is especially reliable during the first time step, where it is calculated based on the observed input and the network’s expectation about local patterns. Since local interactions can trigger or deactivate neurons, the activation map changes over time, reducing its interpretability.
Increasing Inhibition Over Time Stabilizes Patterns.
Figure 7: Reconstruction of a horizontal line interrupted by $7$ pixels: The feature activation and the initial output activity at $t=0$ are interrupted but reconstructed within the two subsequent time steps ($t=1,2$). Afterward, some artifacts triggered at line ends are reduced due to increasing inhibition. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=1.8+0.1t$.
The reconstruction process unfolds across multiple sequential steps, as illustrated in Figure 7. Initially, at time step $t=0$, the net fragments represent the interrupted line. Subsequently, during the second time step ($t=1$), some of the removed pixels are reconstructed, and shortly after, at $t=2$, the entire line is represented by the net fragments. Thus, reconstruction can persist even amid increasing inhibition, given that inactive neurons receive adequate support. This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of certain neurons encourages others to follow.
6.3 Compositionality
A single DNA area models a feature hierarchy through nested net fragments of increasing size (von der Malsburg et al., 2022b). In this study, we demonstrate that these higher-level nets can be composed of local features (small net fragments), even though the underlying patterns had never been observed during training. We show this on the example of curved lines, which share only local features with the straight lines of the training set.
Figure 8: Input images of curved lines are shown on the left, the second and third columns show the corresponding feature and output activation without noise, and the fourth and fifth columns show the same feature and output activation with noise corruption (activation bias $b^{(S2)}=0.3$, inhibition coefficient $\gamma=0.6+0.1t$).
Figure 8 depicts four samples of curved lines with the corresponding neuronal activation within an area, once without noise and once with $18.5\%$ noise added to its afferent input. The cortical area is able to represent the curved lines as composite of net fragments learned from straight lines. Added unstructured clutter is efficiently removed, leading to robust input representation.
7 Discussion and Conclusions
The concepts proposed here may be compared to established neural theory. Nets are based on associative learning but differ fundamentally from Hebbian assemblies (see Appendix A) as described in (Hopfield, 1982; Cohen and Grossberg, 1983; Papadimitriou et al., 2020), which are monolithic and without inner structure, whereas nets can represent an infinity of patterns based on shared fragments and can deal with new entities based on homeomorphy between common compositional structure, thus enabling the systematicity and productivity characterizing cognitive operations (Fodor and Pylyshyn, 1988; von der Malsburg, 2023). In comparison to multi-layered perceptrons, coherent nets replace hierarchies of independent feature neurons by net fragments, share with them the potential for equivariance (shift symmetry), but in distinction to them support the construction of an invariant representation that doesn’t need to be learned object-by-object (see next section). With the transformer architecture (Vaswani et al., 2017), nets share the provision for finding non-local patterns through on-the-fly computation of lateral links.
A decisive difference between DNA and all deep-learning systems (LeCun et al., 2015; Schmidhuber, 2015; Stadelmann et al., 2019) is the mode of learning. ANNs typically rely on backpropagation of error where the goal is reaching consistency (minimal error) between system output and teaching signals. With DNA models, the goal is mutual consistency at each neuron, between the excitatory connections converging there, this consistency being measured by the probability for each connection to reach a highly active neuron when carrying a signal itself. Of all possible connectivity patterns, only a minute fraction has this property of self-consistency. This condition of self-consistency acts as a selection criterion for eligible networks and thus powerfully reduces the search space of learning.
Net fragments and the structured nets they build form a basis for the representation of hierarchical structures. They are effective for feature filtering and for preventing early commitment, which may be of fundamental importance for improving the robustness of current machine learning systems. So far, a neural code based on net fragments has merely been proposed as a theoretical concept (von der Malsburg et al., 2022b), while this study pioneers the implementation of net fragments in a concrete simulation.
Our experiments demonstrate that the proposed DNA architecture builds net fragments that effectively stabilize patterns, enhance robustness by eliminating visual clutter, and can reconstruct removed parts of known patterns. The DNA model can handle additive noise exceptionally well and eliminates up to $96\%$ of it while preserving the initial pattern, even in scenarios where up to $59\%$ of all locations are affected. Figure completion works well for line interrupts for segments of up to $3$ pixels, while for larger interruptions, the reconstruction rate decreases considerably. One reason for this is certainly due to the nature of the used 1D pattern: as it involves straight lines with a width of one pixel, typically, only a fraction of the neurons within the local 2D neighborhood of size $h^{(W2)}\times w^{(W2)}$ are active. With a larger line interrupt, the proportion of active neurons compared to neighboring neurons is vanishingly small, which makes reconstruction more difficult. It is expected that reconstruction will be much more robust with two-dimensional patterns.
We posit that this work, which focuses on the internals of a single cortical area, allows for scaling to multiple areas that will allow the implementation of more complex tasks (see next section). Nevertheless, although this work is intended to be a step towards implementing a robust learning framework and the proposed neural code exhibits promising characteristics, it still has weaknesses that should be addressed in future work. One such weakness is dependence on manual adjustment of the parameters $b$ and $\gamma$. The mechanism of signal control is biologically implausible. It should be replaced by a more elegant, theoretically justified, and self-adjusting mechanism.
8 Future Work
One important purpose of handing structure from one area to another is generalization, a prime example for which in vision is object recognition invariant to translation (and wider transformation groups). While our system only includes one “cortical” area, another area (that would correspond to the cortical inferotemporal cortex (Ito et al., 1995)) could contain invariant object models. In the context of DNA, we see invariant recognition as realized by a process in which an object-representing net in a primary sensory area $A$ forms and activates a homeomorphic net in a secondary area $B$. Two nets are homeomorphic if they contain the same feature types in the same spatial arrangement. Invariance is achieved if differently transformed mutually homeomorphic nets in $A$ activate the same net in $B$ under a homeomorphic mapping. Inter-areal homeomorphic maps are based on feature type-preserving connections, which in the brain are realized by axonal fibers [^2] With the help of several intermediate areas, the number of necessary connections can be brought down to realistic values (Anderson and van Essen, 1987; Olshausen et al., 1993; Wolfrum et al., 2008)).. In fact, connections should not only be feature type-preserving but also restricted to pairs of corresponding alternative neurons in $A$ and $B$, such that they carry not only feature information but also information on connectedness [^3] To further ease the formation of homeomorphic maps in $B$, a mechanism for mutual support between mapping-compatible connections has been proposed and successfully applied (see Section 2.3).. The generation of the net in $B$ is again the result of a two-phase process, the afferent connections from $A$ to $B$ transiently activating many neurons in $B$, most of which getting silenced in phase two, leaving only those neurons active that not only got afferent input but are also supported by lateral excitation within $B$. The net thus formed in $B$ is invariant to homeomorphic transformations in $A$. Object classification, the identification of the net in $B$ as a particular object type or recognition of individual objects needs a further area $C$ containing a fixed net that is activated based on approximate homeomorphy with the invariant net in $B$.
The result of the recognition process is the formation of a larger net composed of a net in $A$, a net in $B$, and one-to-one connections between corresponding neurons in $A$ and $B$. The self-consistency of this larger net is the basis for learning all participating connections, as all connections are stabilized (or generated) by their success in predicting activity in their target neuron.
Through inverse mapping, these stabilizing interactions can be re-imported to the previous area and thus help to discriminate the corresponding nets there from noise and clutter and bridge larger partial occlusions, which would realize the idea of predictive coding Keller and Mrsic-Flogel (2018).
An obvious first step in preparation for a larger-scale system of interacting areas is, however, the training of a single primary-sensory area with natural images to capture as net fragments all textures that occur in them with significant frequency. This project should include the extraction of such different sensory qualities as stereo depth, local motion or surface discontinuities. Net fragments and nets based on these would then have a chance to relate to intrinsic surface properties such as form.
Funding
This research has been funded by the Canton of Zurich, Switzerland, through the Digitalization Initiative of the Canton of Zurich (DIZH) Fellowship project “Stability of self-organizing net fragments as inductive bias for next-generation deep learning.”
Data Availability Statement
The code to generate the data as well as to reproduce the results can be found after the publication of this work on Github at https://github.com/sagerpascal/dynamic-link-architecture.
Appendix
Appendix A Net Fragments vs. Associative Memory, Alternative Neurons
Learning of lateral connections to form net fragments shares aspects with associative memory (Hopfield, 1982; Cohen and Grossberg, 1983). There are, however, important differences. In an associative memory system, neural patterns are stored at once in their entirety with the help of synaptic plasticity, any pair of active neurons within a pattern strengthening their mutual excitatory connections. The resulting connectivity turns each stored pattern into an attractor state that can be recalled (dynamically activated) starting from initial states that come near to it. Such stored and dynamically stabilized neural patterns are often called “assemblies” (Hebb, 1949), and are discussed to this day as fundamental to the function of the brain, see for instance (Papadimitriou et al., 2020).
There are two problems with the concept of associative memory. One is rather fundamental in that a pattern is monolithic and can be recalled only as a rigid whole. This is due to the indiscriminate stickiness with which any pair of simultaneously active neurons is tied together by synaptic plasticity in the one-time storage process. Fragments, in distinction, are formed in a protracted statistical process in which connections are gradually established between neurons that are co-active with statistical significance. Due to the spatial deformation of visual patterns, this can only be the case for neurons relating to neighboring visual points. DNA models distinguish insignificant from significant connections, only the latter being permitted to grow, which they do under competition until reaching equilibrium. As a result, the system doesn’t store rigid global patterns but is able to dynamically stabilize an unlimited number of global states that are composed combinatorially out of overlapping fragments.
The other, more technical, problem of associative memory is that stored states have to be statistically independent (“orthogonal”) to avoid cross-talk between them. (In a simple example of cross-talk, a neuron $c_{1}$ gets connected at different times with neurons $c_{2},c_{3},c_{4}$, although the pattern $\{c_{1},c_{2},c_{3},c_{4}\}$ doesn’t occur with statistical significance during learning. As a result, $c_{1}$ may be activated erroneously if $\{c_{2},c_{3},c_{4}\}$ is active as part of a larger pattern.) Sparsity (according to which a very small fraction of all neurons are active within any given pattern) has been proposed as a general means to avoid cross-talk and to ensure statistical independence of stored patterns (French, 1999).
DNA models are freed from this constraint by the introduction of “alternative neurons,” neurons that are activated by the same input feature but are free to learn different lateral connections to other neurons. (In our example, neuron $c_{1}$ would be replaced by a set $\{c_{1},c_{1^{\prime}},c_{1^{\prime\prime}}\}$ of alternative neurons, which could then independently learn the connections $c_{1}-c_{2}$, $c_{1^{\prime}}-c_{3}$ and $c_{1^{\prime\prime}}-c_{4}$, thus avoiding the above cross-talk.) In the biological case, there is ample room for alternative neurons. There are, for instance, a hundred times more neurons in the primary visual cortex compared to the number of fibers coming out of the retina (Leuba and Kraftsik, 1994).
Appendix B Weight Initialization
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are copied into the $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative cells.
The lateral connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$. When Hebbian plasticity is employed on the weight matrix $\boldsymbol{W}^{(L)}$, these connections ought to grow (as neurons tend to keep their activation state between time steps), but due to the constraint of keeping all connection weights within the range $(0,1)$ (refer to equation 8), they stay put at $1$.
Appendix C Hebbian Plasticity
After processing for $T$ time steps both matrices $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ undergo Hebbian updates (Hebb, 1949):
$\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\alpha\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (8)
where $\alpha$ represents the learning rate, and $\boldsymbol{\rho}_{avg}$ denotes the average correlation between input and output. This average is determined for a given activation over all instances of a given pair of neurons $c,c^{\prime}$ in different positions of the image domain. To compute $\boldsymbol{\rho}_{avg}$, we count the number of instances in which both neurons are on, subtract from it the number of instances where one neuron is on, and the other is off, and divide that sum by the number of all positive and negative instances. The actual modification eq. (8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
The network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant. Consequently, the same lateral connection is applied across multiple output and input neurons. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects. If the resulting average is positive, the corresponding connection increases by $\alpha$ as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced by $\alpha$.
Appendix D Normalization
We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$. These steps are functionally motivated and not by reference to biology.
We first apply a size limit:
$\boldsymbol{a}^{\prime\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
According to this formula, $\boldsymbol{a}^{\prime\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$. (In our experiments we found the value $\lambda=1.3\cdot\frac{h^{(W2)}+w^{(W2)}}{2}$ to work well). This saturation helps to mitigate imbalances between network fragments with differing numbers of participating neurons (as with our stimuli, line segments, where many more feature neurons activate around the end-points than in the middle of segments).
In the next step, we introduce the normalization
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
This normalization confines activations to the interval $(0,1)$, where neurons receiving maximal support are mapped to $1$, while those with lesser support are mapped to correspondingly smaller values. This normalization relative to the maximal value that itself is growing during learning instead of relative to a constant parameter like $\lambda$ is important to give neurons initially (when lateral connections are still small) a chance to fire and later, when lateral links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
Appendix E Alternative Neuron Selection
Only one of a set of alternative neurons is allowed to be active. The selection process is determined by evaluating the correlation between feature patches and alternative kernel filters. Initially, patches matching the size of the feature kernel are extracted from every position in the input $(\boldsymbol{y}^{(S1)};\boldsymbol{y}^{(S2)}[t-1])$. Subsequently, the correlation of these patches with the feature kernels is computed, and at each location, the alternative kernel exhibiting the highest correlation is chosen. Consequently, multiple features may be active at each location, but only one of the competing feature channels — the one best matching a given input patch — achieves activation. At initialization (see Appendix B), where all alternative channels fit equally well to a given local pattern, the first channel is selected. However, after applying a Hebbian update, this channel starts specializing in recognizing a specific local pattern. Consequently, the activation of subsequent local patterns may either trigger the same channel if they have a high correlation or activate another channel better suited to recognizing the new pattern.
Appendix F Feature Extraction
Figure 9: The hand-crafted convolutional kernels of the feature extractor.
The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that, given our specific elementary feature types (cf. Fig. 9), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (“stride”) of $1$ (assuming 00 as input signal beyond the image border: “zero padding”) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using binary neurons $\mathrm{B}()$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 2).
Appendix G Parameters
We train the model with a learning rate of $\alpha=0.2$ for $100$ training cycles (epochs). The model uses $n_{a}=10$ alternative neurons with lateral connections spanning a kernel of $11\times 1$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix F), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot n_{a}=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix H Learned Support Strength
Figure 10: The average support active and inactive neurons receive during training, captured by $\boldsymbol{a}^{\prime}_{j}$. The support strength is shown on the y-axis, and the training progress (epoch) is on the x-axis.
A crucial parameter for gauging the efficacy of net fragments involves quantifying the difference in support received by active and inactive neurons. During the initial stages of training, active neurons receive a support of $1$, and inactive neurons receive a support of 00. The support increases during training as net fragments are formed. Figure 10 visually represents the average support for active and inactive neurons, captured by $\boldsymbol{a}^{\prime}_{j}$ (see eqs. 3, 5). Notably, the graph demonstrates a widening gap between the support for active and inactive neurons. This discrepancy suggests an enhancement in the robustness of statistically relevant patterns over time as active neurons receive increased support and the support of inactive neurons remains close to 00.
Appendix I Quantifying Noise Filtration
Our evaluations focus on the efficacy of net fragments in suppressing added Gaussian noise and complementing missing or occluded patterns. For both evaluation experiments, we report recall and precision: Recall is defined as the proportion of neurons that maintain activity (on) when noise is added, and precision quantifies the extent to which neurons activated in the presence of noise were also active without noise. To assess the system’s noise reduction capability in the case of Gaussian noise, we additionally measure the percentage of flipped neurons that revert to their original activation state, referring to this measurement as the noise reduction rate. To assess subtractive noise, we measure the similarity between the original (created without noise) and reconstructed (created with noise) feature activations at the spatial locations where pixels have been removed. We call this metric the feature reconstruction rate.
Filtering Gaussian Noise.
Figure 11: System performance as function of the noise level. First row: noise reduction rate; second row: recall; third row: precision. Columns differ in the inhibition coefficient $\gamma$ that models inhibition. The colors within each plot represent different activation biases $b^{(S2)}$.
Figure 11 displays noise reduction rates, recall, and precision across different parameter configurations for different levels of added noise. The metrics demonstrate that increasing the inhibition coefficient or the activation bias results in enhanced noise filtration.
[…]
Rconstructing Subtractive Noise.
Figure 12: Reconstruction of partial occlusion. First row: feature reconstruction rate; second row: recall; third row: precision. Columns are distinguished by the inhibition coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$.
Figure 12 displays the feature reconstruction rate, precision, and recall for removed line fragments of different lengths. Having a lower activation bias or inhibition coefficient $\gamma$ leads to a better reconstruction of removed patterns, as these settings allow neurons with less support to activate.
Interpreting Metrics.
While the reported values for precision and recall may appear low, we argue that these fragments are of good quality and point out that precision and recall metrics have to be interpreted differently than in a classification context: A net fragment consisting of many neurons characterizes (a part of) an object and maintains this characterization even when many of its neurons are inversed (and precision and recall are low). This aligns with the findings of Ahmad and Hawkins (2015), highlighting that binary distributed activations demonstrate high robustness, retaining accurate interpretation even in the presence of numerous flipped neurons.
In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$. Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines. Hence, we posit that various of the tested parameter configurations are well suited as they permit this distinction.

### v2

@@
-The Dynamic Net Architecture: Learning Robust and Holistic Visual Representations Through Self-Organizing Networks
+The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns
 Abstract
-We present a novel intelligent-system architecture called “Dynamic Net Architecture” (DNA) that relies on recurrence-stabilized networks and discuss it in application to vision.
-Our architecture models a (cerebral cortical) area wherein elementary feature neurons encode details of visual structures, and coherent nets of such neurons model holistic object structures.
-By interpreting smaller or larger coherent pieces of an area network as complex features, our model encodes hierarchical feature representations essentially different than artificial neural networks (ANNs).
-DNA models operate on a dynamic connectionism principle, wherein neural activations stemming from initial afferent signals undergo stabilization through a self-organizing mechanism facilitated by Hebbian plasticity alongside periodically tightening inhibition.
-In contrast to ANNs, which rely on feed-forward connections and backpropagation of error, we posit that this processing paradigm leads to highly robust representations, as by employing dynamic lateral connections, irrelevant details in neural activations are filtered out, freeing further processing steps from distracting noise and premature decisions.
-We empirically demonstrate the viability of the DNA by composing line fragments into longer lines and show that the construction of nets representing lines remains robust even with the introduction of up to $59\%$ noise at each spatial location.
-Furthermore, we demonstrate the model’s capability to reconstruct anticipated features from partially obscured inputs and that it can generalize to patterns not observed during training.
-In this work, we limit the DNA to one cortical area and focus on its internals while providing insights into a standalone area’s strengths and shortcomings.
-Additionally, we provide an outlook on how future work can implement invariant object recognition by combining multiple areas.
-Keywords net fragments  $\cdot$ neural code  $\cdot$ early commitment  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neuronal networks  $\cdot$ neuroscience  $\cdot$ machine learning
+Nets, cooperative networks of neurons, have been proposed as format for the representation of sensory signals, as physical implementation of the Gestalt phenomenon and as solution to the neural binding problem, while the direct interaction between nets by structure-sensitive matching has been proposed as basis for object-global operations such as object detection.
+The nets are flexibly composed of overlapping net fragments, which are learned from statistical regularities of sensory input.
+We here present the cooperative network architecture (CNA), a concrete model that learns such net structure to represent input patterns and deals robustly with noise, deformation, and out-of-distribution data, thus laying the groundwork for a novel neural architecture.
+Keywords net fragments  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ cooperativity  $\cdot$ neuroscience  $\cdot$ machine learning
 1 Introduction
-We present the first implementation of a novel intelligent-system architecture called “Dynamic Net Architecture” (DNA).
-We motivate it here by contrasting it to the architectures of artificial neural networks (ANNs) that play an essential role in analyzing high-dimensional data (LeCun et al., 2015; Schmidhuber, 2015; Stadelmann et al., 2018).
-An ANN is a static, deeply nested function optimized with respect to a static cost function.
-An input to an ANN is mapped to an output by a succession of nested functions arranged in layers.
-Consequently, representations have the form of a hierarchy that integrates simpler low-level features (outputs of inner functions) to more complex high-level features (outputs of outer functions).
-The representations computed by the inner functions are independent of the outputs of outer functions.
-We posit that this is one of the leading causes of the lack of robustness witnessed in ANNs (Windrim et al., 2016; Li et al., 2019; Wang et al., 2019; Rusak et al., 2020; Drenkow et al., 2022) since the decisions of inner functions cannot be adjusted by plausibility checks based on outer functions.
-This assertion finds support in the formal explanation provided by Goodfellow et al.
-(2015), who argue that adversarial attacks rely on perturbations below the threshold of visibility when these are summed over a sufficiently high input dimensionality (cf.
-Ilyas et al.
-2019).
-Following Marr (2010), we call this phenomenon of lacking robustness due to sequential information processing “the fallacy of early commitment,” i.e., the inability to factor the outputs of the outer functions (high-level features) into the computation of the inner functions (low-level features), so that low-level decisions are in danger of misleading high-level decisions.
-With the proposed DNA, we overcome this limitation by dynamically integrating local and global features.
-The DNA is formulated in terms of cortical areas (two-dimensional sheets of neurons) and projections between them.
-Connectivity within areas is dominated by short-range lateral (“recurrent”) connections, different areas are connected bidirectionally.
-The neurons of a given area are elementary feature detectors.
-Afferent connections from a given input pattern activate a fraction of all neurons in the area.
-Cyclically rising inhibition then silences the majority of the initially activated neurons, leaving only those active that are supporting each other with sufficient strength through excitatory lateral connections.
-This mutual support can only be realized within specifically structured consistent networks, “nets” (von der Malsburg, 2018) and have to be generated by statistical learning.
-Nets can span the entire visual field and connect feature neurons that are spatially distant through chains of locally connected intermediary neurons.
-Nets are flexibly composed of “net fragments” that correspond to often-occurring feature constellations.
-Each position of an area contains many fragments (analogous to the code-book of image compression schemes, Taubman and Marcellin (2002)) from which the input can choose.
-Unlike the pieces of a jigsaw puzzle, each net fragment fits to a choice of possible neighboring fragments (and each neuron can be part of several net fragments), so that a virtually infinite number of nets can be formed from a finite set of neurons and net fragments.
-According to this principle, a two-phase process admits first a transient initial volley of activity, while in a second phase, most of this activity is silenced so that only mutually supporting neurons stay active.
-What is responding to the input of a cortical area are thus not individual neurons but rather organized nets.
-From Gestalt psychology (Wertheimer et al., 1938; Köhler, 1992), we know the visual system can rapidly discern holistic structures, that is, recognize global patterns as arrays of local features that conform consistently to certain “local Gestalt rules” Köhler (1992): from all the local features present in the input, only those are selected by the visual system that fit consistently into a coherent global pattern (coherent in the sense of being composed throughout of overlapping fragments).
-Thus, local decisions are conditioned on global patterns, whereas global patterns need for their constitution local features.
-We offer our nets of lateral connections as the mechanism that mediates between local and global levels: local Gestalt rules are implemented by net fragments, and as long as an input pattern can be covered throughout by fragments, the response it elicits is a coherent net.
-The two-phase process described is able to remove noise and filter out irrelevant details.
-It thus avoids the fallacy of early commitment by conditioning neuron activation on integrability into a globally consistent net.
-An additional effect is that neurons that aren’t receiving afferent input but have sufficient lateral support might get activated, thus complementing figures or textures and allowing the interpretation of occluded objects.
-The connectivity necessary for stable second-phase activity is very specific and needs to be generated by learning.
-The goal is to find and stabilize those patterns of afferent activation that stand out statistically.
-This is possible with the help of Hebbian plasticity (Hebb, 1949) that strengthens short-range excitatory connections between neurons often co-activated by the input.
-In the initial stages of learning, before the establishment of supporting connections, learning has to rely on phase one activity, that is, on the actual statistics of input activation.
-In the course of learning, however, plasticity has to be biased more and more in favor of those neurons that stand out by lateral support, neurons whose activity survives into phase two.
-How do our cortical areas compare to the layers of ANNs?
-Whereas the latter are mainly there to accommodate a hierarchy of feature detectors of higher and higher complexity, already a single cortical area provides for such hierarchy in the form of nested net fragments of growing size (von der Malsburg et al., 2022b), a hierarchy that extends up to input-activated nets representing entire objects.
-One role of higher cortical areas is to permit the construction of representations invariant to transformations affecting lower areas (translation, rotation, scaling, etc.), while the connection between variant and invariant representations is established by structured dynamic sets of connections (see below).
-Areas are dominated by short-range lateral connections, which abound in the cortex Stettler et al.
-(2002); Suzuki et al.
-(2023) but are ignored in multi-layer perceptrons (MLPs) (Prince, 2023) and transformer architectures (Vaswani et al., 2017).
-Lateral, recurrent connections play important roles both for function and for learning.
-As for function, they are the basis for the establishment of nets as high-level (e.g., object-global) representations by a collective computational mechanism Forrest (1990), conditioning global and local decisions on each other, as noted above.
-Such mediation between levels is not possible in ANNs with their purely feed-forward flow of information (which is unavoidable due to their learning principle of backpropagation of error (Rosenblatt, 1962; Linnainmaa, 1970; Rumelhart and McClelland, 1987)).
-Regarding learning, the restriction to connectivity patterns that are self-consistent (connectivity structure supporting activity states that stabilize this same connectivity structure) reduces the search space of learning by many orders of magnitude and has the potential to produce connectivity patterns representing novel structures never seen as examples.
-Moreover, the stabilization of connection weights based on local rules (rather than backpropagation through many layers) decisively relaxes precision requirements put on them.
-In this work, we present the first implementation of a cortical area structured by fragments and apply it to a simple vision task, showcasing its ability to be robust against noise and perform figure completion.
-We believe that this conceptual work lays the foundation for a novel learning system that can fulfill complex tasks such as invariant object recognition when scaled up to multiple areas in subsequent research (see Section 2.3).
-Specifically, we make the following contributions:
-  • We demonstrate the viability of net fragments as neural code, highlighting their role in enhancing the robustness of computational frameworks.
-  • We provide a summary of the relevant neuroscientific hypotheses of net fragments and dynamic linking.
-  • We translate these hypotheses into a computational framework compatible with recent machine learning advances.
-  • We discuss the introduced concepts and describe how extending it with multiple areas can foster invariant object detection.
-2 Related Work
-As outlined above, deep neural networks exhibit a deficiency in robustness.
-In the scientific literature, robustness is typically defined as immunity to noise (additive noise, or subtractive noise in the form of missing information; see Cao et al.
-(2021); Neururer et al.
-(2024)) and adversarial attacks (Drenkow et al., 2022), as well as the capability of dealing with ambiguous inputs (Simmler et al., 2021) and generalization within/across domains (e.g., over image translation, rotation, changing lighting, etc.).
-This work focuses on analyzing net fragments that are formed in each DNA-based area concerning their robustness to small perturbations.
-Related work pertaining to this aspect is summarized in Section 2.1, followed by a review of representation learning.
-The subsequent Section 2.3 examines systems utilizing correspondence-based mapping, which can cope well with ambiguous inputs and exhibit good generalization capabilities.
-These systems closely align conceptually with the DNA proposed in this work, making them highly relevant for future research, especially for scaling to a multi-area DNA.
-2.1 Robustness to Noise
-Adversarial attacks exploit deep networks’ vulnerability to noise by calculating slight perturbations capable of altering the model’s predictions for specific samples (Szegedy et al., 2014; Goodfellow et al., 2015; Papernot et al., 2016a; Carlini and Wagner, ; Deriu et al., 2022) or perturbations applicable across the entire dataset (Moosavi-Dezfooli et al., 2017).
-Adversarial attacks can also be devised without direct access to the model and its parameters, for example, by analyzing a model’s decision boundary (Brendel et al., 2018; Engstrom et al., 2019) or training an auxiliary model (Chen et al., 2017; Cisse et al., 2017; Sarkar et al., 2017).
-Current approaches aim to counter these attacks by modifying the network’s architecture to increase robustness (Papernot et al., 2016b; Gao et al., 2017; Ros and Doshi-Velez, 2018), including adversarial examples to the training data (Moosavi-Dezfooli et al., 2016; Goodfellow et al., 2015), transform the inputs before being processed (Dziugaite et al., 2016; Bhagoji et al., 2018), or adding external models (Meng and Chen, 2017; Xu et al., 2018) or attack detectors (Amirian et al., 2018).
-The lack of robustness also manifests itself in poor generalization between domains, where a change in the statistical distribution of data prevents the applicability of a model trained on a source domain to a slightly different target domain (Tuggener et al., 2022; Yan et al., 2024).
-Addressing such domain shifts typically involves domain alignment (Erfani et al., 2016; Ghifary et al., 2017; Rahman et al., 2020; Sager et al., 2022), meta-learning (Dou et al., 2019; Zhou et al., 2020; Du et al., 2020), or data augmentation (Carlucci et al., 2019; Zhang et al., 2020; Tuggener et al., 2024).
-Despite notable advances in improving robustness and preventing adversarial attacks, existing methods primarily address symptoms by reducing the number of parameters (Papernot et al., 2016b), removing specific features (Gao et al., 2017), regularizing gradients (Ros and Doshi-Velez, 2018), augment or extend the training data (Moosavi-Dezfooli et al., 2016; Dziugaite et al., 2016; Tuggener et al., 2024), align domains (Erfani et al., 2016), or meta-learn features (Dou et al., 2019).
-In contrast, we here attempt to fundamentally rethink the principles underlying learning, network architecture, and information flow, potentially leading to more robust representations.
-In our experiments (see Section 6), we substantiate the efficacy of the proposed net fragments in effectively mitigating random Gaussian noise and dealing with occluded objects (a type of subtractive noise).
-2.2 Representation Learning
-Other research relevant to this work employs other design principles to enhance robustness.
-A crucial principle involves sparse neural activation, which enhances robustness relative to dense representations, as shown for deep neural networks (Guo et al., 2018; Liao et al., 2022; Timpl et al., 2022; Prince, 2023).
-This is particularly true for binary sparse distributions (Ahmad and Scheinkman, 2019).
-Element-wise comparison of two vectors, each with only a few activated neurons, proves to be highly robust.
-Given sufficiently large vector dimensionality, accidental similarity across different binary sparse representations is highly improbable (Ahmad and Hawkins, 2015).
-Also in this work, we exploit binary sparse representations to bolster model robustness.
-In our system, sparsity is not imposed by specific constraints but by design along established principles, including lateral connections (Kothari and Agyepong, 1996; Corchado et al., 2003) and self-organization through local Hebbian updates (Hebb, 1949), coupled with inhibitory mechanisms for activity regulation (Abbott and Nelson, 2000; Luz and Shamir, 2012).
-2.3 Correspondence-Based Mapping
-A highly robust vision framework for object detection is based on correspondence mapping between pixel patterns within an image and a corresponding prototype (Wolfrum et al., 2008).
-Here, “corresponding” refers to neurons relating to the same point on the object’s surface.
-A pivotal concept in this framework are maps implemented as sets of projection fibers (connections that dynamically turn on and off) known as shifter circuits (Anderson and van Essen, 1987; Olshausen et al., 1993) that are composed of maplets (Zhu and von der Malsburg, 2004).
-These maps are activated when point-to-point correlations are detected between coherent nets in two cortical areas.
-The composite nets in the two cortical areas explicitly represent spatial relations, and a map between them can only be formed if they agree on this spatial structure.
-Once a forward mapping to a coherent model of a visual pattern is established, backward mapping to the primary visual domain can stabilize the modeled structure there by complementing missing elements or by lacking to support visual noise elements.
-So far, correspondence-based mapping has been applied only to human face detection and identification (Lades et al., 1993; Wolfrum et al., 2008; Fernandes and von der Malsburg, 2015), in which image and object model match with little deformation.
-For general classes of visual structure, mapping has to tolerate deformation and must be based on the matching of robust feature constellations (Biederman and Kalocsai, 1997).
-Net fragments can be seen as an attempt to model these (von der Malsburg, 2014, 2018; von der Malsburg et al., 2022a).
-So far, however, net fragments have only been proposed as a theoretical concept and have not been modeled in any concrete fashion.
-In this work our focus is, however, not on correspondence-based mapping, but on the learning of fragment-based net formation, hoping to thereby contribute to the establishment of this neural code as a sound conceptual framework translating concepts of biological vision to applications in next-generation machine learning systems.
-3 Net Fragments for Visual Perception
-In the following, we restrict the discussion of DNA to the visual system.
-Light fields captured by the human eyes are translated to retinal images that appear as two-dimensional arrays of activated neurons in the primary visual cortex (Grill-Spector and Malach, 2004).
-There, neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004) and interact through short-range lateral connections (Gilbert et al., 1990).
-Visual patterns that appear with significant frequency within an image patch lead, through Hebbian plasticity, to the formation of net fragments, as visualized in Figure 1.
-Figure 1: Net fragments in a nutshell: The eyes observe a visual scene and activate neurons in the primary visual cortex.
-Neurons that support each other inside net fragments remain active, while neurons initially activated by afferent excitation but lacking such support are silenced by inhibition.
-Among afferent input-activated neurons, only a few are part of net fragments so as to support each other’s activity by exchanging spikes along excitatory lateral connections (Stettler et al., 2002).
-Within milliseconds, neurons that have sufficient lateral support are singled out by neural inhibition (Vogels et al., 2011) from all others that also receive afferent input but lack lateral support.
-All statistically significant visual textures within image patches with the radius of the range of lateral connections are represented by local connectivity patterns (such as in image compression algorithms (Wallace, 1991) by codebook vectors).
-To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (“convolutional” structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift.
-A given feature can be part of a large number of patterns.
-To avoid interference between these patterns, each feature (defined in the primary visual area by a particular receptive field shape in a particular position) is represented by a number of “alternative neurons,” which are free to carry different connections  [^1] Alternative neurons may be seen as analogous to transformer’s Vaswani et al.
-(2017) token embedding vector components, which support the different connections between tokens represented by the query-key mechanism..
-In this way, features can vary their connections to other features by activating different alternative neurons: they are connected by “dynamic links.” For more details on alternative neurons, please refer to Appendix A.
-To project net fragments and coherent nets out from the mass of all neurons that are individually activated by the sensory input needs dynamic processing: After neurons are activated by afferent input, they are subjected to progressively increasing inhibition, which eventually silences all those of them that don’t have sufficient lateral support.
-As many fragments can survive inhibition only if supported by overlapping other fragments, inhibition-resistant responses constitute nets that are coherently composed of overlapping fragments.
-These fragment-composed nets differ fundamentally from associative memories as formalized, for instance, by Hopfield (1982); Cohen and Grossberg (1983).
-In these, entire neural patterns, “assemblies,” are stored (in a single step) and can be retrieved only as a rigid whole, while nets are flexibly composed of overlapping fragments (which have been learned in a statistical process) and are flexibly assembled to model the input.
-We discuss this distinction in more detail in Appendix A.
+Figure 1: Representation of Objects by Nets.
+Objects are captured in their entirety by coherent nets that are composed of overlapping net fragments (schematically symbolized by white circles in the right panel).
+After training on natural images, each patch of visual space contains a complement of net fragments that represent textures that have been encountered with statistical significance.
+Net fragments overlap neuron-wise, and those activated by an object coalesce into a coherent net.
+The conundrum (stated in Olshausen and Field (2005)) that object contours can often not be found with the help of edge detectors due to lack of gray-level contrast (as in places inside the square in the left panel) may find its solution in that contours are defined as borders of the nets covering objects (or covering the background).
+Figure adapted from Olshausen and Field (2005), with permission.
+Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Simmler et al., 2021).
+While many recent AI systems predominantly favor neural representations over symbolic ones, we follow a proposal by von der Malsburg et al.
+(2022) and present the cooperative network architecture (CNA), an architecture that uses structured networks of mutually (‘recurrently’) connected neurons as representation of sensory signals.
+Using structured networks to encode information brings a diverse set of advantages, including high robustness, ability of figure reconstruction, and representing of out-of-domain data, laying the foundations to build novel object recognition frameworks (see Section 6).
+The networks in the CNA are called nets and are dynamically assembled from overlapping net fragments, where net fragments are connectivity patterns that are learned by synaptic plasticity based on the statistics of sensory input signals.
+A major argument in favor of this data architecture is the possibility (as described repeatedly by Hinton (1981); von der Malsburg (1981); Anderson and Van Essen (1987); Kree and Zippelius (1988); Olshausen et al.
+(1995); Wiskott et al.
+(1997); Wolfrum et al.
+(2008); Memisevic and Hinton (2010); Arathorn et al.
+(2013)) of solving the invariant object recognition problem by structure-preserving mapping between nets representing variant images and invariant models of objects.
+The representation of image structure by recurrently connected nets formed between elementary feature detector neurons is a fundamental deviation from the current artificial neural networks (ANNs) paradigm, according to which image structure is represented by sets of complex-feature neurons activated by purely feedforward connections (see Section 2.3 for a comparison of the two approaches).
+In the cited models of invariant object recognition, the formation of object-representing networks within the image domain (the primary cortical area where sensory signals arrive – here called ‘area’) has been sidestepped by letting spatial neighborhood stand in for connectedness.
+Neighborhood, however, is too weak a condition to qualify neurons as having responded to the same object (which is, for instance, apparent when looking at an object through foliage).
+Following von der Malsburg et al.
+(2022), we rather assume that for sets of neurons to be admitted into a net, they have to have been co-activated frequently in the past by structured visual input and consequently have been tied together by Hebbian synaptic plasticity (Hebb, 1949) into net fragments.
+Net fragments play a role somewhat related to the code-book vectors of some image compression schemes (Taubman and Marcellin, 2002) in representing statistically dominant texture patterns within any patch of the visual field; the formation of net fragments is somewhat related to, though fundamentally different from, associative memory (Hopfield, 1982), see Section 2.1.
+CNA maintains that when a region of the area is hit by signals caused by an object, it transiently activates a subset of all the neurons in the area.
+The majority of those neurons are not part of a co-activated net fragment and have to be silenced.
+This is achieved with the help of some periodically rising inhibition, which silences those neurons and leaves active the small subset of all transiently activated neurons that are supporting each other by excitatory connections.
+Due to neuron-wise overlap, net fragments merge into a coherent net, which can span the entire region covered by an object’s image and thus connect feature neurons that are spatially distant through chains of locally connected intermediary neurons, see Figure 1.
+It is this connectedness that unites neurons activated by an object in forming a structured net that separates those neurons from the background and from occluding elements of other objects, thus solving the segmentation problem (von der Malsburg, 2024) and tying them into a network amenable to be mapped to an invariant model network in another cortical area (see the outlook in Section 6).
+This study introduces the cooperative network architecture (CNA), representing the first concrete implementation of some fundamental components proposed in von der Malsburg et al.
+(2022).
+We focus on how to leverage networks as neural code and, specifically, discuss the
+  • establishment of a computational framework comparable to recent machine learning advances,
+  • learning of net fragments,
+  • demonstration of the activation of object-spanning networks (nets),
+  • demonstration of filter properties afforded by nets,
+  • discussion of nets in contrast to established ideas in neural computation and machine learning,
+while relating to the significance of nets for invariant object recognition in multi-area settings only as outlook in the discussion section (i.e.
+we focus here on a single area while object recognition based on multiple areas is proposed for future research).
+We demonstrate that net fragments can be learned without relying on a teaching signal, whether supervised or unsupervised.
+This neural code exhibits robust capabilities for handling out-of-distribution data and effectively represents objects composed of fundamental features encountered during training—a task at which autoencoders (Hinton and Salakhutdinov, 2006) fail.
+2 Relation to Alternative Concepts
+The model we describe below is the first to demonstrate in concrete form the generation of structured nets that form a neural code to represent visual structure.
+To grasp the significance of the approach, it is important to compare it to alternative approaches.
+2.1 Associative Memory
+Associative memory (AM), as formulated by Amari (1972) or Hopfield (1982), shares with the present study the aim of retrieving previously encountered structure, and in both, the structure is recorded with the help of Hebbian synaptic plasticity.
+Whereas AM deals with the storage and retrieval of neural firing patterns, our purpose is to activate networks, that is, connectivity patterns that are structured by previously recorded connectivity pattern fragments.
+Admittedly, in both cases, the (re)construction process results in the activation of subsets of all neurons of the area, but whereas in AM, the connectivity of the set of active neurons is all-to-all, in our case, connectivity is restricted to within the net fragments that have previously been recorded, which in our visual application are restricted to short-range (local) connections and express two-dimensional topological structure.
+It is this structured connectivity that serves as basis for the structure-sensitive matching mentioned in the introduction.
+The monolithic all-to-all connectivity of the AM, by distinction, does not afford such structure-sensitive matching (i.e.
+instead of representing each structure as monolithic whole, the CNA learns to decompose them into local sub-patterns (net fragments), which can then be dynamically composed into nets to represent structures).
+This functional difference is conditioned on a different mode of learning.
+Whereas in AM, a neural pattern is stored in a single event by strengthening the connections between any pair of currently active neurons, our approach distinguishes significant from insignificant connections, significant being those that are strengthened again and again, while insignificant are those that are strengthened only occasionally.
+This learning is driven by exposure to a large number of patterns, none of which may ever repeat, but which contain small repeating subpatterns.
+It is these that generate net fragments.
+In the visual case, it is local texture patches that play this role of repeating sub-patterns.
+A basic limitation of AM is that unless the patterns to be stored are statistically independent (orthogonal), that is, have small neuron-wise overlap, they interfere with each other by the spurious activation of neurons.
+The problem can be reduced (as proposed by Tsodyks and Feigel’man (2007) or Palm (2013)) if patterns are sparse, that is, activate only a small fraction of all neurons.
+The statistical mode of pattern storage proposed here opens a more fundamental solution to the problem of pattern overlap.
+We propose that each input feature (specified by feature type and position) is represented not by one neuron but by a number of ‘alternative neurons’, neurons that have the exact same receptive field (for examples, see Cossell et al.
+(2015)), and thus are activated by the same afferents  [^1] There is ample space for alternative neurons, as there are 100 times more neurons in primary visual cortex than output neurons in the retina (Leuba and Kraftsik, 1994)..
+While these neurons have the same afferent input, they can exhibit different excitatory output.
+Given this expansion, different net fragments with feature-wise overlap can reduce neuron-wise overlap by making use of different alternative neurons.
+AM and our approach differ in yet another aspect, the mode of retrieval (or, in our case, construction) of patterns.
+In the version of Hopfield (1982) of AM, this is realized by minimizing an energy function by sequential switching of neurons.
+In our approach, winning patterns are projected out from the set of all neurons activated by the input and by rising inhibition that silences all those alternative neurons that do not have sufficient recurrent excitatory support, no neuron-sequential process being necessary.
+2.2 Boltzmann Machine
+Some important aspects of the system illustrated by our model are highlighted best by comparing it with the Boltzmann Machine (BM) proposed in Hinton and Sejnowski (1983).
+Both systems share features such as the ability to admitting recurrent connections (as does Mumford (2003)), but they also exhibit significant differences.
+Nevertheless, as our focus is limited to one area and is not the full architecture as proposed in von der Malsburg et al.
+(2022), we highlight only select differences here.
+The BM makes a clear distinction between visible and hidden (‘latent’) neurons.
+In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by inhibition in favor of those neurons that support each other through previously learned connections.
+A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
+In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference that we do not insist on symmetric connections and therefore do not derive an energy function and consequently cannot prove convergence of the neural dynamics (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the missing of up to half of the connections due to lack of symmetry does not jeopardize stability of a state).
+In a single cycle of the periodically rising inhibition, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
+2.3 Artificial Neural Networks (ANN)
+In the concept pursued here, objects and other structures within the input are represented by nets, sets of neurons that are held together by cooperative excitatory connections and are activated as a whole, the activation of individual neurons being conditioned on the activation of other neurons in the same net.
+In contrast, a large variety of ANN versions and learning techniques have been proposed, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), autoencoders (Hinton and Salakhutdinov, 2006), or representation learning (Bengio et al., 2014).
+Common to all of these is that the sensory input is converted through many layers into a representation in the form of a set of neuron states.
+These neurons can be considered as propositions that indicate the presence of complex features (activity patterns) within the input.
+This representation of the input is then exploited by evaluation circuits (often in the form of a few fully connected layers of neurons) to achieve functions such as object classification, next-token prediction, or, in autoencoders, reconstruction of the input.
+The common feature of all ANN systems is that the connections from the input to the representation layer and from there to the output are purely feedforward and are trained by back-propagation of error (Rosenblatt, 1962, p.
+292).
+The CNA proposes a clear separation between input representation (within an area) and its functional evaluation (between areas, see Section 6).
+The representation of the input is trained by its statistics and is thus essentially neutral and without specialization to any specific function, giving it foundational character.
+In ANN systems, the selection of patterns to be extracted from the input is strongly biased to serve the trained function, which may be the reason that a transformer trained on one set of functions cannot continue to be trained on another (Dohare et al., 2024).
+ANN’s hierarchy of layers of neurons responding to features of intermediate complexity is replaced in the CNA by a single layer, in which net fragments of different sizes represent a hierarchy of features of varying complexity (and inter-mesh to form coherent nets standing for whole objects, thus solving the binding problem).
+The purpose of transition to other areas is, in the CNA, not transition to higher complexity but transition to representations that generalize over, e.g., position, size, or orientation of nets in lower areas, as discussed in Section 6.
+In ANN systems, the high-level feature neurons that are activated by an input do not express any relations between them.
+This ‘binding problem’ (von der Malsburg, 1981; Burwick, 2014; Greff et al., 2020)) can give rise to ambiguity, e.g., the hallucination of feature conjunctions (Treisman and Schmidt, 1982).
+This ambiguity can be reduced by additional neurons whose trigger patterns overlap those of other neurons whose relative position would otherwise be in doubt (autoencoders (Hinton and Salakhutdinov, 2006) show that ambiguity can be suppressed altogether by a sufficiently large number of hidden neurons).
+In the CNA, the binding problem is solved in a natural way by net fragments that interconnect sets of features that co-occur with statistical significance in the input and that merge together into nets representing coherent chunks within a given input.
+This simplicity of representation of visual structure in CNA is due to its fundamentally different perspective on neural representation, in that the active state is not only described by the elementary features that are active but also by the connections between them.
+This is made possible by different alternative neurons that represent the same elementary feature but express different connectivity.
 3 Computational Framework
-Figure 2: The dynamics of a single area within the proposed DNA: The input image is fed into stage $S1$ to obtain feature activations.
+Figure 2: The dynamics of a single area within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations.
 The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer.
-In the upper part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the lateral connections within $S2$.
-In the lower part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
-We here restrict ourselves to modeling this functionality within a single cortical area, although the significance of coherent nets in primary sensory areas can be fully appreciated only in the context of different interacting areas.
-A single area of a DNA model comprising net fragments is shown in Figure 2(A).
-Our system is based on an input layer representing observed images by simple feature detectors (which are perhaps to be localized within layer IV of the primary visual cortex, see, for instance, Usrey et al.
-(2003)) and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$) whose neurons are thought to be also localized in the same primary visual cortex area, presumably within layers II and III.
+In part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the recurrent connections within $S2$.
+In part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
+Our framework addresses the visual system, even though it could be applied to other modalities as well.
+In the primary visual cortex, images appear as two-dimensional arrays of activated neurons, and its neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004).
+To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (‘convolutional’ structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift.
+We use these insights to build a CNA with two stages, as shown in Figure 2(A).
+The CNA is based on an input layer representing observed images by simple feature detectors and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$).
 Feature Extraction.
 The system’s input is a retinal 2D image represented by variables $\boldsymbol{x}_{k,i}$ (gray box in Figure 2), where $k\in(1,...,C)$ denotes the channel index of the image ($C=1$ for grayscale, $C=3$ for RGB), and the composite index $i=(h,w)\in(H\times W)$ stands for vertical and horizontal position of pixels within the input image.
-Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their “membrane potential” $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2).
+Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their ‘membrane potential’ $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2).
 The membrane potential $\boldsymbol{a}^{(S1)}$ is calculated by applying a convolutional operation (LeCun et al., 1989) (yellow box $S1$ in Figure 2):
@@
 $\boldsymbol{y}_{c,j}^{(S1)}=\mathrm{B}(\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)})=\begin{cases}1,&\text{ if }\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)}>0\\ 0,&\text{otherwise}\end{cases}$ (2)
-with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (cells fire if the membrane potential is above this threshold).
+with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (neurons fire if the membrane potential is above this threshold).
 Dynamics.
-Stage $S2$ builds net fragments by allowing active neurons to support laterally connected neurons over T time steps (depicted as recurrent connection in Figure 2).
+Stage $S2$ builds net fragments by allowing active neurons to support recurrently connected neurons over T time steps (depicted as recurrent connection in Figure 2).
 A sample image is introduced into the system at time step $t=0$ to activate neurons and remains unchanged until time step $t=T$.
-Throughout this period, neuronal activity undergoes sparsification due to growing inhibition.
+Throughout this period, neuronal activity undergoes sparsification due to growing inhibition (see below).
 Through the sparsification process, the system forms an object representation by leaving active a set of neurons each of which is supported by active neighbors, that is, a net.
-Short-Range Lateral Connections.
-Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feed-forward connections between different layers, we employ them here to implement short-range lateral connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments.
-Given lateral connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$.
-In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state.
-The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by lateral excitation from within $S2$.
+Short-Range recurrent Connections.
+Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feedforward connections between different layers, we employ them here to implement short-range recurrent connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments.
+Given recurrent connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$.
+In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state to project out a net representing a given sensory signal.
+Importantly, the internal state evolves continuously, even when processing the same input pattern, driven by dynamic changes in the support between neurons along recurrent connections.
+The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by recurrent excitation from within $S2$.
 Before digitization, they are modified by saturation, normalization, and inhibition (yellow box $S2$ in Figure 2).
 Before those modifications, its raw form is computed as:
-$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Lateral Connection}}.$ (3)
+$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Recurrent Connection}}.$ (3)
 Similar to $S1$, $c\in C^{(S2)}$ denotes the feature channel of $S2$, the compound index $j$ expands to $j=\{w,h\}$, $w\in W$, $h\in H$, and the compound index $i$ is running in the range $l_{j}=(h,....,h+h^{(W2)};w,....,w+w^{(W2)})$, with $h^{(W2)}$ and $w^{(W2)}$ being the size of the convolution kernel.
-The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the lateral connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e.
-$C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments, see Apendix A).
-In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the lateral connections determines which neurons can be connected within stage 2.
-The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix B).
-The weights are updated using Hebbian Learning (details in Appendix C).
-For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide lateral support:
+The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the recurrent connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e.
+$C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments).
+In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the recurrent connections determines which neurons can be connected within stage 2.
+The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix A).
+The weights are updated using Hebbian Learning (details in Appendix B).
+For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide recurrent support:
 $\boldsymbol{y}_{i}^{(S2)}[t=0]=0$ (4)
@@
 This entire process is visually depicted in the upper part (A) of Figure 2.
-For simplicity in our algorithm, and since we use the same kernel size for the forward and lateral connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
+For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
 $\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(S2)}_{c,k,j-i}\cdot\left(\boldsymbol{y}_{k,i}^{(S1)};\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)$ (5)
@@
 These steps of processing the raw membrane potential $\boldsymbol{a}^{\prime}$ of neurons proved necessary in our simulation experiments to achieve the qualitative function we aimed for.
-The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix D.
+The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix C.
 The differences between activation levels of neurons are accentuated by setting
@@
 The coefficient $\gamma$ regulates the inhibition strength and increases with the time-steps of the inner loop according to $\gamma=1.2+0.2t$, making it harder for neurons to fire (as $\boldsymbol{a}^{\text{norm }(S2)}\leq 1$).
-The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive lateral support, while invalid activations are less supported.
+The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive recurrent support, while invalid activations are less supported.
 While the inhibition coefficient increases over time, many neurons eventually get suppressed, leaving only those active that support each other and form a consistent net.
@@
 $\boldsymbol{y}_{c,j}^{(S2)}[t]=\mathrm{B}\left(\boldsymbol{a}^{(S2)}_{c,j}[t]-b^{(S2)}\right)$ (7)
-The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the lateral connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed.
-To allow features to be part of multiple net fragments and represent various patterns, they are represented by a set of alternative neurons (details in Appendix A).
-All neurons within such sets are free to learn different lateral connections to other neurons.
-They are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero).
-This selection process process is described in the Appendix E.
+The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the recurrent connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed.
+Alternative neurons are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero).
+This selection process process is described in the Appendix D.
+Relation to Nets and Fragments.
+Net fragments are learned through the strengthening of recurrent connections $\boldsymbol{W}^{(L)}$, facilitated by Hebbian plasticity.
+Neurons within a fragment are interconnected, providing mutual support that enhances their membrane potentials.
+By shifting $\boldsymbol{W}^{(L)}$ as convolutional kernels across the input, fragments can be detected at all spatial locations, resulting in numerous overlapping fragments.
+Coherent fragments are interconnected to form a larger structure, a net.
+While multiple nets may exist, projecting out the most coherent one is crucial for robust object representation.
+This is achieved by progressively increasing inhibition over time.
+As inhibition intensifies, neurons with insufficient support (e.g., only a few instead of many neurons within one net fragment were triggered) are deactivated [^2] Neuron activations can be interpreted as hypotheses that require confirmation/support from other neurons with compatible hypotheses - otherwise, they are deactivated..
+This deactivation can propagate, leading to the elimination of entire fragments and leaving a single coherent net to represent the input structure.
+In our experiments (see below), we observe that, initially, neurons operate independently due to the absence of established recurrent connections, with a single neuron having a membrane potential of $1$ being sufficient for activation.
+However, after training, the average membrane potential rises to approximately $12$, reflecting robust support from neurons within a fragment.
+As a result, inconsistent neurons can no longer activate (and neurons with a membrane potential of $1$ are switched off), ensuring a more stable and coherent representation.
 4 Experiments
-We evaluate our framework by assessing the efficacy of net fragments to filter out random Gaussian noise, the capacity to reconstruct occluded objects, i.e., the efficacy of dealing with subtractive noise, and its generalization capacity by composing small net fragments into object structures not observed during training.
-5.1 Dataset
-The dataset comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
-The images are generated when required and their set may vary from one epoch to another.
-During each training cycle, we randomly generate $300$ image instances (each line about $5$ times).
-During evaluation, we sample each of the $59$ distinct lines exactly once.
-In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
-We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$.
-To that end, we probabilistically flip neurons within each channel with a probability of up to $20\%$.
-Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location.
-To evaluate subtractive noise, we create discontinuous lines by deleting a line segment in the middle.
-To demonstrate the generalization capability of DNA areas, we generate validation samples that differ from the ones used during training.
-Specifically, we produce kinked lines by generating four random points within the image and draw straight lines between these points.
+To demonstrate the compositionality of net fragments, we train (updating $\boldsymbol{W}^{(L)}$ with Hebbian plasticity) the model on binary images containing straight-line patterns.
+For evaluation (after training and without updating any connections), we use images of kinked lines, digits, characters, and line drawings of objects such as houses and clocks.
+These complex patterns can be expressed as compositions of straight-line net fragments, even though the CNA was not explicitly trained on such patterns.
+This flexible compositional behavior contrasts strongly with associative memory (see Section 2.1).
+To highlight the differences between ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
+Besides compositionality, we introduce noise to examine the robustness and persistence of coherent structures with statistical relevance.
+Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects, i.e., the efficacy of dealing with subtractive noise.
+We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
+Further experimental details are provided in the Appendix: We describe the training and testing datasets in more detail in Appendix E, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix F, the training details of our CNA model are in Appendix G, and the training details of the autoencoder in Appendix H.
 5 Results
-In the following, we show that the proposed net fragments can suppress Gaussian noise in Section 6.1 and can reconstruct removed pixels, see Section 6.2.
-A description of the feature extraction mechanism to obtain $\boldsymbol{y}^{(S1)}$ is in Appendix F, and details about the used parameters are in Appendix G.
-Additionally, Appendix H contains visualizations of the average support strength received by cells, and Appendix I a quantitative evaluation of suppressed noise.
+Figure 3: Net formed by activated neurons.
+The left of the image shows the neuronal activation $\boldsymbol{y}^{(S2)}$ when observing a line, and the right visualizes the net formed by these neurons.
+For convenience, the neurons are rearranged in a circle, and the connections of two randomly selected neurons, A and B, are visualized in colors (purple and red).
+We empathize that this work focuses on learning a new neural code and, in the following, demonstrates it on a rather simple dataset.
+Although our result figures display activated neurons, we would like to emphasize that these neurons are carriers of connections and form part of nets.
+To make this point for once, Figure 3 depicts this connectivity of activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged as a circle for convenience).
+5.1 Compositionality
+Figure 4: The top of the figure illustrates the straight lines used during the training phase.
+Below, various samples from the evaluation datasets are shown, including kinked lines (rows $1$ and $2$), digits and characters (rows 3 and 4), and line drawings (rows $5$-$8$).
+For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
+The model weights, $\boldsymbol{W}^{(S2)}$, are trained using straight lines (see training data at the top of Figure 4), allowing the model to capture the underlying structure of straight lines within net fragments.
+The model composes these fragments dynamically into nets to represent much more complex structures (see validation data in rows 1-8 of Figure 4).
+For instance, despite not encountering these images during training, the CNA model can represent kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows $5$-$8$).
+The motivation for why we build net fragments ($\boldsymbol{y}^{(S2)}$) and not just use the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise.
+This noise corresponds to invalid patterns not observed during training, which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons.
+Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
 5.2 Filtering Gaussian Noise
-Figure 3: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by lateral connections of net fragments.
+Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by recurrent connections of net fragments.
 In each image, the background is shown in black, and the four different features in distinct colors.
@@
 The illustration in Figure 5 visually demonstrates the system’s effective capability of filtering out isolated noise or small clusters thereof.
-Noise of varying degrees (from $0\%$ to $59\%$ noise probability at each spatial location - see Appendix 5.1) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments.
+Noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix E) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments.
 While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated.
@@
 The plot is created after adding $10\%$ noise to each feature channel (sing a bias $b^{(S2)}=0.7$ and an inhibition coefficient $\gamma=1.2+0.2t$).
-Figure 3 visualizes that neural activity simulating visual clutter without coherent structure is removed after processing for $5$ time-steps.
 Figure 6 depicts this systematic filtering over time.
-The additionally introduced clutter doesn’t activate overlapping net fragments and cannot withstand increasing inhibition.
+The additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time.
 The figure illustrates a substantial noise reduction already at time step $t=0$ due to the filtering by matrix $\boldsymbol{W}^{(F)}$ and initial inhibition.
 In each subsequent time step, additional clutter is removed.
-This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons.
-We thus observe a chain reaction of gradual suppression of incoherent structure as well as the preservation of net structure on the basis of internal support.
-6.2 Reconstructing Subtractive Noise
-Figure 5: Partial occlusion as subtractive noise: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations.
+This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons, demonstrating how overlapping fragments build a coherent net of mutually supporting neurons.
+5.3 Reconstructing Figures
+Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations.
 Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
-If a substantial number of neurons activate and support other neurons within a net fragment, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction.
-Figure 5 shows how inactive neurons activate and thus demonstrates that net fragments can deal with subtractive noise (occluded patterns).
+Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the ‘activation probabilities’ (on the right) after the first time step.
+If a substantial number of neurons activate and support other neurons within a net, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction.
+Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns.
 This reconstruction works only reliably for up to $3$ missing pixels (see Appendix I).
@@
 Membrane Potential Represents Expected Feature Activation.
-Figure 6: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the “activation probabilities” (on the right) after the first time step.
-The activation probability is close to $1$ where a line has been observed and around $50\%$ in the middle where no line was observed but expected.
-Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
-When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ plays a crucial role.
-The potential is higher at spatial locations where line features are observed and lower where they are not observed but expected.
-Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 6.
-Modeling this ambiguity enables having different internal interpretations, such as whether “it is an interrupted line that should be reconstructed” or “there are two co-linear line segments.” We find that the activation map is especially reliable during the first time step, where it is calculated based on the observed input and the network’s expectation about local patterns.
-Since local interactions can trigger or deactivate neurons, the activation map changes over time, reducing its interpretability.
+When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected.
+Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 7B.
+Modeling this ambiguity enables having different internal interpretations, such as whether ‘it is an interrupted line that should be reconstructed’ or ‘there are two co-linear line segments’ and selecting one of them based on the threshold $b^{(S2)}$.
 Increasing Inhibition Over Time Stabilizes Patterns.
@@
 Initially, at time step $t=0$, the net fragments represent the interrupted line.
-Subsequently, during the second time step ($t=1$), some of the removed pixels are reconstructed, and shortly after, at $t=2$, the entire line is represented by the net fragments.
+Subsequently, during the second time step ($t=1$), one of the removed pixels is reconstructed, and shortly after, at $t=2$, the entire line is represented.
 Thus, reconstruction can persist even amid increasing inhibition, given that inactive neurons receive adequate support.
-This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of certain neurons encourages others to follow.
-6.3 Compositionality
-A single DNA area models a feature hierarchy through nested net fragments of increasing size (von der Malsburg et al., 2022b).
-In this study, we demonstrate that these higher-level nets can be composed of local features (small net fragments), even though the underlying patterns had never been observed during training.
-We show this on the example of curved lines, which share only local features with the straight lines of the training set.
-Figure 8: Input images of curved lines are shown on the left, the second and third columns show the corresponding feature and output activation without noise, and the fourth and fifth columns show the same feature and output activation with noise corruption (activation bias $b^{(S2)}=0.3$, inhibition coefficient $\gamma=0.6+0.1t$).
-Figure 8 depicts four samples of curved lines with the corresponding neuronal activation within an area, once without noise and once with $18.5\%$ noise added to its afferent input.
-The cortical area is able to represent the curved lines as composite of net fragments learned from straight lines.
-Added unstructured clutter is efficiently removed, leading to robust input representation.
-7 Discussion and Conclusions
-The concepts proposed here may be compared to established neural theory.
-Nets are based on associative learning but differ fundamentally from Hebbian assemblies (see Appendix A) as described in (Hopfield, 1982; Cohen and Grossberg, 1983; Papadimitriou et al., 2020), which are monolithic and without inner structure, whereas nets can represent an infinity of patterns based on shared fragments and can deal with new entities based on homeomorphy between common compositional structure, thus enabling the systematicity and productivity characterizing cognitive operations (Fodor and Pylyshyn, 1988; von der Malsburg, 2023).
-In comparison to multi-layered perceptrons, coherent nets replace hierarchies of independent feature neurons by net fragments, share with them the potential for equivariance (shift symmetry), but in distinction to them support the construction of an invariant representation that doesn’t need to be learned object-by-object (see next section).
-With the transformer architecture (Vaswani et al., 2017), nets share the provision for finding non-local patterns through on-the-fly computation of lateral links.
-A decisive difference between DNA and all deep-learning systems (LeCun et al., 2015; Schmidhuber, 2015; Stadelmann et al., 2019) is the mode of learning.
-ANNs typically rely on backpropagation of error where the goal is reaching consistency (minimal error) between system output and teaching signals.
-With DNA models, the goal is mutual consistency at each neuron, between the excitatory connections converging there, this consistency being measured by the probability for each connection to reach a highly active neuron when carrying a signal itself.
-Of all possible connectivity patterns, only a minute fraction has this property of self-consistency.
-This condition of self-consistency acts as a selection criterion for eligible networks and thus powerfully reduces the search space of learning.
-Net fragments and the structured nets they build form a basis for the representation of hierarchical structures.
-They are effective for feature filtering and for preventing early commitment, which may be of fundamental importance for improving the robustness of current machine learning systems.
-So far, a neural code based on net fragments has merely been proposed as a theoretical concept (von der Malsburg et al., 2022b), while this study pioneers the implementation of net fragments in a concrete simulation.
-Our experiments demonstrate that the proposed DNA architecture builds net fragments that effectively stabilize patterns, enhance robustness by eliminating visual clutter, and can reconstruct removed parts of known patterns.
-The DNA model can handle additive noise exceptionally well and eliminates up to $96\%$ of it while preserving the initial pattern, even in scenarios where up to $59\%$ of all locations are affected.
-Figure completion works well for line interrupts for segments of up to $3$ pixels, while for larger interruptions, the reconstruction rate decreases considerably.
-One reason for this is certainly due to the nature of the used 1D pattern: as it involves straight lines with a width of one pixel, typically, only a fraction of the neurons within the local 2D neighborhood of size $h^{(W2)}\times w^{(W2)}$ are active.
-With a larger line interrupt, the proportion of active neurons compared to neighboring neurons is vanishingly small, which makes reconstruction more difficult.
-It is expected that reconstruction will be much more robust with two-dimensional patterns.
-We posit that this work, which focuses on the internals of a single cortical area, allows for scaling to multiple areas that will allow the implementation of more complex tasks (see next section).
-Nevertheless, although this work is intended to be a step towards implementing a robust learning framework and the proposed neural code exhibits promising characteristics, it still has weaknesses that should be addressed in future work.
-One such weakness is dependence on manual adjustment of the parameters $b$ and $\gamma$.
-The mechanism of signal control is biologically implausible.
-It should be replaced by a more elegant, theoretically justified, and self-adjusting mechanism.
-8 Future Work
-One important purpose of handing structure from one area to another is generalization, a prime example for which in vision is object recognition invariant to translation (and wider transformation groups).
-While our system only includes one “cortical” area, another area (that would correspond to the cortical inferotemporal cortex (Ito et al., 1995)) could contain invariant object models.
-In the context of DNA, we see invariant recognition as realized by a process in which an object-representing net in a primary sensory area $A$ forms and activates a homeomorphic net in a secondary area $B$.
-Two nets are homeomorphic if they contain the same feature types in the same spatial arrangement.
-Invariance is achieved if differently transformed mutually homeomorphic nets in $A$ activate the same net in $B$ under a homeomorphic mapping.
-Inter-areal homeomorphic maps are based on feature type-preserving connections, which in the brain are realized by axonal fibers [^2] With the help of several intermediate areas, the number of necessary connections can be brought down to realistic values (Anderson and van Essen, 1987; Olshausen et al., 1993; Wolfrum et al., 2008))..
-In fact, connections should not only be feature type-preserving but also restricted to pairs of corresponding alternative neurons in $A$ and $B$, such that they carry not only feature information but also information on connectedness [^3] To further ease the formation of homeomorphic maps in $B$, a mechanism for mutual support between mapping-compatible connections has been proposed and successfully applied (see Section 2.3)..
-The generation of the net in $B$ is again the result of a two-phase process, the afferent connections from $A$ to $B$ transiently activating many neurons in $B$, most of which getting silenced in phase two, leaving only those neurons active that not only got afferent input but are also supported by lateral excitation within $B$.
-The net thus formed in $B$ is invariant to homeomorphic transformations in $A$.
-Object classification, the identification of the net in $B$ as a particular object type or recognition of individual objects needs a further area $C$ containing a fixed net that is activated based on approximate homeomorphy with the invariant net in $B$.
-The result of the recognition process is the formation of a larger net composed of a net in $A$, a net in $B$, and one-to-one connections between corresponding neurons in $A$ and $B$.
-The self-consistency of this larger net is the basis for learning all participating connections, as all connections are stabilized (or generated) by their success in predicting activity in their target neuron.
-Through inverse mapping, these stabilizing interactions can be re-imported to the previous area and thus help to discriminate the corresponding nets there from noise and clutter and bridge larger partial occlusions, which would realize the idea of predictive coding Keller and Mrsic-Flogel (2018).
-An obvious first step in preparation for a larger-scale system of interacting areas is, however, the training of a single primary-sensory area with natural images to capture as net fragments all textures that occur in them with significant frequency.
-This project should include the extraction of such different sensory qualities as stereo depth, local motion or surface discontinuities.
-Net fragments and nets based on these would then have a chance to relate to intrinsic surface properties such as form.
+This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of neurons encourages others to follow.
+5.4 Comparison to an Autoencoder
+Figure 9: Comparison of the outputs generated by the proposed CNA model and an autoencoder across different tasks.
+Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder).
+Rows $1$ to $3$ illustrate how both models encode patterns that were not present in the training dataset.
+Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
+Figure 9 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix H).
+While the autoencoder successfully learns to accurately represent line structures in the training data - achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and decoder reconstruction - it struggles with generalization to data not observed during training.
+For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
+Rows $4$ and $5$ highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced.
+However, as the noise level increases, its performance decreases significantly.
+The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations.
+Metrics on noise reduction, precision, and recall, presented in Appendix I, further substantiate the CNA model’s superior performance.
+In row $6$, the models’ capabilities in reconstructing missing line segments are compared.
+Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed.
+This is confirmed by the metrics in Appendix I: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of more noise at other locations, leading to lower overall precision and recall.
+6 Discussion and Outlook
+The CNA presented here establishes a novel approach to construct coherent nets as neural representations of visual input, enabling arrays of elements to be treated as unified gestalts through the composability of previously learned sub-patterns (net fragments).
+While the concept of such net construction has been proposed theoretically (von der Malsburg et al., 2022), this work marks the first successful implementation and simulation of the idea.
+Our approach offers distinct advantages over existing neural coding frameworks, particularly in terms of robustness and the ability to represent out-of-domain data.
+Using nets of neurons to represent sensory signals allows filtering out noise and figure completion, while using net fragments allows representing unseen objects through composing small feature patches.
+In contrast to past work that relied on manually constructed filter functions (e.g., Geman and Geman, 1984), the filter functions, nets in our scheme, are learned from input statistics resp.
+constructed on the spot, so that they will eventually be able to cover the whole gamut of patterns encountered in natural images.
+An essential next step would be the construction of a system that is trained on natural images, that contains at each position of the visual field a complement of elementary feature detector neurons able to accurately represent the pixels of the input (as in Olshausen and Field, 1996), and whose recurrent connections have sufficient lateral extent to capture the most important signal correlations present in natural images.
+Such systems can then be expected to develop net fragments that capture and represent all textures encountered in natural images on that length scale, thus realizing the ambition expressed in Figure 1.
+A first and important such operation is abstracting over position, size, and in-plane orientation of an object’s image in the input area (primary visual cortex) by the construction of an invariant net representation of the object in a separate area (which would correspond to inferotemporal cortex (Ito et al., 1995)).
+As hinted at in the Introduction, this can be based on structure-sensitive, ‘homeomorphic’, mappings, as proposed in von der Malsburg (1981); Anderson and Van Essen (1987); Olshausen et al.
+(1995); Wiskott et al.
+(1997); Wolfrum et al.
+(2008); Memisevic and Hinton (2010) or Arathorn et al.
+(2013).
+A mapping is homeomorphic if it preserves neural feature type as well as connectivity.
+Object classification or recognition of individual objects can then be based on a further homeomorphic mapping to an area containing a more permanent schematic representation of the object class or a sample of the object to be recognized, as realized in Wolfrum et al.
+(2008).
+The introduction of nets as a neural code and the inclusion of alternative neurons suggest a natural basis for the mapping process by requiring inter-areal connections to be not only specific to feature type but also to alternative neuron identity.
+As this identity contains information on net connectivity, this specificity can implement the homeomorphy constraint of inter-areal mappings in a natural way and overcome the limitations in existing systems (Wolfrum et al., 2008).
+Once the mechanism of homeomorphic mapping has been established, it is available for arbitrary novel objects, thus forming the basis for invariant object recognition at first sight (as observed in humans in Biederman and Cooper (1991)).
+This capability stands in stark contrast to current ANN technology, which requires invariance to be trained individually for each object, highlighting the broader generalization potential of our approach.
+The simple mechanism of net formation that we model here is a first but significant step in the direction of a data architecture grounded in nets connecting sparse sets of neurons within and between areas.
+Nets are singled out by the properties of cooperativity and inner consistency: all neurons of an active net experience the high-order coincidence of converging excitatory signals, realizing predictive coding at each neuron.
+Given high sparsity of neural firing and high sparsity of connectivity, satisfying the requirement of consistency is highly non-trivial and can only be established by a learning process in which single neurons acquire converging connections coming from sets of neurons that fire simultaneously again and again, as realized in our model.
+In summary, the CNA demonstrates the feasibility and potential of a principled and compositional approach to neural coding.
+Its ability to autonomously construct robust representations, handle out-of-domain data, and generalize to novel stimuli underscores its promise as a foundation for more advanced systems in the future, especially when extended with multiple areas.
 Funding
-This research has been funded by the Canton of Zurich, Switzerland, through the Digitalization Initiative of the Canton of Zurich (DIZH) Fellowship project “Stability of self-organizing net fragments as inductive bias for next-generation deep learning.”
+This research has been funded by the Canton of Zurich, Switzerland, through the Digitalization Initiative of the Canton of Zurich (DIZH) Fellowship project ‘Stability of self-organizing net fragments as inductive bias for next-generation deep learning.’
 Data Availability Statement
-The code to generate the data as well as to reproduce the results can be found after the publication of this work on Github at https://github.com/sagerpascal/dynamic-link-architecture.
+The code to generate the data as well as to reproduce the results can be found after the publication of this work on Github at https://github.com/sagerpascal/cna.
 Appendix
-Appendix A Net Fragments vs. Associative Memory, Alternative Neurons
-Learning of lateral connections to form net fragments shares aspects with associative memory (Hopfield, 1982; Cohen and Grossberg, 1983).
-There are, however, important differences.
-In an associative memory system, neural patterns are stored at once in their entirety with the help of synaptic plasticity, any pair of active neurons within a pattern strengthening their mutual excitatory connections.
-The resulting connectivity turns each stored pattern into an attractor state that can be recalled (dynamically activated) starting from initial states that come near to it.
-Such stored and dynamically stabilized neural patterns are often called “assemblies” (Hebb, 1949), and are discussed to this day as fundamental to the function of the brain, see for instance (Papadimitriou et al., 2020).
-There are two problems with the concept of associative memory.
-One is rather fundamental in that a pattern is monolithic and can be recalled only as a rigid whole.
-This is due to the indiscriminate stickiness with which any pair of simultaneously active neurons is tied together by synaptic plasticity in the one-time storage process.
-Fragments, in distinction, are formed in a protracted statistical process in which connections are gradually established between neurons that are co-active with statistical significance.
-Due to the spatial deformation of visual patterns, this can only be the case for neurons relating to neighboring visual points.
-DNA models distinguish insignificant from significant connections, only the latter being permitted to grow, which they do under competition until reaching equilibrium.
-As a result, the system doesn’t store rigid global patterns but is able to dynamically stabilize an unlimited number of global states that are composed combinatorially out of overlapping fragments.
-The other, more technical, problem of associative memory is that stored states have to be statistically independent (“orthogonal”) to avoid cross-talk between them.
-(In a simple example of cross-talk, a neuron $c_{1}$ gets connected at different times with neurons $c_{2},c_{3},c_{4}$, although the pattern $\{c_{1},c_{2},c_{3},c_{4}\}$ doesn’t occur with statistical significance during learning.
-As a result, $c_{1}$ may be activated erroneously if $\{c_{2},c_{3},c_{4}\}$ is active as part of a larger pattern.) Sparsity (according to which a very small fraction of all neurons are active within any given pattern) has been proposed as a general means to avoid cross-talk and to ensure statistical independence of stored patterns (French, 1999).
-DNA models are freed from this constraint by the introduction of “alternative neurons,” neurons that are activated by the same input feature but are free to learn different lateral connections to other neurons.
-(In our example, neuron $c_{1}$ would be replaced by a set $\{c_{1},c_{1^{\prime}},c_{1^{\prime\prime}}\}$ of alternative neurons, which could then independently learn the connections $c_{1}-c_{2}$, $c_{1^{\prime}}-c_{3}$ and $c_{1^{\prime\prime}}-c_{4}$, thus avoiding the above cross-talk.) In the biological case, there is ample room for alternative neurons.
-There are, for instance, a hundred times more neurons in the primary visual cortex compared to the number of fibers coming out of the retina (Leuba and Kraftsik, 1994).
-Appendix B Weight Initialization
+Appendix A Weight Initialization
 The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are copied into the $S2$ neurons in the same position.
-This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative cells.
-The lateral connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$.
+This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative neurons.
+The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$.
 When Hebbian plasticity is employed on the weight matrix $\boldsymbol{W}^{(L)}$, these connections ought to grow (as neurons tend to keep their activation state between time steps), but due to the constraint of keeping all connection weights within the range $(0,1)$ (refer to equation 8), they stay put at $1$.
-Appendix C Hebbian Plasticity
+Appendix B Hebbian Plasticity
 After processing for $T$ time steps both matrices $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ undergo Hebbian updates (Hebb, 1949):
 $\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\alpha\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (8)
-where $\alpha$ represents the learning rate, and $\boldsymbol{\rho}_{avg}$ denotes the average correlation between input and output.
-This average is determined for a given activation over all instances of a given pair of neurons $c,c^{\prime}$ in different positions of the image domain.
-To compute $\boldsymbol{\rho}_{avg}$, we count the number of instances in which both neurons are on, subtract from it the number of instances where one neuron is on, and the other is off, and divide that sum by the number of all positive and negative instances.
-The actual modification eq.
-(8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
-The network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant.
-Consequently, the same lateral connection is applied across multiple output and input neurons.
+where $\alpha$ represents the learning rate.
+Note that the network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant.
+Consequently, the same recurrent connection is applied across multiple output and input neurons.
 The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects.
@@
 Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced by $\alpha$.
-Appendix D Normalization
-We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$. These steps are functionally motivated and not by reference to biology.
+The actual modification eq.
+(8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
+Appendix C Normalization
+We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$.
 We first apply a size limit:
-$\boldsymbol{a}^{\prime\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
-According to this formula, $\boldsymbol{a}^{\prime\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$.
+$\boldsymbol{a}^{\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
+According to this formula, $\boldsymbol{a}^{\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$.
 (In our experiments we found the value $\lambda=1.3\cdot\frac{h^{(W2)}+w^{(W2)}}{2}$ to work well).
-This saturation helps to mitigate imbalances between network fragments with differing numbers of participating neurons (as with our stimuli, line segments, where many more feature neurons activate around the end-points than in the middle of segments).
+This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons.
 In the next step, we introduce the normalization
-$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
+$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
 This normalization confines activations to the interval $(0,1)$, where neurons receiving maximal support are mapped to $1$, while those with lesser support are mapped to correspondingly smaller values.
-This normalization relative to the maximal value that itself is growing during learning instead of relative to a constant parameter like $\lambda$ is important to give neurons initially (when lateral connections are still small) a chance to fire and later, when lateral links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
-Appendix E Alternative Neuron Selection
+This normalization relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later, when recurrent links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
+Appendix D Alternative Neuron Selection
 Only one of a set of alternative neurons is allowed to be active.
@@
 Consequently, multiple features may be active at each location, but only one of the competing feature channels — the one best matching a given input patch — achieves activation.
-At initialization (see Appendix B), where all alternative channels fit equally well to a given local pattern, the first channel is selected.
+At initialization (see Appendix A), where all alternative channels fit equally well to a given local pattern, the first channel is selected.
 However, after applying a Hebbian update, this channel starts specializing in recognizing a specific local pattern.
 Consequently, the activation of subsequent local patterns may either trigger the same channel if they have a high correlation or activate another channel better suited to recognizing the new pattern.
+Appendix E Dataset
+The dataset comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
+The images are generated when required and may vary from one epoch to another.
+During each training cycle, we randomly generate $300$ image instances (each line about $5$ times).
+During evaluation, we sample each of the $59$ distinct lines exactly once.
+In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
+To introduce additive Gaussian noise to the afferent input in stage $S2$, we probabilistically flip neurons within each channel with a probability of up to $20\%$.
+Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location.
+To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
+To demonstrate the generalization capability of CNA areas, we generate validation samples that differ from the ones used during training.
+Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
+Additionally, we generate binary images measuring ($64\times 64$) pixels by hand that represent more complex structures.
+These images include line drawings of digits 00-$9$, characters, and objects such as a house, a star, a wheel, a clock, etc.
+Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
 Appendix F Feature Extraction
@@
 This approach is motivated by the observation that, given our specific elementary feature types (cf.
-Fig.
-9), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons).
+Figure 10), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons).
 The use of such manually designed filters allows a more straightforward interpretation of the results.
 While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
-These filters move across the image with a step size (“stride”) of $1$ (assuming 00 as input signal beyond the image border: “zero padding”) so that the input image and output image are of the same size.
+These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
 Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$.
@@
 We train the model with a learning rate of $\alpha=0.2$ for $100$ training cycles (epochs).
-The model uses $n_{a}=10$ alternative neurons with lateral connections spanning a kernel of $11\times 1$ 11 neurons.
+The model uses $n_{a}=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons.
 Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix F), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot n_{a}=40$.
 The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
-Appendix H Learned Support Strength
-Figure 10: The average support active and inactive neurons receive during training, captured by $\boldsymbol{a}^{\prime}_{j}$.
-The support strength is shown on the y-axis, and the training progress (epoch) is on the x-axis.
-A crucial parameter for gauging the efficacy of net fragments involves quantifying the difference in support received by active and inactive neurons.
-During the initial stages of training, active neurons receive a support of $1$, and inactive neurons receive a support of 00.
-The support increases during training as net fragments are formed.
-Figure 10 visually represents the average support for active and inactive neurons, captured by $\boldsymbol{a}^{\prime}_{j}$ (see eqs.
-3, 5).
-Notably, the graph demonstrates a widening gap between the support for active and inactive neurons.
-This discrepancy suggests an enhancement in the robustness of statistically relevant patterns over time as active neurons receive increased support and the support of inactive neurons remains close to 00.
+Appendix H Autoencoder Parameter
+We compare our CNA model to an ANN, specifically an autoencoder (Hinton and Salakhutdinov, 2006), which is a widely used self-supervised technique for generating input representations.
+An autoencoder is particularly suited for comparison because it includes a built-in decoder, allowing for feature reconstruction and visualization, which is required for feature comparison.
+To ensure a fair comparison, we also employ convolutional filters (LeCun et al., 1989) to the autoencoder and incorporate state-of-the-art design principles (we even permit more data and feature channels to the autoencoder).
+The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model.
+It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ layers.
+The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively.
+ReLU activations are applied after each layer.
+Consequently, the autoencoder is a relatively large model with $777$k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
+The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of $(0.9,0.999)$, and a weight decay of $1\times 10^{-8}$.
+To provide a robust comparison, we increase the number of training samples for the autoencoder.
+Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for $200$ epochs, using $10,000$ images per epoch.
+This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction.
+The autoencoder is trained with a batch size of $512$.
 Appendix I Quantifying Noise Filtration
@@
 The colors within each plot represent different activation biases $b^{(S2)}$.
+The dotted red line is the autoencoder, serving as baseline.
 Figure 11 displays noise reduction rates, recall, and precision across different parameter configurations for different levels of added noise.
@@
 […]
-Rconstructing Subtractive Noise.
+Furthermore, all CNA models (except the one with the lowest power factor and bias) outperform the autoencoder, demonstrating that they achieve performance that is competitive with state-of-the-art deep learning models.
+Reconstructing Subtractive Noise.
 Figure 12: Reconstruction of partial occlusion.
@@
 Columns are distinguished by the inhibition coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$.
+The dotted red line is the autoencoder model.
 Figure 12 displays the feature reconstruction rate, precision, and recall for removed line fragments of different lengths.
 Having a lower activation bias or inhibition coefficient $\gamma$ leads to a better reconstruction of removed patterns, as these settings allow neurons with less support to activate.
+Compared to the autoencoder baseline, the CNA model achieves higher recall and precision but a lower feature reconstruction rate.
+This indicates that the autoencoder is superior at reconstructing the pixels that have been removed but that, in general, much more noise (other neurons are activated/deactivated) is introduced.
 Interpreting Metrics.
@@
 In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$.
-Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines.
-Hence, we posit that various of the tested parameter configurations are well suited as they permit this distinction.
+Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines, which is the case for the CNA but not for the autoencoder (e.g., see Figure 11).

### v3

@@
 Abstract
-Nets, cooperative networks of neurons, have been proposed as format for the representation of sensory signals, as physical implementation of the Gestalt phenomenon and as solution to the neural binding problem, while the direct interaction between nets by structure-sensitive matching has been proposed as basis for object-global operations such as object detection.
-The nets are flexibly composed of overlapping net fragments, which are learned from statistical regularities of sensory input.
-We here present the cooperative network architecture (CNA), a concrete model that learns such net structure to represent input patterns and deals robustly with noise, deformation, and out-of-distribution data, thus laying the groundwork for a novel neural architecture.
-Keywords net fragments  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ cooperativity  $\cdot$ neuroscience  $\cdot$ machine learning
+We introduce the Cooperative Network Architecture (CNA), a model that represents sensory signals using structured, recurrently connected networks of neurons, termed “nets.” Nets are dynamically assembled from overlapping net fragments, which are learned based on statistical regularities in sensory input.
+This architecture offers robustness to noise, deformation, and out-of-distribution data, addressing challenges in current vision systems from a novel perspective.
+We demonstrate that net fragments can be learned without supervision and flexibly recombined to encode novel patterns, enabling figure completion and resilience to noise.
+Our findings establish CNA as a promising paradigm for developing neural representations that integrate local feature processing with global structure formation, providing a foundation for future research on invariant object recognition.
+Keywords net fragments  $\cdot$ nets  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ machine learning
 1 Introduction
 Figure 1: Representation of Objects by Nets.
-Objects are captured in their entirety by coherent nets that are composed of overlapping net fragments (schematically symbolized by white circles in the right panel).
+Objects are captured in their entirety by coherent nets composed of overlapping net fragments (schematically symbolized by white circles in the right panel).
 After training on natural images, each patch of visual space contains a complement of net fragments that represent textures that have been encountered with statistical significance.
 Net fragments overlap neuron-wise, and those activated by an object coalesce into a coherent net.
-The conundrum (stated in Olshausen and Field (2005)) that object contours can often not be found with the help of edge detectors due to lack of gray-level contrast (as in places inside the square in the left panel) may find its solution in that contours are defined as borders of the nets covering objects (or covering the background).
+The conundrum (stated in Olshausen and Field (2005)) that object contours can often not be found with the help of edge detectors due to lack of gray-level contrast (as in places inside the square in the left panel) may be resolved by the idea that contours are defined as borders of the nets covering objects (or covering the background).
 Figure adapted from Olshausen and Field (2005), with permission.
-Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Simmler et al., 2021).
-While many recent AI systems predominantly favor neural representations over symbolic ones, we follow a proposal by von der Malsburg et al.
-(2022) and present the cooperative network architecture (CNA), an architecture that uses structured networks of mutually (‘recurrently’) connected neurons as representation of sensory signals.
-Using structured networks to encode information brings a diverse set of advantages, including high robustness, ability of figure reconstruction, and representing of out-of-domain data, laying the foundations to build novel object recognition frameworks (see Section 6).
-The networks in the CNA are called nets and are dynamically assembled from overlapping net fragments, where net fragments are connectivity patterns that are learned by synaptic plasticity based on the statistics of sensory input signals.
-A major argument in favor of this data architecture is the possibility (as described repeatedly by Hinton (1981); von der Malsburg (1981); Anderson and Van Essen (1987); Kree and Zippelius (1988); Olshausen et al.
-(1995); Wiskott et al.
-(1997); Wolfrum et al.
-(2008); Memisevic and Hinton (2010); Arathorn et al.
-(2013)) of solving the invariant object recognition problem by structure-preserving mapping between nets representing variant images and invariant models of objects.
-The representation of image structure by recurrently connected nets formed between elementary feature detector neurons is a fundamental deviation from the current artificial neural networks (ANNs) paradigm, according to which image structure is represented by sets of complex-feature neurons activated by purely feedforward connections (see Section 2.3 for a comparison of the two approaches).
-In the cited models of invariant object recognition, the formation of object-representing networks within the image domain (the primary cortical area where sensory signals arrive – here called ‘area’) has been sidestepped by letting spatial neighborhood stand in for connectedness.
-Neighborhood, however, is too weak a condition to qualify neurons as having responded to the same object (which is, for instance, apparent when looking at an object through foliage).
-Following von der Malsburg et al.
-(2022), we rather assume that for sets of neurons to be admitted into a net, they have to have been co-activated frequently in the past by structured visual input and consequently have been tied together by Hebbian synaptic plasticity (Hebb, 1949) into net fragments.
-Net fragments play a role somewhat related to the code-book vectors of some image compression schemes (Taubman and Marcellin, 2002) in representing statistically dominant texture patterns within any patch of the visual field; the formation of net fragments is somewhat related to, though fundamentally different from, associative memory (Hopfield, 1982), see Section 2.1.
-CNA maintains that when a region of the area is hit by signals caused by an object, it transiently activates a subset of all the neurons in the area.
-The majority of those neurons are not part of a co-activated net fragment and have to be silenced.
-This is achieved with the help of some periodically rising inhibition, which silences those neurons and leaves active the small subset of all transiently activated neurons that are supporting each other by excitatory connections.
-Due to neuron-wise overlap, net fragments merge into a coherent net, which can span the entire region covered by an object’s image and thus connect feature neurons that are spatially distant through chains of locally connected intermediary neurons, see Figure 1.
-It is this connectedness that unites neurons activated by an object in forming a structured net that separates those neurons from the background and from occluding elements of other objects, thus solving the segmentation problem (von der Malsburg, 2024) and tying them into a network amenable to be mapped to an invariant model network in another cortical area (see the outlook in Section 6).
-This study introduces the cooperative network architecture (CNA), representing the first concrete implementation of some fundamental components proposed in von der Malsburg et al.
-(2022).
-We focus on how to leverage networks as neural code and, specifically, discuss the
-  • establishment of a computational framework comparable to recent machine learning advances,
-  • learning of net fragments,
-  • demonstration of the activation of object-spanning networks (nets),
-  • demonstration of filter properties afforded by nets,
-  • discussion of nets in contrast to established ideas in neural computation and machine learning,
-while relating to the significance of nets for invariant object recognition in multi-area settings only as outlook in the discussion section (i.e.
-we focus here on a single area while object recognition based on multiple areas is proposed for future research).
-We demonstrate that net fragments can be learned without relying on a teaching signal, whether supervised or unsupervised.
-This neural code exhibits robust capabilities for handling out-of-distribution data and effectively represents objects composed of fundamental features encountered during training—a task at which autoencoders (Hinton and Salakhutdinov, 2006) fail.
-2 Relation to Alternative Concepts
-The model we describe below is the first to demonstrate in concrete form the generation of structured nets that form a neural code to represent visual structure.
-To grasp the significance of the approach, it is important to compare it to alternative approaches.
-2.1 Associative Memory
-Associative memory (AM), as formulated by Amari (1972) or Hopfield (1982), shares with the present study the aim of retrieving previously encountered structure, and in both, the structure is recorded with the help of Hebbian synaptic plasticity.
-Whereas AM deals with the storage and retrieval of neural firing patterns, our purpose is to activate networks, that is, connectivity patterns that are structured by previously recorded connectivity pattern fragments.
-Admittedly, in both cases, the (re)construction process results in the activation of subsets of all neurons of the area, but whereas in AM, the connectivity of the set of active neurons is all-to-all, in our case, connectivity is restricted to within the net fragments that have previously been recorded, which in our visual application are restricted to short-range (local) connections and express two-dimensional topological structure.
-It is this structured connectivity that serves as basis for the structure-sensitive matching mentioned in the introduction.
-The monolithic all-to-all connectivity of the AM, by distinction, does not afford such structure-sensitive matching (i.e.
-instead of representing each structure as monolithic whole, the CNA learns to decompose them into local sub-patterns (net fragments), which can then be dynamically composed into nets to represent structures).
-This functional difference is conditioned on a different mode of learning.
-Whereas in AM, a neural pattern is stored in a single event by strengthening the connections between any pair of currently active neurons, our approach distinguishes significant from insignificant connections, significant being those that are strengthened again and again, while insignificant are those that are strengthened only occasionally.
-This learning is driven by exposure to a large number of patterns, none of which may ever repeat, but which contain small repeating subpatterns.
-It is these that generate net fragments.
-In the visual case, it is local texture patches that play this role of repeating sub-patterns.
-A basic limitation of AM is that unless the patterns to be stored are statistically independent (orthogonal), that is, have small neuron-wise overlap, they interfere with each other by the spurious activation of neurons.
-The problem can be reduced (as proposed by Tsodyks and Feigel’man (2007) or Palm (2013)) if patterns are sparse, that is, activate only a small fraction of all neurons.
-The statistical mode of pattern storage proposed here opens a more fundamental solution to the problem of pattern overlap.
-We propose that each input feature (specified by feature type and position) is represented not by one neuron but by a number of ‘alternative neurons’, neurons that have the exact same receptive field (for examples, see Cossell et al.
-(2015)), and thus are activated by the same afferents  [^1] There is ample space for alternative neurons, as there are 100 times more neurons in primary visual cortex than output neurons in the retina (Leuba and Kraftsik, 1994)..
-While these neurons have the same afferent input, they can exhibit different excitatory output.
-Given this expansion, different net fragments with feature-wise overlap can reduce neuron-wise overlap by making use of different alternative neurons.
-AM and our approach differ in yet another aspect, the mode of retrieval (or, in our case, construction) of patterns.
-In the version of Hopfield (1982) of AM, this is realized by minimizing an energy function by sequential switching of neurons.
-In our approach, winning patterns are projected out from the set of all neurons activated by the input and by rising inhibition that silences all those alternative neurons that do not have sufficient recurrent excitatory support, no neuron-sequential process being necessary.
-2.2 Boltzmann Machine
-Some important aspects of the system illustrated by our model are highlighted best by comparing it with the Boltzmann Machine (BM) proposed in Hinton and Sejnowski (1983).
-Both systems share features such as the ability to admitting recurrent connections (as does Mumford (2003)), but they also exhibit significant differences.
-Nevertheless, as our focus is limited to one area and is not the full architecture as proposed in von der Malsburg et al.
-(2022), we highlight only select differences here.
-The BM makes a clear distinction between visible and hidden (‘latent’) neurons.
-In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by inhibition in favor of those neurons that support each other through previously learned connections.
-A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
-In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference that we do not insist on symmetric connections and therefore do not derive an energy function and consequently cannot prove convergence of the neural dynamics (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the missing of up to half of the connections due to lack of symmetry does not jeopardize stability of a state).
-In a single cycle of the periodically rising inhibition, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
-2.3 Artificial Neural Networks (ANN)
-In the concept pursued here, objects and other structures within the input are represented by nets, sets of neurons that are held together by cooperative excitatory connections and are activated as a whole, the activation of individual neurons being conditioned on the activation of other neurons in the same net.
-In contrast, a large variety of ANN versions and learning techniques have been proposed, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), autoencoders (Hinton and Salakhutdinov, 2006), or representation learning (Bengio et al., 2014).
-Common to all of these is that the sensory input is converted through many layers into a representation in the form of a set of neuron states.
-These neurons can be considered as propositions that indicate the presence of complex features (activity patterns) within the input.
-This representation of the input is then exploited by evaluation circuits (often in the form of a few fully connected layers of neurons) to achieve functions such as object classification, next-token prediction, or, in autoencoders, reconstruction of the input.
-The common feature of all ANN systems is that the connections from the input to the representation layer and from there to the output are purely feedforward and are trained by back-propagation of error (Rosenblatt, 1962, p.
-292).
-The CNA proposes a clear separation between input representation (within an area) and its functional evaluation (between areas, see Section 6).
-The representation of the input is trained by its statistics and is thus essentially neutral and without specialization to any specific function, giving it foundational character.
-In ANN systems, the selection of patterns to be extracted from the input is strongly biased to serve the trained function, which may be the reason that a transformer trained on one set of functions cannot continue to be trained on another (Dohare et al., 2024).
-ANN’s hierarchy of layers of neurons responding to features of intermediate complexity is replaced in the CNA by a single layer, in which net fragments of different sizes represent a hierarchy of features of varying complexity (and inter-mesh to form coherent nets standing for whole objects, thus solving the binding problem).
-The purpose of transition to other areas is, in the CNA, not transition to higher complexity but transition to representations that generalize over, e.g., position, size, or orientation of nets in lower areas, as discussed in Section 6.
-In ANN systems, the high-level feature neurons that are activated by an input do not express any relations between them.
-This ‘binding problem’ (von der Malsburg, 1981; Burwick, 2014; Greff et al., 2020)) can give rise to ambiguity, e.g., the hallucination of feature conjunctions (Treisman and Schmidt, 1982).
-This ambiguity can be reduced by additional neurons whose trigger patterns overlap those of other neurons whose relative position would otherwise be in doubt (autoencoders (Hinton and Salakhutdinov, 2006) show that ambiguity can be suppressed altogether by a sufficiently large number of hidden neurons).
-In the CNA, the binding problem is solved in a natural way by net fragments that interconnect sets of features that co-occur with statistical significance in the input and that merge together into nets representing coherent chunks within a given input.
-This simplicity of representation of visual structure in CNA is due to its fundamentally different perspective on neural representation, in that the active state is not only described by the elementary features that are active but also by the connections between them.
-This is made possible by different alternative neurons that represent the same elementary feature but express different connectivity.
-3 Computational Framework
-Figure 2: The dynamics of a single area within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations.
+Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Sager et al., 2025).
+Yet, despite these advances, artificial vision systems continue to struggle with robust object representation and recognition, particularly under varying viewing conditions, occlusion, or noise (Qi et al., 2023; Simmler et al., 2021).
+This limitation highlights a fundamental gap between current AI systems and the flexible, context-aware perception observed in biological vision.
+Existing approaches attempt to address these challenges through methods such as feature or sample engineering (Amirian et al., 2018; Moosavi-Dezfooli et al., 2016), domain adaptation (Csurka, 2017; Sager et al., 2022), and data augmentation (Carlucci et al., 2019; Tuggener et al., 2024).
+However, these techniques largely operate within the conventional paradigm of representing objects as static patterns of neuronal activations, often requiring large datasets and extensive supervision to achieve generalization.
+In contrast, we propose a new paradigm by introducing the Cooperative Network Architecture (CNA), a framework in which object representations emerge from dynamic interactions among recurrently connected neurons.
+Rather than encoding objects as distinct activation vectors, CNA organizes sensory information into “nets” – structured networks of mutually supporting neurons.
+These nets are assembled from overlapping “net fragments,” which encode frequently co-occurring feature constellations and are acquired through Hebbian plasticity (Hebb, 1949).
+This compositional structure supports robust recognition, denoising, and figure completion through cooperative dynamics without external supervision.
+In this paper, we present the first concrete implementation of CNA and demonstrate its application to a simple vision task. Our contributions include
+  • a formal mathematical definition of nets and net fragments,
+  • a computational framework for learning net fragments from sensory input,
+  • an illustration of how these fragments can be flexibly composed into object-spanning nets,
+  • and an analysis of the emergent filtering and generalization properties of nets.
+Our results show that CNA offers a biologically inspired computational framework that integrates local and global information processing.
+It provides a foundation for more robust and adaptable perception systems, aligning with theoretical proposals for dynamic, network-based neural coding (von der Malsburg, 2018; von der Malsburg et al., 2022).
+By establishing the neural code and core computational mechanisms within a single representational area, this work provides the essential groundwork for future architectures in which multiple areas cooperate to support invariant object recognition – thus advancing toward the full potential envisioned in network-based neural coding models of perception.
+2 The Net Fragment Framework
+Before discussing the computational framework, we first introduce its building blocks: neurons, activations, nets, and net fragments.
+These definitions provide the basis for the discussions in the subsequent sections.
+For simplicity, we limit the definitions to a single area even though it could be extended to multiple inter-connected areas.
+Neurons.
+A neuron $i$ is denoted by $n_{i}$.
+Each neuron $n_{i}$ has a local neighborhood, denoted by $N_{r}(n_{i})$, which is the set of neurons $n_{j}$ such that the distance $d(n_{i},n_{j})$ between $n_{i}$ and $n_{j}$ is less than or equal to a threshold radius $r>0$, i.e.,
+$N_{r}\left(n_{i}\right)=\{n_{j}\mid d(n_{i},n_{j})\leq r\}.$ (1)
+At each time step $t$, the activity state of a neuron $n_{i}$ is represented by $y_{i,t}\in\{0,1\}$, where $y_{i,t}=1$ indicates that the neuron is active at $t$.
+The activity of a neuron over time is characterized by an activation pattern, which is a sequence of activation states $\{y_{i,t}\}_{t=1}^{T}$, where $T$ is the total number of time steps (sequential processing steps).
+Connections.
+Neurons interact via connections.
+A direct connection from neuron $n_{i}$ to another neuron $n_{j}$ is defined by a positive weight $w_{i,j}>0$, which facilitates the propagation of neuronal activity from $n_{i}$ to $n_{j}$, so that the activity $y_{i,t}$ of neuron $n_{i}$ can influence the activity $y_{j,t+1}$ of neuron $n_{j}$ at the subsequent time step.
+In this work, we limit the connections to local neighborhoods:
+$\forall w_{i,j}\quad n_{j}\in N_{r}(n_{i})$ (2)
+Additionally, we limit the weights to a range between 00 and $1$, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
+$w_{i,j}\in[0,1]$ (3)
+We define a relation $n_{i}\sim n_{j}$ between two neurons to indicate that their activation patterns are highly correlated, i.e., $y_{i,t}$ often influences $y_{j,t+1}$.
+Under Hebbian learning (Hebb, 1949), the connection strength $w_{i,j}$ increases when there is $n_{i}\sim n_{j}$.
+Net Fragments.
+A net fragment is defined as the set of neurons that provide direct activation support to a given neuron.
+In the context of this work, a neuron $n_{j}$ is said to provide support to another neuron $n_{i}$ if two conditions are met: (1) there exists a high correlation between two neurons $n_{j}$ and $n_{i}$ such that $n_{j}\sim n_{i}\Rightarrow w_{ji}>0$; and (2) neuron $n_{j}$ was active at the previous time step, i.e., $y_{j,t-1}=1$.
+Since synaptic weights are bounded within the interval $w_{j,i}\in[0,1]$, the degree of support a neuron can provide is likewise restricted to this range.
+Formally, the net fragment associated with neuron $n_{i}$ at time $t$ is given by
+$\mathcal{F}_{t}(n_{i})=\{n_{i}\}\cup\{n_{j}\mid w_{j,i}>0\land y_{j,t-1}=1\}.$ (4)
+In this formulation, any neuron $n_{j}\in\mathcal{F}(n_{i})$ contributes to the activation likelihood of $n_{i}$ at time $t$.
+Notably, neuron $n_{i}$ is part of its own fragment, reflecting a form of self-support – if active at time $t$, the neuron may contribute to sustaining its own activation at the subsequent time step.
+Note that $n_{i}\sim n_{j}$ is not an equivalence relation since correlation is not transitive, meaning that one neuron can be part of multiple net fragments.
+Activation Function.
+The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment.
+We call this accumulated input the membrane potential at timestep $t$, which is a function $\mathcal{M}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
+$a^{\mathcal{M}}_{i,t}=\mathcal{M}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
+The membrane potential $a^{\mathcal{M}}_{i,t}\in\mathbb{R}$ is large when many other neurons provide support, meaning that they “confirm” the plausibility of a neuron’s activity by being active themselves and exhibiting a connection that has been strengthened through previous mutual activity and Hebbian updates.
+The activation of a neuron at timestep $t$ is a function of the membrane potential $a_{i,t}$ and environmental conditions $\mathcal{E}$:
+$y_{i,t}=\mathcal{A}\left(a_{i,t},\mathcal{E}\right)\in\{0,1\}$ (6)
+The environment conditions $\mathcal{E}$ include an attenuation signal, a constraint that selectively suppresses weak excitatory inputs by requiring neurons to accumulate a sufficiently large membrane potential before firing.
+Nets.
+Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$).
+A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects.
+For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active:
+$\mathcal{F}^{*}_{t}(n_{i})=\mathcal{F}_{t}(n_{i})\quad\text{with}\quad y_{i,t}=1$ (7)
+To account for both direct and indirect influences on a neuron’s activation, we use a recursive formulation of nets from the perspective of a single neuron.
+The base case is defined as the immediate net fragment:
+$\mathcal{N}_{t}^{1}[n_{i}]=\mathcal{F}^{*}_{t}(n_{i})$ (8)
+For $r\geq 2$, the recursive relation is given by
+$\mathcal{N}_{t}^{r}[n_{i}]=\mathcal{N}_{t}^{r-1}[n_{i}]\bigcup_{n_{j}\in\mathcal{N}_{t}^{r-1}[n_{i}]}\mathcal{F}^{*}_{t}(n_{j})$ (9)
+The complete network accessible from $n_{i}$ is then defined as the closure
+$\mathcal{N}_{t}^{*}[n_{i}]=\bigcup_{r=1}^{\infty}\mathcal{N}_{t}^{r}[n_{i}],$ (10)
+which comprises all neurons in the active subgraph that are reachable from $n_{i}$.
+This recursive formulation illustrates how local net fragments, which can be interpreted as elementary features, are progressively integrated to form higher-order features.
+Moreover, the explicit time dependence in these definitions underscores the dynamic nature of both the net fragments and the overall network, as they are continuously reassembled in response to the evolving support received from neighboring neurons, and increasing attenuation that prevents neurons with a low membrane potential from firing (see next section).
+In the proposed architecture, no symmetry constraint is imposed on the weights, that is, $w_{i,j}\neq w_{j,i}$ in general.
+However, due to the temporal dynamics of the network, activations tend to stabilize over a longer time horizon $T$, such that for sufficiently large $t$, the activations become approximately stationary, i.e., $y_{i,t}\approx y_{i,t+1}$.
+Under these conditions, mutual influence between neurons becomes symmetric, meaning that $n_{i}\sim n_{j}$ implies $n_{j}\sim n_{i}$, and almost symmetric weights emerge naturally over time, even in the absence of explicit symmetry constraints.
+As a result, it can be assumed that for any pair of distinct neurons $n_{i}\neq n_{j}$, their closures either converge to the same set of neurons or remain largely disjoint.
+Formally, this can be expressed as:
+$\mathcal{N}_{t}^{*}[n_{i}]\approx\mathcal{N}_{t}^{*}[n_{j}]\quad\text{or}\quad\mathcal{N}_{t}^{*}[n_{i}]\cap\mathcal{N}_{t}^{*}[n_{j}]\approx\emptyset.$ (11)
+3 Cooperative Network Architecture
+Figure 2: The dynamics within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations.
 The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer.
-In part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the recurrent connections within $S2$.
-In part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
-Our framework addresses the visual system, even though it could be applied to other modalities as well.
-In the primary visual cortex, images appear as two-dimensional arrays of activated neurons, and its neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004).
-To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (‘convolutional’ structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift.
-We use these insights to build a CNA with two stages, as shown in Figure 2(A).
-The CNA is based on an input layer representing observed images by simple feature detectors and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$).
-Feature Extraction.
-The system’s input is a retinal 2D image represented by variables $\boldsymbol{x}_{k,i}$ (gray box in Figure 2), where $k\in(1,...,C)$ denotes the channel index of the image ($C=1$ for grayscale, $C=3$ for RGB), and the composite index $i=(h,w)\in(H\times W)$ stands for vertical and horizontal position of pixels within the input image.
-Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their ‘membrane potential’ $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2).
-The membrane potential $\boldsymbol{a}^{(S1)}$ is calculated by applying a convolutional operation (LeCun et al., 1989) (yellow box $S1$ in Figure 2):
-$\boldsymbol{a}^{(S1)}_{c,j}=\sum_{k\in C}\sum_{i\in l_{j}}\boldsymbol{W}^{(S1)}_{c,k,j-i}\cdot\boldsymbol{x}_{k,i}$ (1)
-Here, $c\in C^{(S1)}$ denotes the feature channel, the compound index $j$, expanded as $j=\{h,w\}$, designates again spatial position $h\in H$, $w\in W$, and the compound index $i=\{h,w\}$ is running through all $(h,w)$-pairs in the two-dimensional range $l_{j}=(h,....,h+h^{(W1)};w,....,w+w^{(W1)})$ where $h^{(W1)}$, $w^{(W1)}$ denote the size of the convolutional filter.
-The size of the convolutional kernel used for feature extraction is $\boldsymbol{W}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times C\times h^{(W1)}\times w^{(W1)}}$.
-Feature Binarization.
-The membrane potential $\boldsymbol{a}^{(S1)}$, together with a bias parameter $b^{(S1)}$, determines the binary output of the neurons in $S1$, denoted as variable $\boldsymbol{y}_{c,j}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$, where $1$ stands for firing, 00 for silence.
-The binary neuron model with the input-output function $\mathrm{B}:\boldsymbol{a},b\rightarrow\boldsymbol{y}$ is defined as
-$\boldsymbol{y}_{c,j}^{(S1)}=\mathrm{B}(\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)})=\begin{cases}1,&\text{ if }\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)}>0\\ 0,&\text{otherwise}\end{cases}$ (2)
-with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (neurons fire if the membrane potential is above this threshold).
-Dynamics.
-Stage $S2$ builds net fragments by allowing active neurons to support recurrently connected neurons over T time steps (depicted as recurrent connection in Figure 2).
-A sample image is introduced into the system at time step $t=0$ to activate neurons and remains unchanged until time step $t=T$.
-Throughout this period, neuronal activity undergoes sparsification due to growing inhibition (see below).
-Through the sparsification process, the system forms an object representation by leaving active a set of neurons each of which is supported by active neighbors, that is, a net.
-Short-Range recurrent Connections.
-Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feedforward connections between different layers, we employ them here to implement short-range recurrent connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments.
-Given recurrent connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$.
-In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state to project out a net representing a given sensory signal.
-Importantly, the internal state evolves continuously, even when processing the same input pattern, driven by dynamic changes in the support between neurons along recurrent connections.
-The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by recurrent excitation from within $S2$.
-Before digitization, they are modified by saturation, normalization, and inhibition (yellow box $S2$ in Figure 2).
-Before those modifications, its raw form is computed as:
-$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Recurrent Connection}}.$ (3)
-Similar to $S1$, $c\in C^{(S2)}$ denotes the feature channel of $S2$, the compound index $j$ expands to $j=\{w,h\}$, $w\in W$, $h\in H$, and the compound index $i$ is running in the range $l_{j}=(h,....,h+h^{(W2)};w,....,w+w^{(W2)})$, with $h^{(W2)}$ and $w^{(W2)}$ being the size of the convolution kernel.
-The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the recurrent connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e.
-$C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments).
-In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the recurrent connections determines which neurons can be connected within stage 2.
-The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix A).
-The weights are updated using Hebbian Learning (details in Appendix B).
-For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide recurrent support:
-$\boldsymbol{y}_{i}^{(S2)}[t=0]=0$ (4)
-Single Weight-Kernel Implementation.
-This entire process is visually depicted in the upper part (A) of Figure 2.
-For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
-$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(S2)}_{c,k,j-i}\cdot\left(\boldsymbol{y}_{k,i}^{(S1)};\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)$ (5)
-The net fragments are implicitly contained in this processing as the active neurons whose activity state is stored in $\boldsymbol{y}^{(S1)}$ and their connection to other neurons.
-Inhibition.
-The raw neural excitation $\boldsymbol{a}^{\prime}$ computed in eq.
-3 (or eq.
-5) is to be subjected to saturation, inhibition, and normalization so it can be binarized using a pre-defined threshold.
-These steps of processing the raw membrane potential $\boldsymbol{a}^{\prime}$ of neurons proved necessary in our simulation experiments to achieve the qualitative function we aimed for.
-The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix C.
-The differences between activation levels of neurons are accentuated by setting
-$\boldsymbol{a}^{(S2)}_{c,j}[t]=\max\left(\left(\boldsymbol{a}^{\text{norm }(S2)}_{c,j}[t]\right)^{\gamma},0\right),$ (6)
-The coefficient $\gamma$ regulates the inhibition strength and increases with the time-steps of the inner loop according to $\gamma=1.2+0.2t$, making it harder for neurons to fire (as $\boldsymbol{a}^{\text{norm }(S2)}\leq 1$).
-The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive recurrent support, while invalid activations are less supported.
-While the inhibition coefficient increases over time, many neurons eventually get suppressed, leaving only those active that support each other and form a consistent net.
-Finally, the membrane potential is converted to a binary activation, neurons becoming active when their inhibited membrane potential is above the activation bias $b^{(S2)}$, which serves to control the firing rate of neurons:
-$\boldsymbol{y}_{c,j}^{(S2)}[t]=\mathrm{B}\left(\boldsymbol{a}^{(S2)}_{c,j}[t]-b^{(S2)}\right)$ (7)
-The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the recurrent connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed.
-Alternative neurons are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero).
-This selection process process is described in the Appendix D.
-Relation to Nets and Fragments.
-Net fragments are learned through the strengthening of recurrent connections $\boldsymbol{W}^{(L)}$, facilitated by Hebbian plasticity.
-Neurons within a fragment are interconnected, providing mutual support that enhances their membrane potentials.
-By shifting $\boldsymbol{W}^{(L)}$ as convolutional kernels across the input, fragments can be detected at all spatial locations, resulting in numerous overlapping fragments.
-Coherent fragments are interconnected to form a larger structure, a net.
-While multiple nets may exist, projecting out the most coherent one is crucial for robust object representation.
-This is achieved by progressively increasing inhibition over time.
-As inhibition intensifies, neurons with insufficient support (e.g., only a few instead of many neurons within one net fragment were triggered) are deactivated [^2] Neuron activations can be interpreted as hypotheses that require confirmation/support from other neurons with compatible hypotheses - otherwise, they are deactivated..
-This deactivation can propagate, leading to the elimination of entire fragments and leaving a single coherent net to represent the input structure.
-In our experiments (see below), we observe that, initially, neurons operate independently due to the absence of established recurrent connections, with a single neuron having a membrane potential of $1$ being sufficient for activation.
-However, after training, the average membrane potential rises to approximately $12$, reflecting robust support from neurons within a fragment.
-As a result, inconsistent neurons can no longer activate (and neurons with a membrane potential of $1$ are switched off), ensuring a more stable and coherent representation.
+In this section, we present a concrete instantiation of the proposed framework, specifically applied to the visual domain – although the framework is general and may be extended to other sensory modalities in the future.
+We refer to this implementation as the Cooperative Network Architecture (CNA).
+To realize our framework, we adopt a standard convolutional neural network (CNN) architecture (LeCun et al., 1989).
+The CNA comprises two processing stages, depicted as yellow squares in Figure 2.
+The stage 1 ($S1$) performs fixed (i.e., not learned) feature extraction directly from the image input.
+The stage 2 ($S2$) implements the core principles of the described computational framework by learning net fragments.
+The network dynamics unfold over discrete time steps $t=[0,\dots,T]$.
+Since the feature extraction stage ($S1$) employs fixed filters, it yields a time-invariant activation pattern for a given input.
+In contrast, the second stage ($S2$) exhibits dynamic temporal evolution, where internal representations are iteratively refined at each time step ($t\to t+1$).
+Specifically, we use a progressively increasing global attenuation signal that only allows neurons with a high membrane potential to remain active.
+During this dynamic processing, a neuron might be deactivated, whereby another neuron’s membrane potential is decreased, triggering a chain reaction that can lead to the deactivation of incoherent nets.
+While these refinements occur within the processing cycle of a single sample ($[0,\dots,T]$), Hebbian learning adapts the weights (connections $w_{i,j}$ within net fragments) on longer timescales, wherein synaptic weight modifications occur between different sample-processing episodes, leading to structural plasticity in the network.
+The following sections discuss these two processing stages.
+3.1 Stage 1: Feature Extraction
+Stage 1 is designed to model the sensory input by extracting initial neural activations from the raw image data.
+This extraction is essential due to the continuous nature of typical image inputs (e.g., grayscale or RGB), whereas the computational framework employed in this work operates on binary activations.
+Consequently, this stage serves as a preprocessing step, enabling the transformation of continuous-valued images into a suitable binary representation.
+The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels ($1$ for grayscale, $3$ for RGB), while $H,W$ stand for the height and width of the image.
+Activation Computation.
+To extract initial activation patterns, we use fixed, non-trainable filters (details see Appendix A).
+These filters are implemented as standard convolutional kernels, used to obtain the binary activation pattern $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)},H,W}$ as follows:
+$\boldsymbol{y}^{(S1)}_{c_{\text{out}}}=\Theta\left(\sum_{c_{\text{in}}=0}^{C}\boldsymbol{W}^{(S1)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{x}_{c_{\text{in}}}\right)$ (12)
+Here, $\star$ denotes the convolution operator between the filters $\boldsymbol{W}^{(S1)}\in\mathbb{R}^{C^{(S1)},C,h,w}$ and the input image $\boldsymbol{x}$, while the parameters $h,w$ define the width and height of the convolutional filters, i.e., the spatial neighborhood from where the features are extracted.
+To binarize the activations, we use the heaviside function $\Theta$, which sets a neuron to one if its input is larger than zero, and to zero otherwise.
+$y_{i}=\Theta(a_{i})=\begin{cases}1\quad\text{if }a_{i}>0\\ 0\quad\text{otherwise}\end{cases}$ (13)
+Note that the width and height between input and output activations remain identical, but the number of channels changes from $C$ to $C^{(S1)}$, allowing the extraction of multiple features at each spatial location.
+3.2 Stage 2: Net Fragments
+$S2$ constructs net fragments from the binary feature activations $\boldsymbol{y}^{(S1)}$ of $S1$ and integrates them into coherent, spatially extended nets over time.
+To ensure stability, single neurons or small groups of neurons are suppressed through an attenuation mechanism, ensuring that only large-scale, consistent activation patterns persist.
+Network Topology.
+Stage 2 uses convolutional layers to implement the connections $w_{j,i}$ within net fragments $n_{j}\in\mathcal{F}(n_{i})$.
+Conceptually, this stage comprises a three-dimensional array of neurons with dimensions $C^{(S2)}\times H\times W$, where $C^{(S2)}$ denotes the number of feature channels in $S2$, and $H$ and $W$ correspond to the spatial dimensions of the input image.
+The spatial distance between two neurons $n_{i}$ and $n_{j}$ is defined using the maximum norm (also known as Chebyshev distance), where adjacent neurons have a unit distance:
+$d(n_{i},n_{j})=\max\left(|h_{i}-h_{j}|,|w_{i}-w_{j}|\right)$ (14)
+Each neuron $n^{(S2)}_{i}$ in stage 2 is connected in two ways: (1) Feedforward connections to the corresponding neuron $n^{(S1)}_{i}$ in stage 1 and to its local neighborhood $Nr(n^{(S1)}_{i})$, and (2) lateral connections to neighboring neurons within stage 2 itself, denoted as:
+$Nr\left(n^{(S2)}_{i}\right)=Nr\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r_{2}\right\}$ (15)
+where $r_{2}$ defines the radius of lateral connections.
+The convolutional architecture realizes these neighborhood structures via convolutional filters, where the filter size directly determines the radius of both feedforward and lateral connections.
+Using convolutional filters not only implements the connectivity restrictions to local neighborhoods but also allows detection of the same pattern at various positions, enabling learning of net fragments independent of position.
+The forward connections from $S1$ to $S2$, denoted as $\boldsymbol{W}^{(F)}$, are necessary to activate cells within $S2$.
+It can be interpreted as a small group of neurons in $S1$ (based on the filter size of $\boldsymbol{W}^{(F)}$) supporting a single neuron within $S2$.
+In contrast, the lateral recurrent connections within $S2$, denoted as $\boldsymbol{W}^{(L)}$, are responsible for building the overlapping net fragments within $S2$, providing support amongst the neurons within $S2$.
+It is important to note that the lateral connection matrix $\boldsymbol{W}^{(L)}$ does not implement recurrent processing in the traditional sense of sequential token-wise updates, as found in RNN-based language models.
+Instead, $\boldsymbol{W}^{(L)}$ iterates over the same neuronal state vector $\boldsymbol{y}^{(S2)}$.
+Specifically, the lateral connections $\boldsymbol{W}^{(L)}$ calculate the state updates of $\boldsymbol{y}^{(S2)}_{t\rightarrow t+1}$ within an iterative refinement process in which individual neuronal activations ${y}_{i}^{(S2)}\in\boldsymbol{y}^{(S2)}$ may flip due to changes in support and attenuation.
+Membrane Potential.
+To determine the support a single neuron $n_{i,t}$ receives from the net fragment $\mathcal{F}_{t-1}(n_{i,t-1})$, we first calculate the membrane potential $\mathcal{M}$ by applying convolutional operations on the forward and lateral connections and then taking the sum of these operations:
+$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\underbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S1)}}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}}}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)}^{\textit{Recurrent Connection}}.$ (16)
+Here, the forward connections have the shape $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},C^{(S1)},h,w}$ and the recurrent connections have the shape $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},C^{(S2)},h,w}$, where $C^{(S2)}$ is the number of channels in $S2$.
+Therefore, the membrane potential $\mathcal{M}$ at time $t$ is increased by the learned amount $w_{j,i}$ from all connected and active neurons ($y_{j,t-1}=1$).
+Note that $\mathcal{F}_{0}(n_{i,0})$ only consists of neurons of $S1$ since all neurons in $S2$ are initialized to silent state ($y_{i,0}=0$).
+Higher support, indicated by larger $a^{\mathcal{M}(S2)}_{i}$, suggests a more likely pattern, as neurons encoding frequent patterns gain reinforcement not only from feedforward connections but also from lateral connections.
+Before we discuss how the weight matrices $\boldsymbol{W}^{(F)},\boldsymbol{W}^{(L)}$ are initialized and trained, we describe how the membrane potential $\boldsymbol{a}^{\mathcal{M}(S2)}$ undergoes attenuation and is transformed into $\boldsymbol{y}^{(S2)}$.
+Attenuation.
+The membrane potential is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{M}},\mathcal{E}\right)$.
+In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{M}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range $[0,1]$ (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active.
+Specifically, attenuation is implemented as:
+$\boldsymbol{a}^{(S2)}_{c,t}=\left(\boldsymbol{a}^{\text{norm }(S2)}_{c,t}\right)^{\gamma},$ (17)
+where $\gamma$ regulates the attenuation strength and increases over time according to $\gamma=\alpha+\beta t$ (when not stated differently, we set $\alpha=1.2$ and $\beta=0.2$), making activation progressively harder.
+This process leads to a sparsification of activations that occurs through both direct attenuation and reduced support from previously silenced neurons .
+The final output is then binarized using the heaviside function $\Theta$:
+$\boldsymbol{y}_{c,t}^{(S2)}=\Theta\left(\boldsymbol{a}^{(S2)}_{c,t}-b^{(S2)}\right)$ (18)
+The activation at the last timestep $\boldsymbol{y}^{(S2)}_{t=T}$ represents a recurrently refined activation pattern where unsupported neurons have been suppressed, yielding a coherent representation.
+Alternative Neurons.
+A single neuron that encodes a specific feature can be activated in multiple contexts.
+For example, a short straight-line segment may appear in various contexts – such as in different line drawings or digits.
+In our framework, each context is represented by a net, whereas individual features correspond to the activity of specific neurons.
+To achieve context disentanglement, it is important to allocate sufficient representational capacity such that the same feature (e.g., a line segment) can be utilized across distinct contexts (e.g., different line drawings) without interference.
+This requires that a feature-representing neuron is capable of participating in multiple, context-specific net fragments.
+Without this capacity, a neuron would be forced to average over competing contextual associations, resulting in a non-specific or “blurry” representation.
+Formally, a neuron representing a feature may be part of multiple net fragments: $n_{i}\in\mathcal{F}(n_{j})$ and $n_{i}\in\mathcal{F}(n_{k})$ where $n_{j}\neq n_{k}$).
+This implies that the neuron must form distinct lateral connections $w_{i,j}$ and $w_{i,k}$, which should not be concurrently active.
+To address this requirement, we introduce “alternative neurons”.
+Each alternative version of a neuron is activated by the same feature from $S1$ but has a distinct set of lateral connections in $S2$ to encode different contexts.
+This is implemented by increasing the channel dimension $C^{(S2)}$ of $\boldsymbol{W}^{(L)}$ by a factor $\kappa$, i.e.,
+$C^{(S2)}:=C^{(S2)}\cdot\kappa$ (19)
+This expansion provides capacity for $\kappa$ alternative version of a neuron.
+Consequently, these alternative neurons receive identical bottom-up input from $S1$ while differing in the lateral support they receive in $S2$.
+A winner-take-all (WTA) competition mechanism ensures that each alternative neuron specializes in a distinct context, with only one being activated in any given instance.
+[…]
+At initialization, neuron selection is random.
+However, after Hebbian learning updates, the alternative neurons progressively specialize, leading to distinct and structured fragment representations.
+Over time, the most correlated alternative neuron is chosen for each input location, ensuring robust feature specialization across different contexts.
+Weight Initialization.
+The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all alternative $S2$ neurons in the same position.
+This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to $1$ if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of alternative neurons.
+The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to $1$.
+Learning of Weights.
+Support within net fragments is learned via Hebbian plasticity (Hebb, 1949), so that frequently occurring patterns that lead to correlated activation patterns between neurons $n_{i}\sim n_{j}$, are “stored” in the system and receive more support:
+$\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\eta\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (20)
+Here, $\eta$ represents the learning rate and $\boldsymbol{W}^{(S2)}$ stands for both $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$.
+Note that the same convolutional kernel is applied at various spatial positions.
+The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects (across all spatial positions).
+If the resulting average is positive, the corresponding connection increases as it connects more simultaneously active neurons than neurons that fire disjointly.
+Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced.
+Learning is applied after processing a sample (after $T$ timesteps) to improve stability.
 4 Experiments
-To demonstrate the compositionality of net fragments, we train (updating $\boldsymbol{W}^{(L)}$ with Hebbian plasticity) the model on binary images containing straight-line patterns.
-For evaluation (after training and without updating any connections), we use images of kinked lines, digits, characters, and line drawings of objects such as houses and clocks.
-These complex patterns can be expressed as compositions of straight-line net fragments, even though the CNA was not explicitly trained on such patterns.
-This flexible compositional behavior contrasts strongly with associative memory (see Section 2.1).
-To highlight the differences between ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
-Besides compositionality, we introduce noise to examine the robustness and persistence of coherent structures with statistical relevance.
-Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects, i.e., the efficacy of dealing with subtractive noise.
-We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
-Further experimental details are provided in the Appendix: We describe the training and testing datasets in more detail in Appendix E, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix F, the training details of our CNA model are in Appendix G, and the training details of the autoencoder in Appendix H.
+To demonstrate the compositionality of net fragments into nets, we train the model on binary images containing straight-line patterns so that the CNA learns net fragments corresponding to straight-line sub-patterns (smaller straight-line segments).
+However, for evaluation (after training and without weight updates), we use much more complex images such as line drawings of objects, and demonstrate that these complex patterns can be expressed as compositions (nets) of straight-line net fragments.
+This flexible compositional behavior contrasts strongly with conventional associative memory that cannot model such out-of-distribution data (see Section 6.1).
+To highlight the differences to ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
+We also examine robustness by introducing noise.
+Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects.
+Further experimental details are provided in the appendix: We describe the training and testing datasets in more detail in Appendix D, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix A, the training details of our CNA model are in Appendix E, and the training details of the autoencoder in Appendix F.
 5 Results
-Figure 3: Net formed by activated neurons.
-The left of the image shows the neuronal activation $\boldsymbol{y}^{(S2)}$ when observing a line, and the right visualizes the net formed by these neurons.
-For convenience, the neurons are rearranged in a circle, and the connections of two randomly selected neurons, A and B, are visualized in colors (purple and red).
-We empathize that this work focuses on learning a new neural code and, in the following, demonstrates it on a rather simple dataset.
-Although our result figures display activated neurons, we would like to emphasize that these neurons are carriers of connections and form part of nets.
-To make this point for once, Figure 3 depicts this connectivity of activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged as a circle for convenience).
+Figure 3: The response of our model to a diagonal line input is displayed as a net of activated neurons.
+The left of the image shows the neuronal activation $\boldsymbol{y}^{(S2)}$ when observing the line, and the right visualizes the net formed by these neurons (activations $\boldsymbol{y}^{(S2)}$ with their connections $\boldsymbol{W}^{(L)}$).
+For convenience, the neurons are rearranged in a circle, and the net fragment (connections) of two randomly selected neurons, $n_{i}$ and $n_{j}$, are visualized in colors (purple and red).
+Although the figures illustrate individual neuron activity, it is important to clarify that our objective is not to interpret patterns merely as isolated neuron activations, but rather as nets – cohesive, interconnected clusters of active neurons.
+To emphasize this perspective, Figure 3 illustrates the connectivity among activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged in a circular layout for clarity), after feeding a diagonal line into the network.
+We select randomly two neurons $n_{i}$ and $n_{j}$ in the left panel and visualize their net fragments $\mathcal{F}(n_{i})$ and $\mathcal{F}(n_{j})$ based on $\boldsymbol{W}^{(L)}$ on the right.
+This figure highlights that individual neurons (e.g., neurons $n_{i}$ and $n_{j}$) are embedded within broader nets of supportive connections.
+Without such network-based support, these neurons would be suppressed via attenuation mechanisms and would not appear in the left panel of Figure 3, which displays neuronal activity.
+Consequently, the subsequent visualizations should be understood as representing supported neurons embedded within nets, even though the underlying connectivity is not explicitly shown.
 5.1 Compositionality
@@
 For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
-The model weights, $\boldsymbol{W}^{(S2)}$, are trained using straight lines (see training data at the top of Figure 4), allowing the model to capture the underlying structure of straight lines within net fragments.
-The model composes these fragments dynamically into nets to represent much more complex structures (see validation data in rows 1-8 of Figure 4).
-For instance, despite not encountering these images during training, the CNA model can represent kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows $5$-$8$).
-The motivation for why we build net fragments ($\boldsymbol{y}^{(S2)}$) and not just use the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise.
-This noise corresponds to invalid patterns not observed during training, which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons.
+The model weights are trained using straight lines (see training data at the top of Figure 4).
+During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 1-8 of Figure 4).
+Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 1 and 2), digits and characters (rows 3 and 4), and line drawings (rows 5–8).
+These results demonstrate that while individual net fragments $\mathcal{F}_{t}$ must be learned (in our case we use Hebbian learning to train the connections within $\mathcal{F}_{t}$), the recursive composition described in eq.
+9 to form $\mathcal{N}^{*}_{t}$ allows the integration of previously uncombined fragments.
+Consequently, the model exhibits combinatorial generalization through the structural recomposition of learned primitives.
+The motivation for why we construct net fragments rather than just using the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise.
+This noise corresponds to invalid patterns not observed during training (i.e., are not captured in net fragments), which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons.
 Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
 5.2 Filtering Gaussian Noise
-Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by recurrent connections of net fragments.
-In each image, the background is shown in black, and the four different features in distinct colors.
-The plots are created using a bias $b^{(S2)}=0.5$ and an inhibition coefficient $\gamma=1.2+0.2t$.
-The illustration in Figure 5 visually demonstrates the system’s effective capability of filtering out isolated noise or small clusters thereof.
-Noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix E) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments.
-While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated.
+Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
+To further investigate noise filtering, we introduce Gaussian noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix D) to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
+The results are visualized in Figure 5: While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated.
 As additional noise is introduced, more undesired neurons receive support and persist in their activity.
 However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$.
-In Appendix I, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
+In Appendix G, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
 Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed.
-However, when inhibition increases over time, even more noise is filtered out in the course of iteration.
-The plot is created after adding $10\%$ noise to each feature channel (sing a bias $b^{(S2)}=0.7$ and an inhibition coefficient $\gamma=1.2+0.2t$).
-Figure 6 depicts this systematic filtering over time.
-The additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time.
-The figure illustrates a substantial noise reduction already at time step $t=0$ due to the filtering by matrix $\boldsymbol{W}^{(F)}$ and initial inhibition.
+However, when attenuation increases over time, even more noise is filtered out.
+The plot is created after adding $10\%$ noise to each feature channel (using a bias $b^{(S2)}=0.7$).
+The reason for the filtering is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time.
+Figure 6 depicts this systematic filtering across timesteps.
+The figure illustrates a substantial noise reduction already at time step $t=0$ due to initial attenuation.
 In each subsequent time step, additional clutter is removed.
-This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons, demonstrating how overlapping fragments build a coherent net of mutually supporting neurons.
+This reduction is due to increased attenuation and falling support as a consequence of the drop-out of supporting neurons.
 5.3 Reconstructing Figures
-Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations.
-Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
-Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the ‘activation probabilities’ (on the right) after the first time step.
-If a substantial number of neurons activate and support other neurons within a net, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction.
+Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations.
+Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$.
+Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
+If a fragmet $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$).
 Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns.
-This reconstruction works only reliably for up to $3$ missing pixels (see Appendix I).
+This reconstruction is reliable for up to $3$ missing pixels (see Appendix G).
 In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
 Membrane Potential Represents Expected Feature Activation.
-When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected.
-Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 7B.
-Modeling this ambiguity enables having different internal interpretations, such as whether ‘it is an interrupted line that should be reconstructed’ or ‘there are two co-linear line segments’ and selecting one of them based on the threshold $b^{(S2)}$.
-Increasing Inhibition Over Time Stabilizes Patterns.
-Figure 8: Reconstruction of a horizontal line interrupted by $7$ pixels: The feature activation and the initial output activity at $t=0$ are interrupted but reconstructed within the two subsequent time steps ($t=1,2$).
-Afterward, some artifacts triggered at line ends are reduced due to increasing inhibition.
-Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=1.8+0.1t$.
-The reconstruction process unfolds across multiple sequential steps, as illustrated in Figure 8.
-Initially, at time step $t=0$, the net fragments represent the interrupted line.
-Subsequently, during the second time step ($t=1$), one of the removed pixels is reconstructed, and shortly after, at $t=2$, the entire line is represented.
-Thus, reconstruction can persist even amid increasing inhibition, given that inactive neurons receive adequate support.
-This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of neurons encourages others to follow.
+In the presence of missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected (as shown in Figure 7B).
+Depending on the context, it may be desirable either to reconstruct the missing line segment – effectively bridging the gap between discontinuous features – or to preserve the separation between distinct line elements.
+This behavior can be modulated by adjusting the hyperparameter $b^{(S2)}$ in eq.
+18, which can govern the balance between feature completion and preservation of discontinuity, thereby allowing the model to selectively favor line reconstruction or separation.
 5.4 Comparison to an Autoencoder
-Figure 9: Comparison of the outputs generated by the proposed CNA model and an autoencoder across different tasks.
+Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder.
 Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder).
@@
 Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
-Figure 9 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix H).
-While the autoencoder successfully learns to accurately represent line structures in the training data - achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and decoder reconstruction - it struggles with generalization to data not observed during training.
+Figure 8 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix F).
+While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training.
 For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
@@
 The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations.
-Metrics on noise reduction, precision, and recall, presented in Appendix I, further substantiate the CNA model’s superior performance.
-In row $6$, the models’ capabilities in reconstructing missing line segments are compared.
+Metrics on noise reduction, precision, and recall, presented in Appendix G, further substantiate the CNA model’s superior performance.
+In row $6$, the models’ capabilities for reconstructing missing line segments are compared.
 Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed.
-This is confirmed by the metrics in Appendix I: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of more noise at other locations, leading to lower overall precision and recall.
+This observation is confirmed by the metrics in Appendix G: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of increased noise elsewhere, leading to lower overall precision and recall for figure reconstruction.
+6 Relation to Alternative Concepts
+To grasp the significance of the approach, it is important to compare it to alternative approaches.
+6.1 Associative Memory
+Associative memory (AM) (Amari, 1972; Hopfield, 1982) and the model presented in this work share the objective of retrieving previously encountered patterns, both utilize Hebbian synaptic plasticity for encoding information.
+However, fundamental differences exist in their structural and functional organization.
+In classical AM, neurons exhibit an all-to-all connectivity pattern, leading to the representation of structures as monolithic wholes, whereas our model constrains connectivity within learned net fragments.
+This restriction enables a more modular representation of complex patterns by dynamically assembling net fragments into structured nets, thereby allowing the formation of representations that extend beyond the training data distribution.
+These differences arise from distinct learning mechanisms.
+In AM, a neural pattern is stored in a single event through the uniform strengthening of connections between co-activated neurons.
+Our approach, however, capitalizes on the observation that a large number of non-repeating patterns (as found in natural images) contain small recurring sub-patterns, such as texture patches.
+These recurring elements are captured within net fragments.
+The learning rule defined in eq.
+20 selectively reinforces connections corresponding to the most relevant sub-patterns.
+Specifically, frequent sub-patterns exhibit a consistently positive average correlation $\rho$, which leads to the progressive increase of their associated synaptic weights.
+In contrast, rare or non-informative patterns exhibit low or even negative average correlations, causing their associated weights to decay toward zero.
+This Hebbian mechanism effectively filters out noise and ensures that only the most salient and statistically dominant sub-patterns are preserved in the learned representation.
+A fundamental limitation of AM is the interference between stored patterns unless they are statistically independent (i.e., orthogonal), meaning that minimal neuron-wise overlap exists.
+Sparse coding, as proposed by Tsodyks and Feigel’man (2007) or Palm (2013), mitigates interference by activating only a small subset of neurons.
+Our model offers a more intrinsic solution by representing each input feature using multiple alternative neurons.
+These neurons possess identical receptive fields (as observed by Cossell et al.
+(2015)), but can exhibit different excitatory outputs.
+This redundancy enables overlapping net fragments to reduce neuron-wise interference by utilizing distinct alternative neurons.
+6.2 Boltzmann Machine
+Key aspects of our model become clearer when compared to the Boltzmann Machine (BM) introduced by Hinton and Sejnowski (1983).
+Both models allow recurrent connections (as also proposed in Mumford (2002)), but they also exhibit significant differences.
+Nevertheless, as our focus is limited to one area with net fragments ($S2$) and is not the full object detection architecture as proposed by von der Malsburg et al.
+(2022), we highlight only select differences here.
+The BM makes a clear distinction between visible and hidden (“latent”) neurons.
+In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by attenuation in favor of those neurons that support each other through previously learned connections.
+A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
+In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference being that we do not insist on symmetric connections and therefore do not derive an energy function (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the absence of up to half of the connections due to lack of symmetry does not jeopardize the stability of a state).
+However, even without an energy function, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
+6.3 Artificial Neural Networks (ANN)
+Artificial neural networks (ANNs) rely on various architectures and learning techniques, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), and autoencoders (Hinton and Salakhutdinov, 2006).
+These systems process sensory input through multiple layers to generate a neuronal representation.
+A defining characteristic of ANN systems is their strict feedforward architecture, where connections from input to representation and from representation to output are trained via backpropagation of error (Rosenblatt, 1962, p.
+292).
+In contrast, the neuronal representation within the CNA consists of nets that activate as a whole, with individual neurons depending on the activation of others in the same net.
+Thus, while ANNs rely on a global structuring of neuronal activations, the CNA relies on a local self-organizing process.
+Moreover, ANNs rely on hierarchical layers where neurons respond to features of increasing complexity.
+The CNA replaces this hierarchy with a single layer, where net fragments of different sizes capture features of varying complexity.
+These fragments interconnect to form larger nets that represent entire objects, naturally solving the binding problem, whereas ANNs do not encode explicit relationships between features, leading to ambiguities such as hallucinated feature conjunctions.
+Furthermore, input representation in the CNA emerges from statistical regularities in the input, making it independent of specific tasks.
+In contrast, ANNs bias feature extraction to serve predefined training objectives, which limits their adaptability when retraining for different functions.
 7 Discussion and Outlook
-The CNA presented here establishes a novel approach to construct coherent nets as neural representations of visual input, enabling arrays of elements to be treated as unified gestalts through the composability of previously learned sub-patterns (net fragments).
-While the concept of such net construction has been proposed theoretically (von der Malsburg et al., 2022), this work marks the first successful implementation and simulation of the idea.
-Our approach offers distinct advantages over existing neural coding frameworks, particularly in terms of robustness and the ability to represent out-of-domain data.
-Using nets of neurons to represent sensory signals allows filtering out noise and figure completion, while using net fragments allows representing unseen objects through composing small feature patches.
-In contrast to past work that relied on manually constructed filter functions (e.g., Geman and Geman, 1984), the filter functions, nets in our scheme, are learned from input statistics resp.
-constructed on the spot, so that they will eventually be able to cover the whole gamut of patterns encountered in natural images.
-An essential next step would be the construction of a system that is trained on natural images, that contains at each position of the visual field a complement of elementary feature detector neurons able to accurately represent the pixels of the input (as in Olshausen and Field, 1996), and whose recurrent connections have sufficient lateral extent to capture the most important signal correlations present in natural images.
-Such systems can then be expected to develop net fragments that capture and represent all textures encountered in natural images on that length scale, thus realizing the ambition expressed in Figure 1.
-A first and important such operation is abstracting over position, size, and in-plane orientation of an object’s image in the input area (primary visual cortex) by the construction of an invariant net representation of the object in a separate area (which would correspond to inferotemporal cortex (Ito et al., 1995)).
-As hinted at in the Introduction, this can be based on structure-sensitive, ‘homeomorphic’, mappings, as proposed in von der Malsburg (1981); Anderson and Van Essen (1987); Olshausen et al.
-(1995); Wiskott et al.
-(1997); Wolfrum et al.
-(2008); Memisevic and Hinton (2010) or Arathorn et al.
-(2013).
-A mapping is homeomorphic if it preserves neural feature type as well as connectivity.
-Object classification or recognition of individual objects can then be based on a further homeomorphic mapping to an area containing a more permanent schematic representation of the object class or a sample of the object to be recognized, as realized in Wolfrum et al.
-(2008).
-The introduction of nets as a neural code and the inclusion of alternative neurons suggest a natural basis for the mapping process by requiring inter-areal connections to be not only specific to feature type but also to alternative neuron identity.
-As this identity contains information on net connectivity, this specificity can implement the homeomorphy constraint of inter-areal mappings in a natural way and overcome the limitations in existing systems (Wolfrum et al., 2008).
-Once the mechanism of homeomorphic mapping has been established, it is available for arbitrary novel objects, thus forming the basis for invariant object recognition at first sight (as observed in humans in Biederman and Cooper (1991)).
-This capability stands in stark contrast to current ANN technology, which requires invariance to be trained individually for each object, highlighting the broader generalization potential of our approach.
-The simple mechanism of net formation that we model here is a first but significant step in the direction of a data architecture grounded in nets connecting sparse sets of neurons within and between areas.
-Nets are singled out by the properties of cooperativity and inner consistency: all neurons of an active net experience the high-order coincidence of converging excitatory signals, realizing predictive coding at each neuron.
-Given high sparsity of neural firing and high sparsity of connectivity, satisfying the requirement of consistency is highly non-trivial and can only be established by a learning process in which single neurons acquire converging connections coming from sets of neurons that fire simultaneously again and again, as realized in our model.
-In summary, the CNA demonstrates the feasibility and potential of a principled and compositional approach to neural coding.
-Its ability to autonomously construct robust representations, handle out-of-domain data, and generalize to novel stimuli underscores its promise as a foundation for more advanced systems in the future, especially when extended with multiple areas.
+The CNA presented here establishes a novel approach for constructing coherent nets as representations of input data.
+This framework enables the encoding of complex patterns through the composability of previously learned sub-patterns (net fragments).
+Compared to existing neural coding frameworks, CNA offers distinct advantages, particularly in terms of robustness – demonstrated through its ability to filter noise and reconstruct figures – and its capacity to generalize to out-of-domain data, as evidenced by its capability to represent previously unseen objects.
+For future research, we outline two key directions: First, an essential next step is the extension of CNA to natural image domains.
+This would allow net fragments to capture and encode texture information encountered in real-world scenarios, thus realizing the ambition expressed in Figure 1.
+Second, the current implementation of CNA is restricted to a single processing area with net fragments.
+Following the theoretical framework proposed by von der Malsburg et al.
+(2022), the architecture could be expanded to incorporate multiple areas, thereby enabling object recognition while achieving invariance to position, scale, and in-plane rotation.
+Object recognition could be realized using structure-sensitive, homeomorphic mappings between nets (von der Malsburg, 1981; Anderson and van Essen, 1987; Olshausen et al., 1995; Wiskott et al., 1997; Wolfrum et al., 2008; Memisevic and Hinton, 2010; Arathorn et al., 2013), which would facilitate the mapping of input representations onto a more stable schematic representation of an object class or an exemplar of the object to be recognized.
+The adoption of homeomorphic mappings, as opposed to alternative methods, preserves the core principle of nets and net fragments within and across multiple processing areas, thereby maintaining structural coherence in representation.
+The proposed neural coding framework not only advances representational robustness but also aligns with principles of self-organization observed in cortical processing.
+Its ability to learn and flexibly recombine net fragments without explicit supervision suggests a pathway toward more adaptive and interpretable neural representations.
+Future extensions to multi-area architectures could bridge the gap between bottom-up feature learning and top-down predictive coding.
+These developments position CNA as a foundational step toward more structured and generalizable neural computation.
 Funding
@@
 Appendix
-Appendix A Weight Initialization
-The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are copied into the $S2$ neurons in the same position.
-This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative neurons.
-The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$.
-When Hebbian plasticity is employed on the weight matrix $\boldsymbol{W}^{(L)}$, these connections ought to grow (as neurons tend to keep their activation state between time steps), but due to the constraint of keeping all connection weights within the range $(0,1)$ (refer to equation 8), they stay put at $1$.
-Appendix B Hebbian Plasticity
-After processing for $T$ time steps both matrices $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ undergo Hebbian updates (Hebb, 1949):
-$\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\alpha\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (8)
-where $\alpha$ represents the learning rate.
-Note that the network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant.
-Consequently, the same recurrent connection is applied across multiple output and input neurons.
-The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects.
-If the resulting average is positive, the corresponding connection increases by $\alpha$ as it connects more simultaneously active neurons than neurons that fire disjointly.
-Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced by $\alpha$.
-The actual modification eq.
-(8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
-Appendix C Normalization
-We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$.
+Appendix A Feature Extraction
+Figure 9: The hand-crafted convolutional kernels of the feature extractor.
+The stage $S1$ serves to extract initial binary features from the images.
+The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location.
+We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal).
+This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons).
+The use of such manually designed filters allows a more straightforward interpretation of the results.
+While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
+These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
+Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$.
+The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels.
+We convert the floating-point output of the convolutional weights to binary activations using the heaviside function $\Theta$ with a bias of $b^{(S1)}=0.5$, i.e.
+neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq.
+12).
+Appendix B Normalization
+The membrane potential in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{M}(S2)}$ to the interval $(0,1)$.
 We first apply a size limit:
-$\boldsymbol{a}^{\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
-According to this formula, $\boldsymbol{a}^{\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$.
-(In our experiments we found the value $\lambda=1.3\cdot\frac{h^{(W2)}+w^{(W2)}}{2}$ to work well).
+$\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (21)
+According to this formula, $\boldsymbol{a}^{\mathcal{M}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{M}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$.
+(In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well).
 This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons.
 In the next step, we introduce the normalization
-$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
-This normalization confines activations to the interval $(0,1)$, where neurons receiving maximal support are mapped to $1$, while those with lesser support are mapped to correspondingly smaller values.
-This normalization relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later, when recurrent links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
-Appendix D Alternative Neuron Selection
-Only one of a set of alternative neurons is allowed to be active.
-The selection process is determined by evaluating the correlation between feature patches and alternative kernel filters.
-Initially, patches matching the size of the feature kernel are extracted from every position in the input $(\boldsymbol{y}^{(S1)};\boldsymbol{y}^{(S2)}[t-1])$.
-Subsequently, the correlation of these patches with the feature kernels is computed, and at each location, the alternative kernel exhibiting the highest correlation is chosen.
-Consequently, multiple features may be active at each location, but only one of the competing feature channels — the one best matching a given input patch — achieves activation.
-At initialization (see Appendix A), where all alternative channels fit equally well to a given local pattern, the first channel is selected.
-However, after applying a Hebbian update, this channel starts specializing in recognizing a specific local pattern.
-Consequently, the activation of subsequent local patterns may either trigger the same channel if they have a high correlation or activate another channel better suited to recognizing the new pattern.
-Appendix E Dataset
-The dataset comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
+$\boldsymbol{a}^{\text{norm }(S2)}_{c,i,t}=\max\left(\frac{\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i,t}}{\max_{i^{\prime}}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i^{\prime},t}},0\right)$ (22)
+Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$.
+This normalization confines activations to the interval $[0,1]$, where neurons receiving maximal support are mapped to $1$, while those with less support are mapped to correspondingly smaller values.
+This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.
+Appendix C Single Weight-Kernel Implementation
+So far, we introduced the weights $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ in $S2$ as two independent convolutional kernels.
+For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}^{(S1)}$ and $\boldsymbol{y}^{(S2)}_{t-1}$, leading to the equivalent formulation of eq.
+16:
+$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(S2)}_{c_{\text{out}},c_{\text{in}}}\star\left(\boldsymbol{y}_{c_{\text{in}}}^{(S1)};\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)$ (23)
+Appendix D Dataset
+The dataset used for training comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
 The images are generated when required and may vary from one epoch to another.
@@
 To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
-To demonstrate the generalization capability of CNA areas, we generate validation samples that differ from the ones used during training.
+To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training.
 Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
@@
 Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
-Appendix F Feature Extraction
-Figure 10: The hand-crafted convolutional kernels of the feature extractor.
-The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location.
-We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 10, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal).
-This approach is motivated by the observation that, given our specific elementary feature types (cf.
-Figure 10), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons).
-The use of such manually designed filters allows a more straightforward interpretation of the results.
-While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
-These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
-Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$.
-The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels.
-We convert the floating-point output of the convolutional weights to binary activations using binary neurons $\mathrm{B}()$ with a bias of $b^{(S1)}=0.5$, i.e.
-neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq.
-2).
-Appendix G Parameters
-We train the model with a learning rate of $\alpha=0.2$ for $100$ training cycles (epochs).
-The model uses $n_{a}=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons.
-Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix F), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot n_{a}=40$.
+Appendix E Parameters
+We train the model with a learning rate of $\eta=0.2$ for $100$ training cycles (epochs).
+The model uses $\kappa=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons.
+Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$.
 The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
-Appendix H Autoencoder Parameter
+Appendix F Autoencoder Parameter
 We compare our CNA model to an ANN, specifically an autoencoder (Hinton and Salakhutdinov, 2006), which is a widely used self-supervised technique for generating input representations.
@@
 The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model.
-It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ layers.
+It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ channels.
 The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively.
@@
 The autoencoder is trained with a batch size of $512$.
-Appendix I Quantifying Noise Filtration
+Appendix G Quantifying Noise Filtration
 Our evaluations focus on the efficacy of net fragments in suppressing added Gaussian noise and complementing missing or occluded patterns.
@@
 First row: noise reduction rate; second row: recall; third row: precision.
-Columns differ in the inhibition coefficient $\gamma$ that models inhibition.
+Columns differ in the attenuation coefficient $\gamma$.
 The colors within each plot represent different activation biases $b^{(S2)}$.
@@
 Figure 10 displays noise reduction rates, recall, and precision across different parameter configurations for different levels of added noise.
-The metrics demonstrate that increasing the inhibition coefficient or the activation bias results in enhanced noise filtration.
+The metrics demonstrate that increasing the attenuation coefficient or the activation bias results in enhanced noise filtration.
 […]
@@
 First row: feature reconstruction rate; second row: recall; third row: precision.
-Columns are distinguished by the inhibition coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$.
+Columns are distinguished by the attenuation coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$.
 The dotted red line is the autoencoder model.
 Figure 11 displays the feature reconstruction rate, precision, and recall for removed line fragments of different lengths.
-Having a lower activation bias or inhibition coefficient $\gamma$ leads to a better reconstruction of removed patterns, as these settings allow neurons with less support to activate.
+Having a lower activation bias or attenuation coefficient $\gamma$ leads to a better reconstruction of removed patterns, as these settings allow neurons with less support to activate.
 Compared to the autoencoder baseline, the CNA model achieves higher recall and precision but a lower feature reconstruction rate.

### v4

@@
 Our findings establish CNA as a promising paradigm for developing neural representations that integrate local feature processing with global structure formation, providing a foundation for future research on invariant object recognition.
-Keywords net fragments  $\cdot$ nets  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ machine learning
+Keywords net fragments  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ representation learning
 1 Introduction
@@
 This limitation highlights a fundamental gap between current AI systems and the flexible, context-aware perception observed in biological vision.
-Existing approaches attempt to address these challenges through methods such as feature or sample engineering (Amirian et al., 2018; Moosavi-Dezfooli et al., 2016), domain adaptation (Csurka, 2017; Sager et al., 2022), and data augmentation (Carlucci et al., 2019; Tuggener et al., 2024).
+Existing approaches attempt to address these challenges through methods such as feature or sample engineering (Amirian et al., 2018; Moosavi-Dezfooli et al., 2016; Meyer et al., 2025), domain adaptation (Csurka, 2017; Sager et al., 2022), and data augmentation (Carlucci et al., 2019; Tuggener et al., 2024).
 However, these techniques largely operate within the conventional paradigm of representing objects as static patterns of neuronal activations, often requiring large datasets and extensive supervision to achieve generalization.
-In contrast, we propose a new paradigm by introducing the Cooperative Network Architecture (CNA), a framework in which object representations emerge from dynamic interactions among recurrently connected neurons.
+In contrast, we propose a new paradigm by introducing the Cooperative Network Architecture (CNA), a model in which object representations emerge from dynamic interactions among recurrently connected neurons.
 Rather than encoding objects as distinct activation vectors, CNA organizes sensory information into “nets” – structured networks of mutually supporting neurons.
@@
   • and an analysis of the emergent filtering and generalization properties of nets.
-Our results show that CNA offers a biologically inspired computational framework that integrates local and global information processing.
-It provides a foundation for more robust and adaptable perception systems, aligning with theoretical proposals for dynamic, network-based neural coding (von der Malsburg, 2018; von der Malsburg et al., 2022).
-By establishing the neural code and core computational mechanisms within a single representational area, this work provides the essential groundwork for future architectures in which multiple areas cooperate to support invariant object recognition – thus advancing toward the full potential envisioned in network-based neural coding models of perception.
+CNA offers the conceptual advance of a biologically inspired computational framework that integrates local and global information processing.
+Our results indicate that it provides a foundation for more robust and adaptable perception systems, aligning with theoretical proposals for dynamic, network-based neural coding (von der Malsburg, 2018; von der Malsburg et al., 2022).
+In this work, we establish the neural code and core computational mechanisms within a single representational area (a layer of neurons in which nets and their constituent fragments are formed, see below).
+Such a single area learns input-driven net fragments that bind together co-occurring sensory features, enabling the assembly of coherent perceptual nets.
+Future architectures will extend this principle to multiple interacting areas, in which additional areas can maintain object-associated net fragments that act as top-down hypotheses.
+Through reciprocal interactions between input-driven and object-memory areas, such systems would support invariant object recognition and more complex perceptual inference, realizing the broader vision of network-based neural coding models of perception (see our discussion in Section 6).
+The remainder of this paper is structured as follows.
+In Section 2, we present the theoretical foundations of our approach and provide the first formal mathematical definition of nets and net fragments, concepts that have previously been described only in prose (von der Malsburg, 2018; von der Malsburg et al., 2022).
+Section 3 introduces a concrete computational implementation of this framework.
+In Section 4, we report experimental results demonstrating the model’s ability to generalize to out-of-distribution inputs and to perform noise suppression.
+Section 5 situates our approach within the broader literature, contrasting it with related models such as associative memory networks, Boltzmann machines, and modern artificial neural networks.
+Finally, Section 6 discusses the implications of our findings and outlines how the proposed principles could be extended toward multi-area architectures for object recognition.
 2 The Net Fragment Framework
-Before discussing the computational framework, we first introduce its building blocks: neurons, activations, nets, and net fragments.
-These definitions provide the basis for the discussions in the subsequent sections.
-For simplicity, we limit the definitions to a single area even though it could be extended to multiple inter-connected areas.
+Before presenting the computational framework, we introduce its foundational components: neurons, activations, nets, net fragments, and competition.
+These theoretical definitions serve as the basis for the subsequent sections.
+Implementation details for the introduced formalisms are provided in the next section.
 Neurons.
 A neuron $i$ is denoted by $n_{i}$.
-Each neuron $n_{i}$ has a local neighborhood, denoted by $N_{r}(n_{i})$, which is the set of neurons $n_{j}$ such that the distance $d(n_{i},n_{j})$ between $n_{i}$ and $n_{j}$ is less than or equal to a threshold radius $r>0$, i.e.,
-$N_{r}\left(n_{i}\right)=\{n_{j}\mid d(n_{i},n_{j})\leq r\}.$ (1)
+Each neuron $n_{i}$ has a local neighborhood, denoted by the closed ball $\mathcal{B}_{r}(n_{i})$, which is the set of neurons $n_{j}$ such that the distance $d(n_{i},n_{j})$ between $n_{i}$ and $n_{j}$ is less than or equal to a threshold radius $r>0$, i.e.,
+$\mathcal{B}_{r}\left(n_{i}\right)=\left\{n_{j}\mid d(n_{i},n_{j})\leq r\right\}.$ (1)
 At each time step $t$, the activity state of a neuron $n_{i}$ is represented by $y_{i,t}\in\{0,1\}$, where $y_{i,t}=1$ indicates that the neuron is active at $t$.
@@
 In this work, we limit the connections to local neighborhoods:
-$\forall w_{i,j}\quad n_{j}\in N_{r}(n_{i})$ (2)
-Additionally, we limit the weights to a range between 00 and $1$, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
+$\forall w_{i,j}\quad n_{j}\in\mathcal{B}_{r}(n_{i})$ (2)
+Additionally, we limit the weights to a range between 0 and 111, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
 $w_{i,j}\in[0,1]$ (3)
@@
 A net fragment is defined as the set of neurons that provide direct activation support to a given neuron.
-In the context of this work, a neuron $n_{j}$ is said to provide support to another neuron $n_{i}$ if two conditions are met: (1) there exists a high correlation between two neurons $n_{j}$ and $n_{i}$ such that $n_{j}\sim n_{i}\Rightarrow w_{ji}>0$; and (2) neuron $n_{j}$ was active at the previous time step, i.e., $y_{j,t-1}=1$.
+A neuron $n_{j}$ is said to provide support to another neuron $n_{i}$ if two conditions are met: (1) there exists a high correlation between two neurons $n_{j}$ and $n_{i}$ such that $n_{j}\sim n_{i}\Rightarrow w_{ji}>0$; and (2) neuron $n_{j}$ was active at the previous time step, i.e., $y_{j,t-1}=1$.
 Since synaptic weights are bounded within the interval $w_{j,i}\in[0,1]$, the degree of support a neuron can provide is likewise restricted to this range.
@@
 The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment.
-We call this accumulated input the membrane potential at timestep $t$, which is a function $\mathcal{M}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
-$a^{\mathcal{M}}_{i,t}=\mathcal{M}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
-The membrane potential $a^{\mathcal{M}}_{i,t}\in\mathbb{R}$ is large when many other neurons provide support, meaning that they “confirm” the plausibility of a neuron’s activity by being active themselves and exhibiting a connection that has been strengthened through previous mutual activity and Hebbian updates.
-The activation of a neuron at timestep $t$ is a function of the membrane potential $a_{i,t}$ and environmental conditions $\mathcal{E}$:
-$y_{i,t}=\mathcal{A}\left(a_{i,t},\mathcal{E}\right)\in\{0,1\}$ (6)
-The environment conditions $\mathcal{E}$ include an attenuation signal, a constraint that selectively suppresses weak excitatory inputs by requiring neurons to accumulate a sufficiently large membrane potential before firing.
+We call this accumulated input the pre-synaptic activity at timestep $t$, which is a function $\mathcal{I}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
+$\boldsymbol{a}^{\mathcal{I}}_{i,t}=\mathcal{I}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
+The pre-synaptic activity $a^{\mathcal{I}}_{i,t}\in\mathbb{R}$ is large when many other neurons provide support.
+Since supporting neurons exhibit connections that have been selectively strengthened through previous mutual activity and Hebbian updates within specific contexts, their activity effectively serves to confirm the contextual plausibility of a given neuron’s activation.
+The very general formulation of $\mathcal{I}$ shows that the pre-synaptic activity can be calculated with any desired function.
+In this work, we specifically define $\mathcal{I}$ as the weighted sum of prior activations from supporting neurons:
+$\boldsymbol{a}^{\mathcal{I}}_{i,t}=\sum_{n_{j}\in\mathcal{F}_{t}(n_{i})}\boldsymbol{w}_{j,i}\cdot\boldsymbol{y}_{j,t-1}.$ (6)
+The activation of a neuron at timestep $t$ is determined by the activation function $\mathcal{A}$ that depends on the pre-synaptic activity $a_{i,t}$ and environmental conditions $\mathcal{E}$:
+$\boldsymbol{y}_{i,t}=\mathcal{A}\left(\boldsymbol{a}_{i,t},\mathcal{E}\right)\in\{0,1\}$ (7)
+The environmental conditions $\mathcal{E}$ include an attenuation signal, a constraint that selectively suppresses weak excitatory inputs by requiring neurons to accumulate a sufficiently large pre-synaptic activity before firing.
+Again, we here provide a very general formulation, using $\mathcal{A}$ and $\mathcal{E}$, to emphasize that any activation function or environmental mechanism can be used.
+In this work, we instantiate $\mathcal{E}$ using a combination of normalization (see Appendix B) and attenuation (see Eq.
+28), and define $\mathcal{A}$ as the heaviside step function (see Eq.
+29).
 Nets.
@@
 A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects.
-For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active:
+For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active as well:
 $\mathcal{F}^{*}_{t}(n_{i})=\mathcal{F}_{t}(n_{i})\quad\text{with}\quad y_{i,t}=1$ (8)
 To account for both direct and indirect influences on a neuron’s activation, we use a recursive formulation of nets from the perspective of a single neuron.
-The base case is defined as the immediate net fragment:
+In the base case, a net $\mathcal{N}_{t}^{1}[n_{i}]$ is defined as the immediate net fragment:
 $\mathcal{N}_{t}^{1}[n_{i}]=\mathcal{F}^{*}_{t}(n_{i})$ (9)
@@
 which comprises all neurons in the active subgraph that are reachable from $n_{i}$.
-This recursive formulation illustrates how local net fragments, which can be interpreted as elementary features, are progressively integrated to form higher-order features.
-Moreover, the explicit time dependence in these definitions underscores the dynamic nature of both the net fragments and the overall network, as they are continuously reassembled in response to the evolving support received from neighboring neurons, and increasing attenuation that prevents neurons with a low membrane potential from firing (see next section).
+This recursive formulation illustrates how support relationships propagate through time: local net fragments $\mathcal{F}_{t}(n_{i})$, which reflect immediate sources of support for a neuron, are combined recursively to form increasingly larger subnetworks $\mathcal{N}_{t}^{r}[n_{i}]$.
+These progressively include neurons with indirect but recurrent influence on $n_{i}$’s activation.
+This process reflects how localized evidence (elementary activations) can aggregate into a coherent larger pattern (composite activation), corresponding to a higher-order feature representation in the network.
+Moreover, the explicit time dependence in these definitions underscores the dynamic nature of both the net fragments and the overall net, as they are continuously reassembled in response to the evolving support received from neighboring neurons, and increasing attenuation that prevents neurons with a low pre-synaptic activity from firing (see next section).
 In the proposed architecture, no symmetry constraint is imposed on the weights, that is, $w_{i,j}\neq w_{j,i}$ in general.
 However, due to the temporal dynamics of the network, activations tend to stabilize over a longer time horizon $T$, such that for sufficiently large $t$, the activations become approximately stationary, i.e., $y_{i,t}\approx y_{i,t+1}$.
-Under these conditions, mutual influence between neurons becomes symmetric, meaning that $n_{i}\sim n_{j}$ implies $n_{j}\sim n_{i}$, and almost symmetric weights emerge naturally over time, even in the absence of explicit symmetry constraints.
+Under these conditions, the mutual influence between neurons becomes symmetric, meaning that $n_{i}\sim n_{j}$ implies $n_{j}\sim n_{i}$, and almost symmetric weights emerge naturally over time, even in the absence of explicit symmetry constraints.
 As a result, it can be assumed that for any pair of distinct neurons $n_{i}\neq n_{j}$, their closures either converge to the same set of neurons or remain largely disjoint.
@@
 $\mathcal{N}_{t}^{*}[n_{i}]\approx\mathcal{N}_{t}^{*}[n_{j}]\quad\text{or}\quad\mathcal{N}_{t}^{*}[n_{i}]\cap\mathcal{N}_{t}^{*}[n_{j}]\approx\emptyset.$ (12)
+Competitive Neurons.
+One issue with the current formulation is that the mapping
+$\mathcal{A}\circ\mathcal{I}:2^{\mathcal{N}}\;\longrightarrow\;\{0,1\}$ (13)
+treats any two net fragments that yield the same activation identically. In particular, we say
+$\mathcal{F}_{t_{k}}(n_{i})\simeq\mathcal{F}_{t_{l}}(n_{i})\quad\Longleftrightarrow\quad\mathcal{A}\bigl{(}\mathcal{I}(\mathcal{F}_{t_{k}}(n_{i})),\mathcal{E}\bigr{)}=\mathcal{A}\bigl{(}\mathcal{I}(\mathcal{F}_{t_{l}}(n_{i})),\mathcal{E}\bigr{)},$ (14)
+meaning different contexts collapse to the same output for $n_{i}$. To break this symmetry and make each neuron context-aware, we introduce competitive neurons.
+For each neuron $n_{i}$, We introduce a set of alternatives $\{n_{i}^{1},\dots,n_{i}^{\kappa}\}$, all located at the same spatial position $\left(\mathcal{B}_{r}(n_{i})=\mathcal{B}_{r}(n_{i}^{k})\right)$.
+These alternatives undergo competition and one of them is permitted to be active within a given context.
+This exclusivity encourages these competitive neurons to specialize, over time, into representations of different contextual variations.
+To formalize these competitive neurons, we first define the competition class of a neuron $n_{i}$:
+$\mathcal{C}(n_{i})=\bigl{\{}n_{i}^{k}\,\text{ where }\,\mathcal{B}_{r}(n_{i})=\mathcal{B}_{r}(n_{i}^{k})\bigr{\}},$ (15)
+These neurons are mutually exclusive candidates to represent the context-dependent expression of a feature at a fixed spatial position.
+During each time step, after computing the pre-synaptic activities $\boldsymbol{a}_{j,t}^{\mathcal{I}}$ for $n_{j}\in\mathcal{C}(n_{i})$, we select a small winner set
+$W_{t}(n_{i})=\mathrm{Comp}\bigl{(}\{\,\boldsymbol{a}_{j,t}^{\mathcal{I}}\mid n_{j}\in\mathcal{C}(n_{i})\}\bigr{)}.$ (16)
+While various rules to determine the winning set can be used, we apply winner-takes-all selection in this work (for implementation details, see Section 3.2).
+Only neurons in $W_{t}(n_{i})$ are eligible to fire:
+$y_{i,t}\;=\;\begin{cases}\mathcal{A}\bigl{(}\boldsymbol{a}_{i,t}^{\mathcal{I}},\mathcal{E}\bigr{)}&\text{if }n_{i}\in W_{t}(n_{i}),\\ 0&\text{otherwise.}\end{cases}$ (17)
+By enforcing mutual exclusivity among neurons with overlapping receptive fields, competition ensures that different contexts activate distinct neuronal subsets.
 3 Cooperative Network Architecture
-Figure 2: The dynamics within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations.
+Figure 2: Dynamics within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations.
 The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer.
@@
 The CNA comprises two processing stages, depicted as yellow squares in Figure 2.
-The stage 1 ($S1$) performs fixed (i.e., not learned) feature extraction directly from the image input.
-The stage 2 ($S2$) implements the core principles of the described computational framework by learning net fragments.
+Stage 1 ($S1$) performs fixed (i.e., not learned) feature extraction directly from the image input.
+This fixed feature stage is used here for simplicity and to facilitate clearer diagnostics of the subsequent results by combining well-characterized, predefined features.
+However, these features could also be learned in future work, for example, through biologically inspired Hebbian learning schemes that have been shown to induce convolutional, Gabor-like filters (e.g., Miconi, 2021).
+Stage 2 ($S2$) implements the core principles of the described computational framework by learning net fragments.
 The network dynamics unfold over discrete time steps $t=[0,\dots,T]$.
-Since the feature extraction stage ($S1$) employs fixed filters, it yields a time-invariant activation pattern for a given input.
+Since the feature extraction stage ($S1$) employs four fixed filters, it yields a time-invariant activation pattern for a given (fixed) input.
 In contrast, the second stage ($S2$) exhibits dynamic temporal evolution, where internal representations are iteratively refined at each time step ($t\to t+1$).
-Specifically, we use a progressively increasing global attenuation signal that only allows neurons with a high membrane potential to remain active.
-During this dynamic processing, a neuron might be deactivated, whereby another neuron’s membrane potential is decreased, triggering a chain reaction that can lead to the deactivation of incoherent nets.
-While these refinements occur within the processing cycle of a single sample ($[0,\dots,T]$), Hebbian learning adapts the weights (connections $w_{i,j}$ within net fragments) on longer timescales, wherein synaptic weight modifications occur between different sample-processing episodes, leading to structural plasticity in the network.
+Specifically, we use a progressively increasing global attenuation signal that only allows neurons with a high pre-synaptic activity to remain active.
 The following sections discuss these two processing stages.
@@
 Consequently, this stage serves as a preprocessing step, enabling the transformation of continuous-valued images into a suitable binary representation.
-The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels ($1$ for grayscale, $3$ for RGB), while $H,W$ stand for the height and width of the image.
+The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels (111 for grayscale, 333 for RGB), while $H,W$ stand for the height and width of the image.
 Activation Computation.
-To extract initial activation patterns, we use fixed, non-trainable filters (details see Appendix A).
+To extract initial activation patterns, we use four fixed, non-trainable filters (details see Appendix A).
 These filters are implemented as standard convolutional kernels, used to obtain the binary activation pattern $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)},H,W}$ as follows:
@@
 To binarize the activations, we use the heaviside function $\Theta$, which sets a neuron to one if its input is larger than zero, and to zero otherwise.
-$y_{i}=\Theta(a_{i})=\begin{cases}1\quad\text{if }a_{i}>0\\ 0\quad\text{otherwise}\end{cases}$ (13)
+$y_{i}=\Theta(\boldsymbol{a}_{i})=\begin{cases}1\quad\text{if }\boldsymbol{a}_{i}>0\\ 0\quad\text{otherwise}\end{cases}$ (19)
 Note that the width and height between input and output activations remain identical, but the number of channels changes from $C$ to $C^{(S1)}$, allowing the extraction of multiple features at each spatial location.
 3.2 Stage 2: Net Fragments
-$S2$ constructs net fragments from the binary feature activations $\boldsymbol{y}^{(S1)}$ of $S1$ and integrates them into coherent, spatially extended nets over time.
+$S2$ constructs net fragments from the binary feature activations $\boldsymbol{y}^{(S1)}$ of $S1$ and integrates them into coherent nets over time.
 To ensure stability, single neurons or small groups of neurons are suppressed through an attenuation mechanism, ensuring that only large-scale, consistent activation patterns persist.
@@
 $d(n_{i},n_{j})=\max\left(|h_{i}-h_{j}|,|w_{i}-w_{j}|\right)$ (20)
-Each neuron $n^{(S2)}_{i}$ in stage 2 is connected in two ways: (1) Feedforward connections to the corresponding neuron $n^{(S1)}_{i}$ in stage 1 and to its local neighborhood $Nr(n^{(S1)}_{i})$, and (2) lateral connections to neighboring neurons within stage 2 itself, denoted as:
-$Nr\left(n^{(S2)}_{i}\right)=Nr\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r_{2}\right\}$ (15)
-where $r_{2}$ defines the radius of lateral connections.
+Each neuron $n^{(S2)}_{i}$ in stage 2 is connected in two ways: (1) Feedforward connections to the corresponding local neighborhood $\mathcal{B}_{r}(n^{(S1)}_{i})$ in $S1$, and (2) lateral connections to neighboring neurons within stage 2 itself, $\mathcal{B}_{r}(n^{(S2)}_{i})$, denoted as:
+$\mathcal{B}_{r}\left(n^{(S2)}_{i}\right)$ $=\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r\right\}$ (21)
+$\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)$ =$=\left\{n_{j}^{(S1)}\mid d(n_{i}^{(S2)},n_{j}^{(S1)})\leq r\right\}$
+where $r$ defines the radius of lateral connections.
 The convolutional architecture realizes these neighborhood structures via convolutional filters, where the filter size directly determines the radius of both feedforward and lateral connections.
@@
 Specifically, the lateral connections $\boldsymbol{W}^{(L)}$ calculate the state updates of $\boldsymbol{y}^{(S2)}_{t\rightarrow t+1}$ within an iterative refinement process in which individual neuronal activations ${y}_{i}^{(S2)}\in\boldsymbol{y}^{(S2)}$ may flip due to changes in support and attenuation.
-Membrane Potential.
-To determine the support a single neuron $n_{i,t}$ receives from the net fragment $\mathcal{F}_{t-1}(n_{i,t-1})$, we first calculate the membrane potential $\mathcal{M}$ by applying convolutional operations on the forward and lateral connections and then taking the sum of these operations:
-$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\underbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S1)}}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}}}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)}^{\textit{Recurrent Connection}}.$ (16)
+Pre-Synaptic Activity.
+To determine the support a single neuron $n_{i,t}$ receives from the net fragment $\mathcal{F}_{t-1}(n_{i,t-1})$, we first calculate the pre-synaptic activity $\mathcal{I}$ by applying convolutional operations on the forward and lateral connections and then taking the sum of these operations:
+$\boldsymbol{a}^{\mathcal{I}(S2)}_{c_{\text{out}},t}$ $=\underbrace{\sum_{c_{\text{in}}=0}^{C^{(S1)}}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}}}^{(S1)}}_{\textit{forward input }(\boldsymbol{a}^{\mathcal{I}(F)})}+\underbrace{\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}}_{\textit{lateral input }(\boldsymbol{a}^{\mathcal{I}(L)})}$ (22)
+$\boldsymbol{a}^{\mathcal{I}(F)}_{c_{\text{out}}}$ $=\sum_{c_{\text{in}}=0}^{C^{(S1)}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}},h,w}\cdot\boldsymbol{y}^{(S1)}_{c_{\text{in}},h^{\prime}+h,w^{\prime}+w}$
+$\boldsymbol{a}^{\mathcal{I}(L)}_{c_{\text{out}},t}$ $=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}},h,w}\cdot\boldsymbol{y}^{(S2)}_{c_{\text{in}},h^{\prime}+h,w^{\prime}+w,t-1}$
 Here, the forward connections have the shape $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},C^{(S1)},h,w}$ and the recurrent connections have the shape $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},C^{(S2)},h,w}$, where $C^{(S2)}$ is the number of channels in $S2$.
-Therefore, the membrane potential $\mathcal{M}$ at time $t$ is increased by the learned amount $w_{j,i}$ from all connected and active neurons ($y_{j,t-1}=1$).
-Note that $\mathcal{F}_{0}(n_{i,0})$ only consists of neurons of $S1$ since all neurons in $S2$ are initialized to silent state ($y_{i,0}=0$).
-Higher support, indicated by larger $a^{\mathcal{M}(S2)}_{i}$, suggests a more likely pattern, as neurons encoding frequent patterns gain reinforcement not only from feedforward connections but also from lateral connections.
-Before we discuss how the weight matrices $\boldsymbol{W}^{(F)},\boldsymbol{W}^{(L)}$ are initialized and trained, we describe how the membrane potential $\boldsymbol{a}^{\mathcal{M}(S2)}$ undergoes attenuation and is transformed into $\boldsymbol{y}^{(S2)}$.
+Therefore, the pre-synaptic activity $\mathcal{I}$ at time $t$ is increased by the learned amount $\boldsymbol{w}_{j,i}$ from all connected and active neurons ($\boldsymbol{y}_{j,t-1}=1$).
+Higher support, indicated by larger $\boldsymbol{a}^{\mathcal{I}(S2)}_{i}$, suggests a more likely pattern, as neurons encoding frequent patterns gain reinforcement not only from feedforward connections but also from lateral connections.
+Next, we extend this convolutional-based formulation of forward and lateral connections by introducing competitive neurons.
+Afterward, we describe how the pre-synaptic activity $\boldsymbol{a}^{\mathcal{I}(S2)}$ undergoes attenuation, how it is transformed into $\boldsymbol{y}^{(S2)}$, and how the weight matrices $\boldsymbol{W}^{(F)},\boldsymbol{W}^{(L)}$ are initialized and trained.
+Competitive Neurons.
+To address the challenge of context entanglement, wherein individual neurons may participate in multiple fragments, we implement competitive neurons as follows: For each base feature channel $c\in C^{(S2)}$, we introduce $\kappa$ parallel alternatives that share the same receptive field structure but are allowed to develop distinct connections.
+Following our formal definition, each base neuron $n_{c,h^{\prime},w^{\prime}}$ at spatial location $(h^{\prime},w^{\prime})$ has the competitive class:
+$\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)=\left\{n_{c,h^{\prime},w^{\prime}}^{1},\dots,n_{c,h^{\prime},w^{\prime}}^{\kappa}\right\}$ (23)
+We implement this by introducing an additional competition dimension of size $\kappa$ to the tensors in $S2$:
+$C^{(S2)}$ $\to(C^{(S2)},\,\kappa)$ (24)
+$\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},\,C^{(S1)},\,h,\,w}$ $\to\mathbb{R}^{C^{(S2)},\,\kappa,\,C^{(S1)},\,h,\,w}$
+$\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},\,C^{(S2)},\,h,\,w}$ $\to\mathbb{R}^{C^{(S2)},\,\kappa,\,C^{(S2)},\,h,\,w}$
+$\boldsymbol{a}^{\mathcal{I}(L)}\in\mathbb{R}^{C^{(S2)},H,W}$ $\to\mathbb{R}^{C^{(S2)},\,\kappa,\,H,W}$
+Although this transformation increases the number of total feature maps by a factor of $\kappa$, only a single competitive neuron per base channel is active at any given spatial location (one neuron within $\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)$).
+This sparsity is enforced through a group-wise winner-take-all (WTA) mechanism along the $\kappa$-dimension.
+At each time step $t$, competition operates independently within each class $\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)$.
+For each base neuron, we compute pre-activation scores for all competitive neurons and select the winning competitive index:
+$k^{*}(c,h^{\prime},w^{\prime})\;=\;\operatorname*{arg\,max}_{k\in\{0,\dots,\kappa-1\}}\;\left\langle\boldsymbol{W}_{c,k,:,:,:,:},\;\boldsymbol{z}_{h^{\prime},w^{\prime},t}\right\rangle.$ (25)
+where $\boldsymbol{z}_{h^{\prime},w^{\prime},t}$ is the concatenated input patch:
+$\boldsymbol{z}_{h^{\prime},w^{\prime},t}\;=\;\begin{bmatrix}\boldsymbol{y}^{(S1)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W}\\ \boldsymbol{y}^{(S2)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W,\,t-1}\end{bmatrix}.$ (26)
+Thus, at each spatial location and time step, the index $k^{*}\in\{1,\dots,\kappa\}$ maximizing the correlation between the corresponding filter and the local input patch is chosen.
+Once the winning index $k^{*}$ is selected, the activation is computed exclusively for the corresponding neuron:
+$\boldsymbol{a}^{\mathcal{I}(S2)}_{c,h^{\prime},w^{\prime},t}$ $=\boldsymbol{a}^{\mathcal{I}(F)}_{c,k^{*},h^{\prime},w^{\prime}}+\boldsymbol{a}^{\mathcal{I}(L)}_{c,k^{*},h^{\prime},w^{\prime},t}$ (27)
+$\boldsymbol{a}^{\mathcal{I}(F)}_{c,k^{*},h^{\prime},w^{\prime}}$ $=\sum_{c_{\text{in}}=0}^{C^{(S1)}-1}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(F)}_{c,k^{*},\,c_{\text{in}},\,h,\,w}\cdot\boldsymbol{y}^{(S1)}_{c_{\text{in}},\,h^{\prime}+h,\,w^{\prime}+w}$
+$\boldsymbol{a}^{\mathcal{I}(L)}_{c,k^{*},h^{\prime},w^{\prime},t}$ $=\sum_{c_{\text{in}}=0}^{C^{(S2)}-1}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(L)}_{c,k^{*},\,c_{\text{in}},\,h,\,w}\cdot\boldsymbol{y}^{(S2)}_{c_{\text{in}},\,h^{\prime}+h,\,w^{\prime}+w,\,t-1}$
+Initially, all competitive neurons $n_{c,h^{\prime},w^{\prime}}^{k}$ are identical due to shared initialization.
+However, Hebbian learning induces specialization: each competitive neuron within $\mathcal{C}\left(n_{c,h^{\prime},w^{\prime}}\right)$ gradually diverges, learning to encode different contextual variations of the same base feature.
+In a practical implementation, we merge the $\kappa$ and channel dimensions, resulting in a flat channel axis of size $C^{(S2)}\cdot\kappa$.
+This allows the use of efficient standard 3D convolutional layers, where non-winning neurons are simply masked out at each spatial location by setting them to 0, preserving the theoretical definition.
 Attenuation.
-The membrane potential is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{M}},\mathcal{E}\right)$.
-In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{M}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range $[0,1]$ (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active.
+The pre-synaptic activity is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{I}},\mathcal{E}\right)$.
+In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{I}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range [0,1][0,1][ 0 , 1 ] (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active.
 Specifically, attenuation is implemented as:
@@
 The activation at the last timestep $\boldsymbol{y}^{(S2)}_{t=T}$ represents a recurrently refined activation pattern where unsupported neurons have been suppressed, yielding a coherent representation.
-Alternative Neurons.
-A single neuron that encodes a specific feature can be activated in multiple contexts.
-For example, a short straight-line segment may appear in various contexts – such as in different line drawings or digits.
-In our framework, each context is represented by a net, whereas individual features correspond to the activity of specific neurons.
-To achieve context disentanglement, it is important to allocate sufficient representational capacity such that the same feature (e.g., a line segment) can be utilized across distinct contexts (e.g., different line drawings) without interference.
-This requires that a feature-representing neuron is capable of participating in multiple, context-specific net fragments.
-Without this capacity, a neuron would be forced to average over competing contextual associations, resulting in a non-specific or “blurry” representation.
-Formally, a neuron representing a feature may be part of multiple net fragments: $n_{i}\in\mathcal{F}(n_{j})$ and $n_{i}\in\mathcal{F}(n_{k})$ where $n_{j}\neq n_{k}$).
-This implies that the neuron must form distinct lateral connections $w_{i,j}$ and $w_{i,k}$, which should not be concurrently active.
-To address this requirement, we introduce “alternative neurons”.
-Each alternative version of a neuron is activated by the same feature from $S1$ but has a distinct set of lateral connections in $S2$ to encode different contexts.
-This is implemented by increasing the channel dimension $C^{(S2)}$ of $\boldsymbol{W}^{(L)}$ by a factor $\kappa$, i.e.,
-$C^{(S2)}:=C^{(S2)}\cdot\kappa$ (19)
-This expansion provides capacity for $\kappa$ alternative version of a neuron.
-Consequently, these alternative neurons receive identical bottom-up input from $S1$ while differing in the lateral support they receive in $S2$.
-A winner-take-all (WTA) competition mechanism ensures that each alternative neuron specializes in a distinct context, with only one being activated in any given instance.
-[…]
-At initialization, neuron selection is random.
-However, after Hebbian learning updates, the alternative neurons progressively specialize, leading to distinct and structured fragment representations.
-Over time, the most correlated alternative neuron is chosen for each input location, ensuring robust feature specialization across different contexts.
 Weight Initialization.
-The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all alternative $S2$ neurons in the same position.
-This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to $1$ if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of alternative neurons.
-The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to $1$.
+The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all competitive $S2$ neurons in the same position.
+This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to 111 if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of competitive neurons.
+The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to 111.
 Learning of Weights.
-Support within net fragments is learned via Hebbian plasticity (Hebb, 1949), so that frequently occurring patterns that lead to correlated activation patterns between neurons $n_{i}\sim n_{j}$, are “stored” in the system and receive more support:
+Support within net fragments is learned via Hebbian plasticity, so that frequently occurring patterns that lead to correlated activation patterns between neurons $n_{i}\sim n_{j}$ are “stored” in the system and receive more support:
 $\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\eta\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (30)
@@
 Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced.
-Learning is applied after processing a sample (after $T$ timesteps) to improve stability.
-4 Experiments
-To demonstrate the compositionality of net fragments into nets, we train the model on binary images containing straight-line patterns so that the CNA learns net fragments corresponding to straight-line sub-patterns (smaller straight-line segments).
-However, for evaluation (after training and without weight updates), we use much more complex images such as line drawings of objects, and demonstrate that these complex patterns can be expressed as compositions (nets) of straight-line net fragments.
-This flexible compositional behavior contrasts strongly with conventional associative memory that cannot model such out-of-distribution data (see Section 6.1).
-To highlight the differences to ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
-We also examine robustness by introducing noise.
-Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects.
-Further experimental details are provided in the appendix: We describe the training and testing datasets in more detail in Appendix D, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix A, the training details of our CNA model are in Appendix E, and the training details of the autoencoder in Appendix F.
-5 Results
+Importantly, the weight update is applied only once per sample, after the full sequence of $T$ timesteps has been processed.
+Learning proceeds over multiple epochs, meaning each sample is revisited multiple times.
+Due to the small learning rate of $\eta=0.2$, these weight changes accumulate gradually.
+4 Experimental Results
+4.1 Experimental Setup
+Our experiments are designed to rigorously probe the core representational capability of the CNA: its ability to form net fragments during training and recombine them into larger nets during inference.
+To isolate this property, we deliberately avoid task-specific or dataset-driven complexity and instead focus on the model’s behavior in its most fundamental setting.
+We therefore train CNA on binary images of straight lines at varying orientations.
+Straight lines serve as the most elementary visual primitives: they are simple enough to avoid confounds yet naturally compose into richer structures such as line drawings of objects.
+This controlled regime directly tests whether CNA can learn fragmentary representations and later reassemble them into unseen global patterns.
+By contrast, benchmarks such as MNIST Lecun et al.
+(1998), CIFAR Krizhevsky and Hinton (2009), or ImageNet Deng et al.
+(2009) involve dense, texture-heavy images where the link between local features and global structure is opaque.
+Evaluating CNA on such data at this stage would obscure its core representational mechanism, and, as we discuss in Section 6, would require introducing an additional object-representing area and more advanced filters, which we defer to future work.
+Evaluation targets two key properties:
+  • Compositionality: Can learned line fragments recombine into novel structures, such as object line drawings unseen during training?
+  • Robustness: Can nets be retrieved reliably under noise or occlusion?
+To this end, we present the model after training and without updating any parameters with (i) line drawings of objects composed of straight-line segments, (ii) inputs corrupted with Gaussian noise, and (iii) partially occluded patterns.
+For comparison, we also train a conventional autoencoder on the same line-stimuli dataset to highlight representational differences.
+Further experimental details are provided in the appendix: We describe the training and testing datasets in more detail in Appendix D, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ in Appendix A, the training details of our CNA model in Appendix E, and the training details of the autoencoder in Appendix F.
+4.2 Interpreting Figures: General Considerations
 Figure 3: The response of our model to a diagonal line input is displayed as a net of activated neurons.
@@
 For convenience, the neurons are rearranged in a circle, and the net fragment (connections) of two randomly selected neurons, $n_{i}$ and $n_{j}$, are visualized in colors (purple and red).
-Although the figures illustrate individual neuron activity, it is important to clarify that our objective is not to interpret patterns merely as isolated neuron activations, but rather as nets – cohesive, interconnected clusters of active neurons.
-To emphasize this perspective, Figure 3 illustrates the connectivity among activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged in a circular layout for clarity), after feeding a diagonal line into the network.
-We select randomly two neurons $n_{i}$ and $n_{j}$ in the left panel and visualize their net fragments $\mathcal{F}(n_{i})$ and $\mathcal{F}(n_{j})$ based on $\boldsymbol{W}^{(L)}$ on the right.
-This figure highlights that individual neurons (e.g., neurons $n_{i}$ and $n_{j}$) are embedded within broader nets of supportive connections.
-Without such network-based support, these neurons would be suppressed via attenuation mechanisms and would not appear in the left panel of Figure 3, which displays neuronal activity.
-Consequently, the subsequent visualizations should be understood as representing supported neurons embedded within nets, even though the underlying connectivity is not explicitly shown.
+A defining feature of CNA is that it represents patterns not through individual neuron activations, but through nets – cohesive, interconnected clusters of mutually supportive neurons.
+This network-based organization means that an individual neuron’s activation is meaningful only in the context of its embedding within a larger net.
+Understanding CNAs’ outputs, therefore, requires a shift in perspective: from isolated neuron activity to the connectivity and structure of the nets they form.
+Figure 3 illustrates this principle.
+After presenting a diagonal line input, the left panel shows the resulting neuronal activity $\boldsymbol{y}^{(S2)}$, while the right panel visualizes the underlying net structure derived from recurrent connections $\boldsymbol{W}^{(L)}$.
+For clarity, neurons are arranged in a circular layout, and the net fragments $\mathcal{F}(n_{i})$ and $\mathcal{F}(n_{j})$ of two randomly selected neurons $n_{i}$ and $n_{j}$ are highlighted in color.
+This visualization demonstrates two important points.
+First, individual neurons (e.g., $n_{i}$ and $n_{j}$) remain active only if embedded within supportive nets: without sufficient recurrent support, they are suppressed by attenuation mechanisms and would appear inactive (black) in the left panel.
+Second, this shows that observed activity patterns should always be interpreted as the visible footprint of underlying nets, even when those nets are not explicitly drawn.
+Consequently, all subsequent visualizations of neural activity must be understood in this net-based context: the squares represent neurons that survive attenuation precisely because they are supported by their respective nets.
+All our visualizations of neural activity are color-coded according to the competing channel groups $\mathcal{C}(n_{i})$.
+Each color corresponds to a feature type (e.g., horizontal, vertical, or diagonal lines; see Appendix A).
+However, these colors do not distinguish between different competitive channels $k\in{1,\dots,\kappa}$ within the same feature group, meaning that two neurons with the same color may belong to different competitive subnetworks with distinct connectivity patterns.
 4.3 Compositionality
-Figure 4: The top of the figure illustrates the straight lines used during the training phase.
-Below, various samples from the evaluation datasets are shown, including kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows $5$-$8$).
+Figure 4: Training and evaluation data.
+The top of the figure illustrates the straight lines used during the training phase.
+Below, various samples from the evaluation datasets are shown, including kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888).
 For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
-The model weights are trained using straight lines (see training data at the top of Figure 4).
-During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 1-8 of Figure 4).
-Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 1 and 2), digits and characters (rows 3 and 4), and line drawings (rows 5–8).
+The model weights are trained using straight lines (see training data examples at the top of Figure 4).
+During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 111888 of Figure 4).
+Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888).
 These results demonstrate that while individual net fragments $\mathcal{F}_{t}$ must be learned (in our case we use Hebbian learning to train the connections within $\mathcal{F}_{t}$), the recursive composition described in Eq.
@@
 Consequently, the model exhibits combinatorial generalization through the structural recomposition of learned primitives.
+4.4 Filtering Gaussian Noise
 The motivation for why we construct net fragments rather than just using the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise.
-This noise corresponds to invalid patterns not observed during training (i.e., are not captured in net fragments), which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons.
+This noise corresponds to invalid patterns not observed during training (i.e., are not captured in net fragments) that consequently lack sufficient recurrent support, leading to the deactivation of the corresponding neurons.
 Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
-5.2 Filtering Gaussian Noise
-Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
-To further investigate noise filtering, we introduce Gaussian noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix D) to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
-The results are visualized in Figure 5: While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated.
-As additional noise is introduced, more undesired neurons receive support and persist in their activity.
-However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$.
-In Appendix G, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
-Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed.
-However, when attenuation increases over time, even more noise is filtered out.
-The plot is created after adding $10\%$ noise to each feature channel (using a bias $b^{(S2)}=0.7$).
-The reason for the filtering is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time.
+Figure 5: Noise filtering efficacy: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
+[…]
+[…]
+[…]
+The reason that the filtering works is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time.
 Figure 6 depicts this systematic filtering across timesteps.
@@
 4.5 Reconstructing Figures
-Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations.
+Figure 7: Recovery from occlusion.
+Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations.
 Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$.
-Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
-If a fragmet $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$).
+Part B: A diagonal line with 666 missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the pre-synaptic activity $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
+If a net fragment $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$).
 Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns.
-This reconstruction is reliable for up to $3$ missing pixels (see Appendix G).
-In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
-Membrane Potential Represents Expected Feature Activation.
-In the presence of missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected (as shown in Figure 7B).
+This reconstruction is reliable for up to 333 missing pixels (see Appendix G).
+In some cases, as in this figure, the line is fully reconstructed for up to 666 removed pixels and partly reconstructed for 777 pixels.
+Pre-Synaptic Activity Represents Expected Feature Activation.
+In the presence of missing pattern elements, the pre-synaptic activity map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected (as shown in Figure 7B).
 Depending on the context, it may be desirable either to reconstruct the missing line segment – effectively bridging the gap between discontinuous features – or to preserve the separation between distinct line elements.
 This behavior can be modulated by adjusting the hyperparameter $b^{(S2)}$ in Eq.
-18, which can govern the balance between feature completion and preservation of discontinuity, thereby allowing the model to selectively favor line reconstruction or separation.
-5.4 Comparison to an Autoencoder
+29, which can govern the balance between feature completion and preservation of discontinuity, thereby allowing the model to selectively favor line reconstruction or separation.
+In future work, incorporating additional areas that maintain object-associated representations and act as top-down hypotheses could enable this selectivity to be learned rather than manually tuned.
+4.6 Comparison to Autoencoders
 Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder.
 Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder).
-Rows $1$ to $3$ illustrate how both models encode patterns that were not present in the training dataset.
-Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
-Figure 8 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix F).
+Rows 111333 illustrate how both models encode patterns that were not present in the training dataset.
+Rows 444555 let one compare the models’ performance in denoising tasks, while row 666 displays their ability to reconstruct lines.
+Figure 8 compares performance between the proposed CNA model and an autoencoder (for model and training details, refer to Appendix F).
 While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training.
-For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
-Rows $4$ and $5$ highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced.
-However, as the noise level increases, its performance decreases significantly.
-The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations.
-Metrics on noise reduction, precision, and recall, presented in Appendix G, further substantiate the CNA model’s superior performance.
-In row $6$, the models’ capabilities for reconstructing missing line segments are compared.
+For instance, rows 111333 demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
+[…]
+In row 666, the models’ capabilities for reconstructing missing line segments are compared.
 Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed.
@@
 5 Relation to Alternative Concepts
-To grasp the significance of the approach, it is important to compare it to alternative approaches.
+The central issue behind our model is how the structure of the environment is represented.
+In biology, this structure would be accessed by the senses in the form of sensory patterns — large active subsets of sensory elements (e.g., retinal ganglion cells).
+While global sensory patterns vary across different conditions and viewpoints, local sensory structure tends to exhibit consistency.
+For instance, the sensory pattern associated with a static object may appear different under varying eye positions or perspectives, yet local fragments of the pattern remain stable and coherent (we demonstrate this in our experiments by having constant straight-line patterns that are stabilized, while changing Gaussian noise is ignored).
+The task of learning – both in biology and in a model like ours – is to pick up durable structures from sensory patterns in spite of their fleeting nature.
+We proceed here with a simple assumption regarding this durable structure: There is an important class of coherent sensory patterns that are composed of a finite set of overlapping pattern fragments.
+We assume that these pattern fragments can be represented by the recurrent connectivity of net fragments and that this connectivity can be learned from the statistics of sensory patterns with the help of Hebbian plasticity.
+This is the background against which we discuss some models of neural computation currently dominating the literature.
 5.1 Associative Memory
-Associative memory (AM) (Amari, 1972; Hopfield, 1982) and the model presented in this work share the objective of retrieving previously encountered patterns, both utilize Hebbian synaptic plasticity for encoding information.
-However, fundamental differences exist in their structural and functional organization.
-In classical AM, neurons exhibit an all-to-all connectivity pattern, leading to the representation of structures as monolithic wholes, whereas our model constrains connectivity within learned net fragments.
-This restriction enables a more modular representation of complex patterns by dynamically assembling net fragments into structured nets, thereby allowing the formation of representations that extend beyond the training data distribution.
-These differences arise from distinct learning mechanisms.
-In AM, a neural pattern is stored in a single event through the uniform strengthening of connections between co-activated neurons.
-Our approach, however, capitalizes on the observation that a large number of non-repeating patterns (as found in natural images) contain small recurring sub-patterns, such as texture patches.
-These recurring elements are captured within net fragments.
-The learning rule defined in eq.
-20 selectively reinforces connections corresponding to the most relevant sub-patterns.
-Specifically, frequent sub-patterns exhibit a consistently positive average correlation $\rho$, which leads to the progressive increase of their associated synaptic weights.
-In contrast, rare or non-informative patterns exhibit low or even negative average correlations, causing their associated weights to decay toward zero.
-This Hebbian mechanism effectively filters out noise and ensures that only the most salient and statistically dominant sub-patterns are preserved in the learned representation.
-A fundamental limitation of AM is the interference between stored patterns unless they are statistically independent (i.e., orthogonal), meaning that minimal neuron-wise overlap exists.
-Sparse coding, as proposed by Tsodyks and Feigel’man (2007) or Palm (2013), mitigates interference by activating only a small subset of neurons.
-Our model offers a more intrinsic solution by representing each input feature using multiple alternative neurons.
-These neurons possess identical receptive fields (as observed by Cossell et al.
-(2015)), but can exhibit different excitatory outputs.
-This redundancy enables overlapping net fragments to reduce neuron-wise interference by utilizing distinct alternative neurons.
+The central concept in the classic works of Hebb (1949) and Hayek (1952) is the assembly, a set of neurons supporting each other’s activity by mutual excitatory connections, connections that have been created by synaptic, “Hebbian”, plasticity.
+This concept has found a concrete mathematical formulation as “associative memory” (AM) (Amari, 1972; Hopfield, 1982), where stored patterns are encoded by enhancing excitatory connections between all pairs of co-active neurons.
+While AM mechanisms are effective for storing and retrieving known patterns, they often fail when applied to perceptual tasks such as object recognition, where previously unseen structures must be classified based on prior experience.
+This limitation arises because AM tends to store complete patterns as static “snapshots.” The reason for this rigidity of AM states is their mode of recording: The excitatory connections between all pairs of neurons active at a particular moment are strengthened simultaneously, without regard to whether a connection is significant in the sense of applying to many patterns (that is, being part of a repeating pattern fragment) or insignificant, applying only to this particular pattern.
+In our model, in contrast, the synaptic weights are gradually modified in response to the statistical properties of a large ensemble of sensory patterns.
+According to our Eq.
+30, a particular synapse can only grow if the pair of neurons it connects is simultaneously active again and again (indicated by positive correlation $\boldsymbol{\rho}_{avg}$), or it decays due to negative correlation.
+This processing allows the system to gradually extract and store re-occurring (statistically relevant) sub-patterns rather than memorizing isolated global patterns.
+A functional restriction on AM is that different stored states have to be statistically independent, that is, as sets of active neurons, they have to have low pairwise overlap.
+This is an awkward constraint, as sensory patterns are composed of repeating pattern fragments and thus are bound to have a large overlap.
+In classical AM, this constraint has been addressed just by requiring patterns to be small subsets of all neurons (Tsodyks and Feigel’man, 2007; Palm, 2013).
+Our model can address the small-overlap constraint in a systematic way with the help of multiple neurons (“competing neurons”) that respond to the same trigger feature but compete with each other for activity so that pairs of pattern fragments that have high feature-wise overlap can select different competing neurons, resulting in the corresponding net fragments having low neuron-wise overlap.
+This is made possible by the statistical mode of strengthening synapses and is not possible in AM with its one-shot learning mode.
+Our model also addresses an open issue of AM: The selection of coherent sensory patterns for storage.
+As coherence is defined as being composed of a fixed set of overlapping pattern fragments (stored in net fragments), it is only natural to require that the learning process converges to a connectivity state in which only sets of neurons that segment the input into coherent sensory patterns can be activated stably (von der Malsburg, 2024).
+The canonical formulation (Amari, 1972; Hopfield, 1982) of the AM’s state dynamic defines an “energy” that acts as a Lyapunov function.
+Beginning from any initial state, neurons are visited sequentially, and their state is flipped to gradually reduce the energy (simultaneous switching of the neurons would lead to non-sensical oscillations).
+Our model’s dynamics is not based on a Lyapunov function and a time-consuming gradual process, relying instead on an externally imposed periodic attenuation control parameter that lets the set of neurons originally activated by the sensory input collapse to a subset formed by neurons that support each other by excitatory connections.
 5.2 Boltzmann Machine
-Key aspects of our model become clearer when compared to the Boltzmann Machine (BM) introduced by Hinton and Sejnowski (1983).
-Both models allow recurrent connections (as also proposed in Mumford (2002)), but they also exhibit significant differences.
-Nevertheless, as our focus is limited to one area with net fragments ($S2$) and is not the full object detection architecture as proposed by von der Malsburg et al.
-(2022), we highlight only select differences here.
-The BM makes a clear distinction between visible and hidden (“latent”) neurons.
-In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by attenuation in favor of those neurons that support each other through previously learned connections.
-A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
-In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference being that we do not insist on symmetric connections and therefore do not derive an energy function (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the absence of up to half of the connections due to lack of symmetry does not jeopardize the stability of a state).
-However, even without an energy function, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
+Regarding Boltzmann Machines (BM), we refer to Hinton and Sejnowski (1983); Hinton et al.
+(1984), taking the latter as the authoritative reference.
+Both the BM and our model describe networks of recurrent connections between neurons, where some neurons get activated by sensory input (“visible units”), and by employing synaptic plasticity, both approaches aim to turn the system into a constraint satisfaction network that uses inner connections to stabilize the patterns suggested by the sensory input.
+The BM as described in Hinton et al.
+(1984) and our model differ, however, in a number of ways.
+Inspired by statistical mechanics, the BM speaks of equilibrium macro-states, each of which is realized as an ensemble of micro-states, the probability of a micro-state being proportional to a negative exponential of a state “energy” (Boltzmann distribution), which is defined as in Amari-Hopfield’s Associative Memory as a function of neural activities and connection weights.
+The BM further assumes that the system compares two modes, clamped (in which the sensory neurons are determined by the input) and free running (no sensory input), and lets synaptic plasticity equalize the signal correlations of pairs of neurons observed in one of the two modes with those observed in the other.
+In our model, neural activity is not modeled in terms of probabilistic ensembles of states but as discrete-time trajectories over a fixed number of time steps $T$, during which the same sensory input is constantly refined.
+These trajectories evolve deterministically according to update rules that depend on both the input and recurrent connectivity.
+Importantly, the dynamics are externally modulated: At each step, a control mechanism attenuates the activity of neurons that lack sufficient recurrent support from other active neurons; see Eqs.
+28 and 29.
+In contrast, Boltzmann Machines rely on stochastic sampling over ensembles of network states, governed by a distribution derived from an energy function.
+While the statistical mechanics framework underlying Boltzmann Machines is elegant and conceptually rich, it is computationally inefficient in settings where a large number of intermediate states must be generated and compared to approximate equilibrium.
+Our model avoids this bottleneck by converging rapidly through deterministic dynamics over a small fixed number of steps.
+A necessary precondition for the feasibility of the goals of the BM and our model is the existence of a system of constraints on the sensory patterns that can be learned from a finite number of such patterns and which then suffice to stabilize all future patterns.
+Given the bias-variance dilemma (Geman et al., 1992) and the no-free-lunch theorems (Wolpert, 1996), a learning system’s architecture needs to drastically constrain its search space by a bias that is tuned to the application domain (Mitchell, 1997).
+No such bias is built into the BM’s architecture in its generality  [^1] The concrete application described in Hinton et al.
+(1984) (the ‘Encoder Problem’) provides such bias, but is trivial in the sense of explicitly defining the full (and very small) set of sensory patterns., while our model’s architecture explicitly incorporates the bias that the sensory patterns of its application domain, vision, are all coherent composites of a finite set of pattern fragments, a bias that is effective yet captures an essentially infinite number of sensory patterns.
+The realization that the BM in its generality would imply unfeasible learning times led to the introduction of Restricted Boltzmann Machines (RBM) (Smolenksy, 1987; Hinton and Salakhutdinov, 2006).
+RBMs consist of two layers, one visible and one latent, with bidirectional connections between them, but no connections within them.
+A stack of RBMs is used in (Hinton and Salakhutdinov, 2006) to construct an autoencoder in which RBM learning is used to pre-train the connections between layers in a sequence from bottom (input) to top (the sparse coding layer).
+Then, the whole encoder-decoder stack is trained end-to-end by backpropagation of error.
+We discuss the differences between our model and networks trained by backpropagation of error (including RBMs) in the next sub-section.
 5.3 Artificial Neural Networks (ANN)
-Artificial neural networks (ANNs) rely on various architectures and learning techniques, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), and autoencoders (Hinton and Salakhutdinov, 2006).
-These systems process sensory input through multiple layers to generate a neuronal representation.
-A defining characteristic of ANN systems is their strict feedforward architecture, where connections from input to representation and from representation to output are trained via backpropagation of error (Rosenblatt, 1962, p.
-292).
-In contrast, the neuronal representation within the CNA consists of nets that activate as a whole, with individual neurons depending on the activation of others in the same net.
-Thus, while ANNs rely on a global structuring of neuronal activations, the CNA relies on a local self-organizing process.
-Moreover, ANNs rely on hierarchical layers where neurons respond to features of increasing complexity.
-The CNA replaces this hierarchy with a single layer, where net fragments of different sizes capture features of varying complexity.
-These fragments interconnect to form larger nets that represent entire objects, naturally solving the binding problem, whereas ANNs do not encode explicit relationships between features, leading to ambiguities such as hallucinated feature conjunctions.
-Furthermore, input representation in the CNA emerges from statistical regularities in the input, making it independent of specific tasks.
-In contrast, ANNs bias feature extraction to serve predefined training objectives, which limits their adaptability when retraining for different functions.
+Artificial neural networks (ANNs), as concerned here, are layered structures with feedforward layer-to-layer connections from input to output, and these connections are trained by back-propagation of error (Rosenblatt, 1962, p 292).
+Prominent examples are multilayer perceptrons (MLPs) (Prince, 2023), CNNs (Lecun et al., 1998), autoencoders (Hinton and Salakhutdinov, 2006), and transformers (Vaswani et al., 2017).
+In these systems, individual neurons in one layer represent feature compositions that occur in the layer below.
+The complexity and variability of features represented by individual neurons grow from layer to layer (as shown in reconstructions, see Zeiler and Fergus (2013)).
+The hierarchy of layers thus constitutes a hierarchy of nested pattern fragments.
+The architectures of both ANNs and our model implement the bias that a virtually infinite set of potential input patterns can be represented in the form of flexible compositions of a finite set of pattern fragments.
+They differ fundamentally, however, in how these pattern fragments are implemented, how they are combined, and how they are learned.
+Beginning with the latter, ANNs select the features to be represented on the basis of their effect on lowering the system’s prediction error, whereas in our model the selection of features to be represented by net fragments is entirely a matter of the statistics of input patterns (as is, incidentally, the case with the RBM pretraining of layer connections in Hinton and Salakhutdinov (2006)).
+Regarding implementation, ANNs represent elementary pattern fragments in the first latent layer by feature neurons.
+In our model, which has only one layer, elementary pattern fragments are represented by minimal net fragments, each composed of a neuron and a number of excitatory connections from other neurons within the same layer.
+That a neuron remains active under attenuation indicates that it has recognized the pattern formed by these other also activated neurons.
+ANNs represent a nested hierarchy of pattern fragments, as higher-level neurons respond to combinations of active lower-level neurons.
+Our model equally represents a nested hierarchy of pattern fragments within one layer, in this case composed of a neuron, its immediate supporters, the supporters of the supporters, and so on.
+The rules for composing larger structures out of smaller pattern fragments correspondingly differ between ANNs and our model.
+In CNNs, this flexibility is mostly implemented by max pooling, or in transformers by the attention mechanism that selects active neurons in one layer to contribute to the activity of neurons in the next.
+This freedom of choosing pattern fragments in lower levels is exerted independently by the neurons in a given level, without regard to the mutual compatibility of those selected lower-level patterns.
+This lack of coherence is even more drastic between ANNs’ high-level feature neurons, which don’t express in any way mutual overlap in terms of constituent sub-patterns (the binding problem, see Greff et al.
+(2020)), the likely cause of vulnerability of ANNs to adversarial attack (Leadholm and Stringer, 2022).
+In our model, in contrast, the net fragments that respond to different pattern fragments have to fit together into one consistent net, in which each active neuron is sufficiently supported by other active neurons.
+This expresses the bias that significant patterns in the visual environment are composed of pattern fragments in a coherent way.
+We consider this as the solution to the binding problem (Greff et al., 2020) and as an explanation of the Gestalt phenomenon (Wagemans et al., 2012).
 6 Discussion and Outlook
-The CNA presented here establishes a novel approach for constructing coherent nets as representations of input data.
-This framework enables the encoding of complex patterns through the composability of previously learned sub-patterns (net fragments).
-Compared to existing neural coding frameworks, CNA offers distinct advantages, particularly in terms of robustness – demonstrated through its ability to filter noise and reconstruct figures – and its capacity to generalize to out-of-domain data, as evidenced by its capability to represent previously unseen objects.
-For future research, we outline two key directions: First, an essential next step is the extension of CNA to natural image domains.
-This would allow net fragments to capture and encode texture information encountered in real-world scenarios, thus realizing the ambition expressed in Figure 1.
-Second, the current implementation of CNA is restricted to a single processing area with net fragments.
-Following the theoretical framework proposed by von der Malsburg et al.
-(2022), the architecture could be expanded to incorporate multiple areas, thereby enabling object recognition while achieving invariance to position, scale, and in-plane rotation.
-Object recognition could be realized using structure-sensitive, homeomorphic mappings between nets (von der Malsburg, 1981; Anderson and van Essen, 1987; Olshausen et al., 1995; Wiskott et al., 1997; Wolfrum et al., 2008; Memisevic and Hinton, 2010; Arathorn et al., 2013), which would facilitate the mapping of input representations onto a more stable schematic representation of an object class or an exemplar of the object to be recognized.
-The adoption of homeomorphic mappings, as opposed to alternative methods, preserves the core principle of nets and net fragments within and across multiple processing areas, thereby maintaining structural coherence in representation.
-The proposed neural coding framework not only advances representational robustness but also aligns with principles of self-organization observed in cortical processing.
-Its ability to learn and flexibly recombine net fragments without explicit supervision suggests a pathway toward more adaptive and interpretable neural representations.
-Future extensions to multi-area architectures could bridge the gap between bottom-up feature learning and top-down predictive coding.
-These developments position CNA as a foundational step toward more structured and generalizable neural computation.
+The Cooperative Network Architecture (CNA) introduced here proposes a novel approach to neural representation by constructing structured, composable nets from input patterns.
+It is based on the vision to overcome the tension that has eternally been plaguing the brain/mind issue and AI, the tension between low-level and high-level, between sub-symbolic and symbolic descriptions (cf.
+Fodor and Pylyshyn (1988)).
+Both perspectives have their strengths and their weaknesses (as discussed by von der Malsburg (2023)).
+While our model preserves global, high-level pattern structure through the composition of net fragments (the symbolic aspect), it also retains sensitivity to local detail via the fine-grained connectivity of its elements (the sub-symbolic aspect).
+This synthesis of descriptive levels is the basis for perceptual segmentation (as in the Gestalt phenomenon), noise filtering, pattern completion, and for pattern-global decisions such as classification or recognition (see below).
+Empirically, we demonstrated that CNA generalizes beyond its training distribution, successfully reconstructing unseen objects and filtering substantial noise.
+These findings validate CNA’s ability to form holistic, yet flexible, representations.
+Furthermore, the reuse and recombination of learned sub-patterns provide a strong inductive bias, allowing for compositional generalization from limited data.
+Conceptually, our model deviates in essential ways from classical neural models.
+In contrast to the perceptron (which is currently dominating AI), the structure of patterns is expressed by recurrent and not feedforward connections, the entirety of patterns is expressed as structured nets instead of by structureless individual neurons, and learning of connectivity is based on Hebbian plasticity and input statistics instead of backpropagation of error and human-provided sample (and ground truth) material.
+In contrast to the associative memory model, patterns have a flexible articulated structure, being composed of generic pattern fragments, instead of monolithic rigidity.
+From the Boltzmann Machine, our model differs by being tuned to the natural environment by a powerful bias, which permits to capture a virtual infinity of patterns after learning from a finite set of examples.
+What is the functional role of nets within intelligent systems?
+According to a long-standing proposal, invariant object recognition can be realized by structure-sensitive, homeomorphic mappings between networks (von der Malsburg, 1981; Anderson and van Essen, 1987; Olshausen et al., 1995; Wiskott et al., 1997; Arathorn, 2002; Wolfrum et al., 2008; Memisevic and Hinton, 2010; von der Malsburg et al., 2022).
+This calls for the existence of a pair or a series of layers, the lowest of which is the sensory layer we have modeled here, while higher layers harbor nets of growing invariance or generality.
+Nets in higher layers can be invariant against transformations such as translation, rotation, or scaling in the image plane, or can be more abstract by leaving out submodalities such as color or texture, or by accommodating distortion.
+Dynamic mappings that can keep up with rapid changes within the input layer are realized in the cited literature by the rapid selection of connections that connect neurons of equal feature type and that preserve the neighborhood relationships within the connected nets.
+Homeomorphic mappings can themselves be considered as nets of a second kind, and they also are coherently composed of overlapping map fragments (Anderson and van Essen, 1987; Zhu and von der Malsburg, 2004).
+Several nets tied together by mappings can be considered as nets on a higher level.
+The purpose of the model we have presented here is to fill a sore gap in this line of work in providing a clear picture of how nets that are eligible for mapping are dynamically formed.
+Urgent future work will have to expand the present model to cope with a full range of natural visual input and a rich range of visual submodalities (texture, color, stereo depth, motion), will have to expand to other modalities (e.g., audition, motor control), and will have to combine the functional elements modeled here and in the cited work into fully functional systems.
 Funding
@@
 Appendix A Feature Extraction
-Figure 9: The hand-crafted convolutional kernels of the feature extractor.
+Figure 9: Features: The hand-crafted convolutional kernels of the feature extractor.
 The stage $S1$ serves to extract initial binary features from the images.
 The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location.
-We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal).
+We utilize hand-crafted filters with a size of $\times 5$ 5 as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal).
 This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons).
@@
 While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
-These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
-Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$.
+These filters move across the image with a step size (‘stride’) of 111 (assuming 0 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
+Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size (1×32×32)(1$\times$ 32$\times$ 32)( 1 × 32 × 32 ) and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size (4×32×32)(4$\times$ 32$\times$ 32)( 4 × 32 × 32 ).
 The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels.
@@
 neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq.
-12).
+18).
 Appendix B Normalization
-The membrane potential in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{M}(S2)}$ to the interval $(0,1)$.
+The pre-synaptic activity in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{I}(S2)}$ to the interval (0,1)(0,1)( 0 , 1 ).
 We first apply a size limit:
-$\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (21)
-According to this formula, $\boldsymbol{a}^{\mathcal{M}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{M}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$.
+$\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (31)
+According to this formula, $\boldsymbol{a}^{\mathcal{I}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{I}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope 1/2-1/2- 1 / 2.
 (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well).
@@
 In the next step, we introduce the normalization
-$\boldsymbol{a}^{\text{norm }(S2)}_{c,i,t}=\max\left(\frac{\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i,t}}{\max_{i^{\prime}}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i^{\prime},t}},0\right)$ (22)
+$\boldsymbol{a}^{\text{norm }(S2)}_{c,i,t}=\max\left(\frac{\boldsymbol{a}^{\mathcal{I}(S2)}_{c,i,t}}{\max_{i^{\prime}}\boldsymbol{a}^{\mathcal{I}(S2)}_{c,i^{\prime},t}},0\right)$ (32)
 Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$.
-This normalization confines activations to the interval $[0,1]$, where neurons receiving maximal support are mapped to $1$, while those with less support are mapped to correspondingly smaller values.
+This normalization confines activations to the interval [0,1][0,1][ 0 , 1 ], where neurons receiving maximal support are mapped to 111, while those with less support are mapped to correspondingly smaller values.
 This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.
@@
 For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}^{(S1)}$ and $\boldsymbol{y}^{(S2)}_{t-1}$, leading to the equivalent formulation of eq.
-16:
-$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(S2)}_{c_{\text{out}},c_{\text{in}}}\star\left(\boldsymbol{y}_{c_{\text{in}}}^{(S1)};\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)$ (23)
+22:
+$\boldsymbol{a}^{\mathcal{I}(S2)}_{c_{\text{out}},t}=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(S2)}_{c_{\text{out}},c_{\text{in}}}\star\left(\boldsymbol{y}_{c_{\text{in}}}^{(S1)};\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)$ (33)
 Appendix D Dataset
-The dataset used for training comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
+The dataset used for training comprises binary images measuring ($\times 32$ 32) pixels, each sample depicting a straight line going through the image center, starting and ending 222 pixels from the image boundary (leading to 595959 distinct lines of different angles).
 The images are generated when required and may vary from one epoch to another.
-During each training cycle, we randomly generate $300$ image instances (each line about $5$ times).
-During evaluation, we sample each of the $59$ distinct lines exactly once.
+During each training cycle, we randomly generate 300300300 image instances (each line about 555 times).
+During evaluation, we sample each of the 595959 distinct lines exactly once.
 In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
-To introduce additive Gaussian noise to the afferent input in stage $S2$, we probabilistically flip neurons within each channel with a probability of up to $20\%$.
-Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location.
-To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
+[…]
 To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training.
-Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
-Additionally, we generate binary images measuring ($64\times 64$) pixels by hand that represent more complex structures.
-These images include line drawings of digits 00-$9$, characters, and objects such as a house, a star, a wheel, a clock, etc.
+Specifically, we produce binary images measuring ($\times 32$ 32) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
+Additionally, we generate binary images measuring ($\times 64$ 64) pixels by hand that represent more complex structures.
+These images include line drawings of digits 0-999, characters, and objects such as a house, a star, a wheel, a clock, etc.
 Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
-Appendix E Parameters
-We train the model with a learning rate of $\eta=0.2$ for $100$ training cycles (epochs).
-The model uses $\kappa=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons.
-Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$.
+Appendix E Training Parameters
+We train the model with a learning rate of $\eta=0.2$ for 100100100 training cycles (epochs).
+The model uses $\kappa=10$ competitive neurons with recurrent connections spanning a kernel of $\times 11$ 11 neurons.
+Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of competitive channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$.
 The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
-Appendix F Autoencoder Parameter
+Appendix F Autoencoder Training
 We compare our CNA model to an ANN, specifically an autoencoder (Hinton and Salakhutdinov, 2006), which is a widely used self-supervised technique for generating input representations.
@@
 The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model.
-It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ channels.
-The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively.
+It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of 333 and a stride of 222, with progressively increasing channel counts of 323232, 646464, 128128128, and 256256256 channels.
+The decoder mirrors this structure, using transposed convolutional layers with 256256256, 128128128, 646464, and 323232 channels, respectively.
 ReLU activations are applied after each layer.
-Consequently, the autoencoder is a relatively large model with $777$k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
-The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of $(0.9,0.999)$, and a weight decay of $1\times 10^{-8}$.
+Consequently, the autoencoder is a relatively large model with 777777777k parameters, utilizing up to 256256256 channels, whereas the CNA model operates with only 444444 channels.
+The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 ), and a weight decay of $1\times 10^{-8}$.
 To provide a robust comparison, we increase the number of training samples for the autoencoder.
-Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for $200$ epochs, using $10,000$ images per epoch.
+Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for 200200200 epochs, using 10,00010,00010 , 000 images per epoch.
 This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction.
-The autoencoder is trained with a batch size of $512$.
+The autoencoder is trained with a batch size of 512512512.
 Appendix G Quantifying Noise Filtration
@@
 This aligns with the findings of Ahmad and Hawkins (2015), highlighting that binary distributed activations demonstrate high robustness, retaining accurate interpretation even in the presence of numerous flipped neurons.
-In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$.
-Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines, which is the case for the CNA but not for the autoencoder (e.g., see Figure 10).
+[…]

### v5

@@
 We introduce the Cooperative Network Architecture (CNA), a model that represents sensory signals using structured, recurrently connected networks of neurons, termed “nets.” Nets are dynamically assembled from overlapping net fragments, which are learned based on statistical regularities in sensory input.
-This architecture offers robustness to noise, deformation, and out-of-distribution data, addressing challenges in current vision systems from a novel perspective.
+This architecture offers robustness to noise, deformation, and generalization to out-of-distribution data, addressing challenges in current vision systems from a novel perspective.
 We demonstrate that net fragments can be learned without supervision and flexibly recombined to encode novel patterns, enabling figure completion and resilience to noise.
@@
 Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Sager et al., 2025).
-Yet, despite these advances, artificial vision systems continue to struggle with robust object representation and recognition, particularly under varying viewing conditions, occlusion, or noise (Qi et al., 2023; Simmler et al., 2021).
+Yet, despite these advances, artificial vision systems continue to struggle with robust object representation and recognition, particularly under varying viewing conditions, occlusion, or noise (Fan et al., 2023; Simmler et al., 2021).
 This limitation highlights a fundamental gap between current AI systems and the flexible, context-aware perception observed in biological vision.
@@
 $\forall w_{i,j}\quad n_{j}\in\mathcal{B}_{r}(n_{i})$ (2)
-Additionally, we limit the weights to a range between 0 and 111, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
+Additionally, we limit the weights to a range between 0 and $1$, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
 $w_{i,j}\in[0,1]$ (3)
@@
 The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment.
-We call this accumulated input the pre-synaptic activity at timestep $t$, which is a function $\mathcal{I}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
+We call this accumulated input the pre-synaptic activity at timestep $t$, which is a function $\mathcal{I}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep t1t-1.
 $\boldsymbol{a}^{\mathcal{I}}_{i,t}=\mathcal{I}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
@@
 Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$).
-A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects.
+A set of directly and indirectly supported neurons is called a net, implementing the basic structure used in this work to represent objects.
 For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active as well:
@@
 Consequently, this stage serves as a preprocessing step, enabling the transformation of continuous-valued images into a suitable binary representation.
-The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels (111 for grayscale, 333 for RGB), while $H,W$ stand for the height and width of the image.
+The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels ($1$ for grayscale, $3$ for RGB), while $H,W$ stand for the height and width of the image.
 Activation Computation.
@@
 $\mathcal{B}_{r}\left(n^{(S2)}_{i}\right)$ $=\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r\right\}$ (21)
-$\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)$ =$=\left\{n_{j}^{(S1)}\mid d(n_{i}^{(S2)},n_{j}^{(S1)})\leq r\right\}$
+$\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)$ $=\left\{n_{j}^{(S1)}\mid d(n_{i}^{(S2)},n_{j}^{(S1)})\leq r\right\}$
 where $r$ defines the radius of lateral connections.
@@
 where $\boldsymbol{z}_{h^{\prime},w^{\prime},t}$ is the concatenated input patch:
-$\boldsymbol{z}_{h^{\prime},w^{\prime},t}\;=\;\begin{bmatrix}\boldsymbol{y}^{(S1)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W}\\ \boldsymbol{y}^{(S2)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W,\,t-1}\end{bmatrix}.$ (26)
+$\boldsymbol{z}_{h^{\prime},w^{\prime},t}\;=\;\begin{bmatrix}\boldsymbol{y}^{(S1)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W}\\ \boldsymbol{y}^{(S2)}_{:,\,h^{\prime}:h^{\prime}+H,\,w^{\prime}:w^{\prime}+W,\,t-1}\end{bmatrix}$.
+(26)
 Thus, at each spatial location and time step, the index $k^{*}\in\{1,\dots,\kappa\}$ maximizing the correlation between the corresponding filter and the local input patch is chosen.
@@
 The pre-synaptic activity is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{I}},\mathcal{E}\right)$.
-In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{I}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range [0,1][0,1][ 0 , 1 ] (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active.
+In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{I}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range $[0,1]$ (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active.
 Specifically, attenuation is implemented as:
@@
 The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all competitive $S2$ neurons in the same position.
-This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to 111 if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of competitive neurons.
-The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to 111.
+This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to $1$ if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of competitive neurons.
+The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to $1$.
 Learning of Weights.
@@
 The top of the figure illustrates the straight lines used during the training phase.
-Below, various samples from the evaluation datasets are shown, including kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888).
+Below, various samples from the evaluation datasets are shown, including kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows 5588).
 For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
 The model weights are trained using straight lines (see training data examples at the top of Figure 4).
-During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 111888 of Figure 4).
-Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888).
+During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 1188 of Figure 4).
+Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows 5588).
 These results demonstrate that while individual net fragments $\mathcal{F}_{t}$ must be learned (in our case we use Hebbian learning to train the connections within $\mathcal{F}_{t}$), the recursive composition described in Eq.
@@
 Figure 5: Noise filtering efficacy: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
-[…]
-[…]
-[…]
+To further investigate noise filtering, we introduce Gaussian noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - resulting from adding up to $20\%$ noise per channel across four independent channels; see Appendix D) to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
+The results are visualized in Figure 5: While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated.
+As additional noise is introduced, more undesired neurons receive support and persist in their activity.
+However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$.
+In Appendix G, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
+Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed.
+However, when attenuation increases over time, even more noise is filtered out.
+The plot is created after adding $10\%$ noise to each feature channel (using a bias $b^{(S2)}=0.7$).
 The reason that the filtering works is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time.
@@
 Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$.
-Part B: A diagonal line with 666 missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the pre-synaptic activity $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
+Part B: A diagonal line with 66 missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the pre-synaptic activity $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
 If a net fragment $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$).
 Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns.
-This reconstruction is reliable for up to 333 missing pixels (see Appendix G).
-In some cases, as in this figure, the line is fully reconstructed for up to 666 removed pixels and partly reconstructed for 777 pixels.
+This reconstruction is reliable for up to $3$ missing pixels (see Appendix G).
+In some cases, as in this figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
 Pre-Synaptic Activity Represents Expected Feature Activation.
@@
 Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder).
-Rows 111333 illustrate how both models encode patterns that were not present in the training dataset.
-Rows 444555 let one compare the models’ performance in denoising tasks, while row 666 displays their ability to reconstruct lines.
+Rows 1133 illustrate how both models encode patterns that were not present in the training dataset.
+Rows 4455 let one compare the models’ performance in denoising tasks, while row 66 displays their ability to reconstruct lines.
 Figure 8 compares performance between the proposed CNA model and an autoencoder (for model and training details, refer to Appendix F).
 While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training.
-For instance, rows 111333 demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
-[…]
-In row 666, the models’ capabilities for reconstructing missing line segments are compared.
+For instance, rows 1133 demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
+Rows 4455 highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced.
+However, as the noise level increases, its performance decreases significantly.
+The CNA model, in contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations.
+Metrics on noise reduction, precision, and recall, presented in Appendix G, further substantiate the CNA model’s superior performance.
+In row $6$, the models’ capabilities for reconstructing missing line segments are compared.
 Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed.
@@
 In our model, neural activity is not modeled in terms of probabilistic ensembles of states but as discrete-time trajectories over a fixed number of time steps $T$, during which the same sensory input is constantly refined.
-These trajectories evolve deterministically according to update rules that depend on both the input and recurrent connectivity.
+In distinction to the Boltzmann Machine, these trajectories evolve deterministically, according to update rules that depend on both the input and recurrent connectivity.
 Importantly, the dynamics are externally modulated: At each step, a control mechanism attenuates the activity of neurons that lack sufficient recurrent support from other active neurons; see Eqs.
@@
 (1984) (the ‘Encoder Problem’) provides such bias, but is trivial in the sense of explicitly defining the full (and very small) set of sensory patterns., while our model’s architecture explicitly incorporates the bias that the sensory patterns of its application domain, vision, are all coherent composites of a finite set of pattern fragments, a bias that is effective yet captures an essentially infinite number of sensory patterns.
-The realization that the BM in its generality would imply unfeasible learning times led to the introduction of Restricted Boltzmann Machines (RBM) (Smolenksy, 1987; Hinton and Salakhutdinov, 2006).
+The realization that the BM in its generality would imply unfeasible learning times led to the introduction of Restricted Boltzmann Machines (RBM) (Smolensky, 1987; Hinton and Salakhutdinov, 2006).
 RBMs consist of two layers, one visible and one latent, with bidirectional connections between them, but no connections within them.
@@
 The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location.
-We utilize hand-crafted filters with a size of $\times 5$ 5 as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal).
-This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons).
+We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal).
+This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $45^{\circ}$ diagonal feature neurons).
 The use of such manually designed filters allows a more straightforward interpretation of the results.
 While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
-These filters move across the image with a step size (‘stride’) of 111 (assuming 0 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
-Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size (1×32×32)(1$\times$ 32$\times$ 32)( 1 × 32 × 32 ) and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size (4×32×32)(4$\times$ 32$\times$ 32)( 4 × 32 × 32 ).
+These filters move across the image with a step size (‘stride’) of $1$ (assuming 0 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size.
+Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$.
 The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels.
@@
 Appendix B Normalization
-The pre-synaptic activity in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{I}(S2)}$ to the interval (0,1)(0,1)( 0 , 1 ).
+The pre-synaptic activity in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{I}(S2)}$ to the interval $(0,1)$.
 We first apply a size limit:
 $\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (31)
-According to this formula, $\boldsymbol{a}^{\mathcal{I}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{I}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope 1/2-1/2- 1 / 2.
+According to this formula, $\boldsymbol{a}^{\mathcal{I}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{I}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope 1/2-1/2.
 (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well).
@@
 Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$.
-This normalization confines activations to the interval [0,1][0,1][ 0 , 1 ], where neurons receiving maximal support are mapped to 111, while those with less support are mapped to correspondingly smaller values.
+This normalization confines activations to the interval $[0,1]$, where neurons receiving maximal support are mapped to $1$, while those with less support are mapped to correspondingly smaller values.
 This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.
@@
 Appendix D Dataset
-The dataset used for training comprises binary images measuring ($\times 32$ 32) pixels, each sample depicting a straight line going through the image center, starting and ending 222 pixels from the image boundary (leading to 595959 distinct lines of different angles).
+The dataset used for training comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
 The images are generated when required and may vary from one epoch to another.
-During each training cycle, we randomly generate 300300300 image instances (each line about 555 times).
-During evaluation, we sample each of the 595959 distinct lines exactly once.
+During each training cycle, we randomly generate $300$ image instances (each line about $5$ times).
+During evaluation, we sample each of the $59$ distinct lines exactly once.
 In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
-[…]
+To introduce additive Gaussian noise to the afferent input in stage $S2$, we probabilistically flip neurons within each channel with a probability of up to $20\%$.
+Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location.
+To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
 To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training.
-Specifically, we produce binary images measuring ($\times 32$ 32) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
-Additionally, we generate binary images measuring ($\times 64$ 64) pixels by hand that represent more complex structures.
-These images include line drawings of digits 0-999, characters, and objects such as a house, a star, a wheel, a clock, etc.
+Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
+Additionally, we generate binary images measuring ($64\times 64$) pixels by hand that represent more complex structures.
+These images include line drawings of digits 0-99, characters, and objects such as a house, a star, a wheel, a clock, etc.
 Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
 Appendix E Training Parameters
-We train the model with a learning rate of $\eta=0.2$ for 100100100 training cycles (epochs).
-The model uses $\kappa=10$ competitive neurons with recurrent connections spanning a kernel of $\times 11$ 11 neurons.
+We train the model with a learning rate of $\eta=0.2$ for $100$ training cycles (epochs).
+The model uses $\kappa=10$ competitive neurons with recurrent connections spanning a kernel of $11\times 11$ neurons.
 Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of competitive channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$.
@@
 The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model.
-It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of 333 and a stride of 222, with progressively increasing channel counts of 323232, 646464, 128128128, and 256256256 channels.
-The decoder mirrors this structure, using transposed convolutional layers with 256256256, 128128128, 646464, and 323232 channels, respectively.
+It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ channels.
+The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively.
 ReLU activations are applied after each layer.
-Consequently, the autoencoder is a relatively large model with 777777777k parameters, utilizing up to 256256256 channels, whereas the CNA model operates with only 444444 channels.
-The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 ), and a weight decay of $1\times 10^{-8}$.
+Consequently, the autoencoder is a relatively large model with 777777k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
+The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of $(0.9,0.999)$, and a weight decay of $1\times 10^{-8}$.
 To provide a robust comparison, we increase the number of training samples for the autoencoder.
-Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for 200200200 epochs, using 10,00010,00010 , 000 images per epoch.
+Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for $200$ epochs, using $10,000$ images per epoch.
 This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction.
-The autoencoder is trained with a batch size of 512512512.
+The autoencoder is trained with a batch size of $512$.
 Appendix G Quantifying Noise Filtration
@@
 The metrics demonstrate that increasing the attenuation coefficient or the activation bias results in enhanced noise filtration.
-[…]
+The overall efficacy of noise reduction remains consistently high across most parameter configurations, exhibiting a constant noise reduction rate of $\geq\approx 95\%$.
+This rate persists regardless of the magnitude of the introduced noise.
+As the noise suppression rate is approximately constant, this means that if there is more noise in the input, there will also be more noise in the output, albeit limited to around 5% of the input noise.
 Furthermore, all CNA models (except the one with the lowest power factor and bias) outperform the autoencoder, demonstrating that they achieve performance that is competitive with state-of-the-art deep learning models.
@@
 This aligns with the findings of Ahmad and Hawkins (2015), highlighting that binary distributed activations demonstrate high robustness, retaining accurate interpretation even in the presence of numerous flipped neurons.
-[…]
+In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$.
+Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines, which is the case for the CNA but not for the autoencoder (e.g., see Figure 10).